### v2

# and knot coloring invariants

Abstract.
Generalized Legendrian racks are nonassociative algebraic structures based on the Legendrian Reidemeister moves. We study algebraic aspects of GL-racks and coloring invariants of Legendrian links.
We answer an open question characterizing the group of GL-structures on a given rack. As applications, we classify several infinite families of GL-racks. We also compute automorphism groups of dihedral GL-quandles and the categorical center of GL-racks.
Then we construct an equivalence of categories between racks and GL-quandles.
We also study tensor products of racks and GL-racks coming from universal algebra. Surprisingly, the categories of racks and GL-racks have tensor units. The induced symmetric monoidal structure on medial racks is closed, and similarly for medial GL-racks.
Answering another open question, we use GL-racks to distinguish Legendrian knots whose classical invariants are identical. In particular, we complete the classification of Legendrian $8_{13}$ knots.
Finally, we use exhaustive search algorithms to classify GL-racks up to order 8.
Key words and phrases:
Algebraic theory, center, coloring invariant, Legendrian knot, medial quandle, rack, tensor product
2020 Mathematics Subject Classification:
Primary 08A35, 57K12; Secondary 18B99, 20B25, 57K10, 57K33
1. Introduction
Generalized Legendrian racks, also called GL-racks or bi-Legendrian racks, are a nonassociative algebraic structure used to distinguish Legendrian links in $\mathbb{R}^{3}$ or $S^{3}$. GL-racks can be traced back to algebraic structures called kei, which Takasaki [takasaki] introduced in 1942 to study symmetric spaces; quandles, which Joyce [joyce] introduced in 1982 to study conjugation in groups and links in $\mathbb{R}^{3}$ and $S^{3}$; and racks, which Fenn and Rourke [fenn] introduced in 1992 to study framed links in $3$-manifolds. Kei, quandles, and racks have enjoyed significant study as link invariants in geometric topology and in their own rights in quantum algebra and group theory.
More recently, various authors have introduced variants of racks suitable for studying Legendrian links. In 2017, Kulkarni and Prathamesh [original] introduced rack invariants of Legendrian knots. In 2021, Ceniceros et al. [ceniceros] refined these invariants by introducing Legendrian racks. In 2023, Karmakar et al. [karmakar] and Kimura [bi] independently strengthened these constructions by introducing GL-racks, which are racks equipped with additional structure.
In this paper, we further develop the theory and applications of GL-racks, answer open questions about GL-racks and Legendrian knot classifications, and classify GL-racks up to order $8$. In particular, we provide the first examples of GL-racks that distinguish Legendrian knots that are not distinguishable by their classical invariants, answering a question of Kimura [bi]*Section 4.
1.1. The structure of this paper
In Section 2, we define racks and quandles, consider several examples, and discuss a canonical rack automorphism $\theta$ that plays a fundamental role in the theory.
In Section 3, we give a simplified definition of GL-racks, show its equivalence to the definition in the literature, and discuss examples and universal-algebraic aspects of GL-racks.
Sections 46 study GL-racks from group-theoretic, categorical, and universal-algebraic perspectives, respectively; readers only interested in applications to Legendrian knot theory may safely skip these sections. In Section 4, we determine all GL-structures on a given rack, answering a question of Karmakar et al. [karmakar]*Section 3. As applications, we classify all GL-structures on permutation racks, conjugation quandles of abelian and centerless groups, Takasaki kei of abelian $2$-torsion-free groups, and dihedral quandles. We also characterize GL-rack automorphism groups and compute them for dihedral quandles. Finally, we compute the centers of the categories of GL-racks, GL-quandles, Legendrian racks, and Legendrian quandles.
In Section 5, we construct an equivalence of categories between the categories of racks and GL-quandles. This equivalence restricts to an equivalence between the full subcategories of medial racks and medial GL-quandles. Our construction corresponds to an isomorphism of algebraic theories.
In Section 6, we study tensor products of racks and GL-racks coming from universal algebra. We show that the categories of racks and GL-racks have tensor units. This is unusual for noncommutative algebraic theories; for example, the tensor product in the category of groups does not have a tensor unit. We show that the induced symmetric monoidal structure on medial racks is closed, and similarly for medial GL-racks.
In Section 7, we discuss the applications of GL-racks to Legendrian knot theory. We use GL-rack coloring invariants to distinguish several conjecturally distinct Legendrian knots with the same classical invariants. This gives a positive answer to a question of Kimura [bi]*Section 4; we also show in Appendix A.3 that Legendrian racks give a positive answer. In particular, we add to the classification of Legendrian $8_{10}$ knots and complete the classification of Legendrian $8_{13}$ knots in the extended Legendrian knot atlases of Bhattacharyya et al. [atlas-site] and Petkova and Schwartz [new-atlas].
In Section 8, we propose questions for further research based on our results.
In Appendix A, we describe algorithms that can classify finite GL-racks of a given order up to isomorphism. These algorithms use Vojtěchovský and Yang’s [library] classification of racks up to order $11$. We provide implementations of our algorithms in GAP [GAP4] and the data we were able to compute and enumerate for all $n\leq 8$. We also provide an algorithm that computes GL-rack coloring invariants of Legendrian links. As an application, we show that Legendrian racks positively answer the question of Kimura [bi]*Section 4 answered in Section 7 for GL-racks.
In Appendix B, we use the algorithms in Appendix A.2 to tabulate all GL-racks of orders $2\leq n\leq 4$ up to isomorphism. Due to length considerations, we give the tabulations for $5\leq n\leq 8$ in a GitHub repository [my-code].
Acknowledgments.
I am indebted to Sam Raskin, my undergraduate thesis adviser, for his mentorship, support, and many insights throughout the research and writing processes. I thank Samantha Pezzimenti and Wing Hong Tony Wong for introducing me to Legendrian knot theory and mathematics research, Jose Ceniceros for informing me of the categorical literature on racks, and Peyton Wood for introducing me to GL-racks. I also thank Noam Scully for helpful discussions, Qiaochu Yuan for a terminological correction, and an anonymous Stack Exchange user for referencing me to [alg-theories]*Theorem 3.10.3.
This paper was written in partial fulfillment of the requirements for the bachelor’s degree in mathematics at Yale. I thank Patrick Devlin, Miki Havlíčková, Kati Hubley, Matthew King, Andrew Neitzke, and Carol Rutschman for inspiring me to pursue and complete this degree.
2. Racks
2.1. Racks
In this subsection, we define racks and discuss several major examples.
2.1.1. Notation
Given a set $X$, we denote the permutation group of $X$ by $S_{X}$. In the case that $X=\{1,2,\dots,n\}$, we denote the symmetric group on $n$ letters by $S_{n}$. We also denote the composition of functions $\varphi:X\to Y$ and $\psi:Y\to Z$ by $\psi\varphi$.
[…]
$x\triangleright y=s_{y}(x).$
In this paper, we adopt the convention using permutations due to its convenience for abstract proofs and exhaustive search algorithms.
2.1.2. Racks and quandles
[…]
Definition 2.1.
Let $X$ be a set, let $s:X\to S_{X}$ be a map, and write $s_{x}:=s(x)$ for all elements $x\in X$. We call the pair $(X,s)$ a rack if
$s_{x}s_{y}=s_{s_{x}(y)}s_{x}$
for all $x,y\in X$, in which case we call $s$ a rack structure on $X$. If in addition $s_{x}(x)=x$ for all $x\in X$, then we say that $(X,s)$ is a quandle. We also say that $|X|$ is the order of $(X,s)$. Finally, if $Y\subseteq X$ and $s_{y}^{\pm 1}(z)\in Y$ for all $y,z\in Y$, then we say that $(Y,s|_{Y})$ is a subrack of $(X,s)$.
Example 2.2.
[quandlebook]*Example 99 Let $X$ be a set, and fix $\sigma\in S_{X}$. Define $s:X\rightarrow S_{X}$ by $x\mapsto\sigma$, so that $s_{x}(y)=\sigma(y)$ for all $x,y\in X$. Then $(X,\sigma)_{\mathrm{perm}}:=(X,s)$ is a rack called a permutation rack or constant action rack. (Our notation $(X,\sigma)_{\mathrm{perm}}$ is nonstandard, unlike the notation in the next three examples.) Note that $(X,\sigma)_{\mathrm{perm}}$ is a quandle if and only if $\sigma=\operatorname{id}_{X}$, in which case we call $(X,\operatorname{id}_{X})_{\mathrm{perm}}$ a trivial quandle.
Example 2.3.
[quandlebook]*Example 54 Let $A$ be an abelian additive group. Define $s:A\to S_{A}$ by $b\mapsto s_{b}$ with $s_{b}(a):=2b-a$ for all elements $a,b\in A$. Then $T(A):=(A,s)$ is a quandle called a Takasaki kei. Takasaki kei are the earliest examples of racks in the literature; Takasaki [takasaki] introduced them in 1943 to study symmetric spaces.
Definition 2.4.
Given two racks $R:=(X,s)$ and $(Y,t)$, we say that a map $\varphi:X\to Y$ is a rack homomorphism if
$\varphi s_{x}=t_{\varphi(x)}\varphi$
for all $x\in X$. A rack isomorphism is a bijective rack homomorphism. Rack endomorphisms and automorphisms are defined in the obvious ways, and we denote the automorphism group of $R$ by $\operatorname{Aut}R$. Finally, the inner automorphism group of $R$ is the subgroup $\operatorname{Inn}R:=\langle s_{x}\mid x\in X\rangle$ of $\operatorname{Aut}R$.
Example 2.5.
[book]*Example 2.13 Let $X$ be a union of conjugacy classes in a group $G$, and define $c^{G}:X\rightarrow S_{X}$ by sending $x$ to the conjugation map $c^{G}_{x}:=[y\mapsto xyx^{-1}]$. Then $\operatorname{Conj}X:=(X,c^{G})$ is a quandle called a conjugation quandle or conjugacy quandle.
All group homomorphisms $\varphi:G\to H$ are rack homomorphisms from $\operatorname{Conj}G$ to $\operatorname{Conj}H$ since
$\varphi c^{G}_{x}(y)=\varphi(xyx^{-1})=\varphi(x)\varphi(y)\varphi(x)^{-1}=c^{H}_{\varphi(x)}\varphi(y)$
for all $x,y\in G$. Moreover, if $G$ is abelian, then for any subset $X\subseteq G$, the identity map $\operatorname{id}_{X}$ is a rack isomorphism from $\operatorname{Conj}X$ to the trivial quandle $(X,\operatorname{id}_{X})_{\mathrm{perm}}$.
Example 2.6.
For all racks $(X,s)$, the rack structure $s:X\to S_{X}$ is a rack homomorphism from $(X,s)$ to $\operatorname{Conj}S_{X}$ because
$ss_{x}(y)=s_{s_{x}(y)}=s_{s_{x}(y)}s_{x}s_{x}^{-1}=s_{x}s_{y}s_{x}^{-1}=c^{S_{X}}_{s_{x}}(s_{y})=c^{S_{X}}_{s(x)}s(y)$
for all $x,y\in X$.
Example 2.7.
[quandlebook]*Example 66 Let $n\geq 3$ be an integer. We call the Takasaki kei $R_{n}:=T(\mathbb{Z}/n\mathbb{Z})$ a dihedral quandle due to the following observation: if $\Sigma=\{s,rs,\dots,r^{n-1}s\}$ is the set of all reflections in the dihedral group $D_{n}=\langle r,s\mid r^{n}=s^{2}=1,\,srs=r^{-1}\rangle$ of order $2n$, then the map $\varphi:\mathbb{Z}/n\mathbb{Z}\to\Sigma$ defined by $k\mapsto r^{k}s$ is a rack isomorphism from $R_{n}$ to $\operatorname{Conj}\Sigma$.
2.1.3. Dual racks
Every rack has a canonical dual rack, which we define as follows.
Definition 2.8.
[dual]*Definition 2.6 Given a rack $R=(X,s)$, define $s^{\prime}:X\to S_{X}$ by $x\mapsto s_{x}^{-1}$. Then $R^{\operatorname{op}}:=(X,s^{\prime})$ is a rack called the dual rack of $R$. (Our notation $R^{\operatorname{op}}$ is nonstandard.)
Example 2.9.
For all permutation racks $(X,\sigma)_{\mathrm{perm}}$, we have $(X,\sigma)_{\mathrm{perm}}^{\operatorname{op}}=(X,\sigma^{-1})_{\mathrm{perm}}$.
Proposition 2.10.
For all racks $R=(X,s)$, we have $\operatorname{Aut}R=\operatorname{Aut}R^{\operatorname{op}}$ and $\operatorname{Inn}R=\operatorname{Inn}R^{\operatorname{op}}$. In particular, $s_{x}\in\operatorname{Aut}R^{\operatorname{op}}$ and $s_{x}^{-1}\in\operatorname{Aut}R$ for all $x\in X$.
Proof.
For all $\varphi\in\operatorname{Aut}R$ and $x\in X$, we have $\varphi s_{x}=s_{\varphi(x)}\varphi$. Since $\varphi$ and $s_{x}$ are bijections, we can take inverses of both sides to get
$s_{x}^{-1}\varphi^{-1}=\varphi^{-1}s_{\varphi(x)}^{-1}.$
Equivalently, $\varphi s_{x}^{-1}=s_{\varphi(x)}^{-1}\varphi$, so $\varphi\in\operatorname{Aut}R^{\operatorname{op}}$. Thus, $\operatorname{Aut}R\subseteq\operatorname{Aut}R^{\operatorname{op}}$; a dual argument shows the reverse containment and, hence, the final claim. The equality $\operatorname{Inn}R=\operatorname{Inn}R^{\operatorname{op}}$ is clear. ∎
2.1.4. Medial racks
Let $\mathsf{Rack}$ denote the category of racks and rack homomorphisms, and let $\mathsf{Qnd}$ be the full subcategory of $\mathsf{Rack}$ whose objects are quandles. From the perspective of universal algebra, racks are an algebraic theory with two binary operations $s_{-}(-)$ and $s^{-1}_{-}(-)$, and $\mathsf{Rack}$ is the category of models of this algebraic theory in $\mathsf{Set}$, the category of sets and set maps.
Thus, $\mathsf{Rack}$ is complete and cocomplete (see [alg-theories]*Theorem 3.4.5), as we will need later. In particular, the Cartesian product of racks has a natural rack structure, so we can define the following.
Definition 2.11.
[rack-roll]*Section 3 A rack $(X,s)$ is called medial or abelian if the map $X\times X\to X$ defined by $(x,y)\mapsto s_{y}(x)$ is a rack homomorphism. Equivalently, the subgroup $\langle s_{x}s_{y}^{-1}\mid x,y\in X\rangle$ of $\operatorname{Inn}X$ is abelian; see, for example, [medial-quandles]*Proposition 2.1. This subgroup is called the transvection group or displacement group of $(X,s)$. Equivalently, for all $x,y,z\in X$,
(1) $s_{s_{x}(z)}s_{y}=s_{s_{x}(y)}s_{z}.$
Let $\mathsf{Rack}_{\mathrm{med}}$ be the full subcategory of $\mathsf{Rack}$ whose objects are medial.
Note that mediality is not synonymous with the much rarer condition that $(X,s)$ is commutative, which requires that $s_{x}(y)=s_{y}(x)$ for all $x,y\in X$.
Example 2.12.
All permutation racks $(X,\sigma)_{\mathrm{perm}}$ are medial; since $s_{x}=\sigma$ for all $x\in X$, the transvection group of $(X,\sigma)_{\mathrm{perm}}$ is trivial.
Example 2.13.
All Takasaki kei $T(A)$ are medial because
$s_{s_{x}(z)}s_{y}(a)=4x-2z-2y+a=s_{s_{x}(y)}s_{z}(a)$
for all $x,y,z,a\in A$; that is, equation (1) always holds.
Example 2.14.
Up to isomorphism, there is exactly one nonmedial rack of order $4$ or lower. This rack, which is listed in the penultimate row of Table LABEL:tab:tab4 in Appendix B, is defined as follows. Let $X:=\{1,2,3,4\}$. In cycle notation, define $s:X\to S_{4}$ by $i\mapsto s_{i}$ with
$s_{1}:=\operatorname{id}_{X},\quad s_{2}:=(34),\quad s_{3}:=(24),\quad\text{and }s_{4}:=(23).$
Then $(X,s)$ is a nonmedial quandle because, for example,
$s_{s_{1}(3)}s_{2}=(24)(34)\neq(34)(24)=s_{s_{1}(2)}s_{3}.$
That is, equation (1) does not always hold in $(X,s)$.
Since the transvection groups of a rack $R$ and its dual $R^{\operatorname{op}}$ are equal, we have the following.
Lemma 2.15.
A rack $R$ is medial if and only if the dual rack $R^{\operatorname{op}}$ is medial.
2.2. The canonical automorphism of a rack
In this subsection, we discuss a collection $\Theta$ of canonical rack automorphisms $\theta_{R}\in\operatorname{Aut}R$ that generates the center of the category $\mathsf{Rack}$.
2.2.1. Definition of $\theta_{R}$
Given a rack $R=(X,s)$, define $\theta_{R}:X\to X$ by ${x\mapsto s_{x}(x)}$; see [center]*Proposition 2.5.
Evidently, $R$ is a quandle if and only if $\theta_{R}=\operatorname{id}_{X}$. Thus, we can loosely think of $\theta_{R}$ as measuring the failure of $R$ to be a quandle. Whenever there is no ambiguity, we will suppress the subscript and only write $\theta:=\theta_{R}$.
2.2.2. Categorical centers
Recall that the center of a category $\mathcal{C}$ is the commutative monoid $Z(\mathcal{C})$ of natural endomorphisms of the identity functor $\mathbf{1}_{\mathcal{C}}$. Concretely, $\eta\in Z(\mathcal{C})$ if and only if, for all objects $R,S$ and morphisms $\varphi:R\to S$ in $\mathcal{C}$, the component $\eta_{R}$ is an endomorphism of $R$, and $\eta_{S}\varphi=\varphi\eta_{R}$.
For example, if $A\text{-}\mathsf{mod}$ denotes the category of modules over a ring $A$, then the categorical center $Z(A\text{-}\mathsf{mod})$ is isomorphic to the ring-theoretic center $Z(A)$ of $A$.
2.2.3. Properties of $\theta$
Let $\Theta$ denote the collection of canonical automorphisms $\theta_{R}$ for all racks $R$. Szymik [center] proved the claims in the following proposition using the binary operation $\triangleright$. We offer a new proof of claim 3 and rewrite Szymik’s proofs of the other claims in terms of permutations.
Proposition 2.16.
For all racks $R=(X,s)$ and all integers $k\in\mathbb{Z}$, we have the following:
  1. (A1)
    $\theta:X\to X$ is a bijection with inverse $\theta^{-1}$ defined by $x\mapsto s_{x}^{-1}(x)$.
  2. (A2)
    $\theta^{\pm 1}_{S}\varphi=\varphi\theta^{\pm 1}_{R}$ for all racks $S=(Y,t)$ and rack homomorphisms $\varphi\in\operatorname{Hom}_{\mathsf{Rack}}(R,S)$.
  3. (A3)
    $s_{\theta^{k}(x)}=s_{x}$ for all $x\in X$.
  4. (A4)
    $\Theta^{k}\in Z(\mathsf{Rack})$.
Proof.
Define $\theta^{-1}$ as in claim 1, and fix $x\in X$. We deduce from Proposition 2.10 that
$\theta^{-1}\theta(x)=\theta^{-1}s_{x}(x)=s^{-1}_{s_{x}(x)}s_{x}(x)=s_{x}s^{-1}_{x}(x)=x,$
as desired. Dually, $\theta\theta^{-1}(x)=x$, which proves claim 1. To prove claim 2, observe that
$\theta_{S}\varphi(x)=t_{\varphi(x)}\varphi(x)=\varphi s_{x}(x)=\varphi\theta_{R}(x).$
Since $\theta_{S}$ and $\theta_{R}$ are bijections, we obtain claim 2.
To prove claim 3, recall from Example 2.6 that the rack structure $s:X\to S_{X}$ is a rack homomorphism from $R$ to $\operatorname{Conj}S_{X}$. By claim 2 and the fact that $\operatorname{Conj}S_{X}$ is a quandle,
$s_{\theta^{k}_{R}(x)}=s\theta^{k}_{R}(x)=\theta^{k}_{\operatorname{Conj}S_{X}}s(x)=s(x)=s_{x}$
for all integers $k\in\mathbb{Z}$, which proves claim 3. Now, claims 2 and 3 yield
$s_{\theta(x)}\theta=\theta s_{\theta(x)}=\theta s_{x},$
so $\theta$ is a rack endomorphism. Combined with claims 1 and 2, this proves claim 4. ∎
Hence, $\Theta$ generates a cyclic subgroup of $Z(\mathsf{Rack})$. In fact, this subgroup is $Z(\mathsf{Rack})$; see [center]*Theorem 5.4.
The following identities will be useful later on.
Proposition 2.17.
Let $R=(X,s)$ be a rack. Then the following hold for all $k\in\mathbb{Z}$ and $x\in X$:
  1. (B1)
    $\theta^{k}s_{x}=s_{x}\theta^{k}$.
  2. (B2)
    $\theta^{k}(x)=s^{k}_{x}(x)$.
Proof.
Claim 1 is immediate from the inclusions $s_{x}\in\operatorname{Aut}R$ and $\Theta^{\pm 1}\in Z(\mathsf{Rack})$.
To prove claim 2, we induct on $k$. The base case $k=0$ is trivial. For $k>0$, we have
$\theta^{k}(x)=\theta^{k-1}\theta(x)=\theta^{k-1}s_{x}(x)=s_{x}\theta^{k-1}(x)=s_{x}s^{k-1}_{x}(x)=s^{k}_{x}(x)$
as desired; the second equality uses the definition of $\theta$, while the third equality follows from claim 1. Thanks to part 1 of Proposition 2.16, a similar argument proves claim 2 for $k<0$. ∎
2.3. Racks as an algebraic theory
Recall that racks are an algebraic theory with two binary operations $s_{-}(-)$ and $s^{-1}_{-}(-)$. In this subsection, we discuss universal-algebraic free objects in $\mathsf{Rack}$ and rack structures on certain hom-sets of racks. While we provide the relevant results here, we also direct the reader to [alg-theories] for a reference on the underlying universal algebra.
2.3.1. Examples of quotient racks
Recall that $\mathsf{Rack}$ is the category of models of the algebraic theory of racks in $\mathsf{Set}$, so $\mathsf{Rack}$ is complete and cocomplete. Thus, we can consider quotients of racks by congruences; see [alg-theories]*Lemma 3.5.1 and cf. [dual]*Proposition 3.6. This allows us to state the following definitions, which are actually left adjoints of the inclusion functors from $\mathsf{Qnd}$ and $\mathsf{Rack}_{\mathrm{med}}$ into $\mathsf{Rack}$.
Definition 2.18.
[fenn]*Section 2, [joyce]*Section 10 Given a rack $R=(X,s)$, the associated quandle of $R$ is the quandle $R_{\mathrm{qnd}}:=R/{\sim}$, where $\sim$ is the smallest congruence on $X$ such that $s_{x}(x)=x$ for all $x\in X$. Similarly, the medialization or abelianization of $R$ is the medial rack $R_{\mathrm{med}}:=R/{\sim}$, where $\sim$ is the smallest congruence on $X$ such that
(2) ${s_{s_{x}(z)}s_{y}(a)=s_{s_{x}(y)}s_{z}(a)}$
for all $x,y,z,a\in X$.
2.3.2. Free racks
In analogy to free groups or free modules, the free rack $\operatorname{Free}_{\mathsf{Rack}}(X)$ on a set $X$ exists and is uniquely (up to isomorphism) characterized by a universal property; see, for example, [alg-theories]*Corollary 3.7.8. Namely, there exists a set map $i:X\to\operatorname{Free}_{\mathsf{Rack}}(X)$ such that for all racks $R=(Y,t)$ and all set maps $\varphi_{0}:X\to Y$, there exists a unique rack homomorphism ${\varphi\in\operatorname{Hom}_{\mathsf{Rack}}(\operatorname{Free}_{\mathsf{Rack}}(X),R)}$ such that $\varphi i=\varphi_{0}$. That is, the following diagram commutes:
${X}$${{\operatorname{Free}_{\mathsf{Rack}}(X)}}$${Y}$$\scriptstyle{i}$$\scriptstyle{\varphi_{0}}$$\scriptstyle{\varphi}$
For an explicit set-theoretic construction of free racks, see [free-z]*Proposition 1.3. We can also define the free quandle on $X$ to be $(\operatorname{Free}_{\mathsf{Rack}}(X))_{\mathrm{qnd}}$; cf. [fenn]*Section 7.1.
Example 2.19.
[free-z]*Example 1.6 The free rack $F$ on one generator is canonically isomorphic to the permutation rack $(\mathbb{Z},\sigma)_{\mathrm{perm}}$, where $\sigma(k)=k+1$ for all $k\in\mathbb{Z}$. Under this identification, $\theta_{F}=\sigma$. By contrast, the free quandle on one generator is the trivial quandle with one element.
For fun, we encourage the reader to prove that $F\cong(\mathbb{Z},\sigma)_{\mathrm{perm}}$. To do this, appeal to the universal property of $\operatorname{Free}_{\mathsf{Rack}}(\{x\})$ with $i(x):=0$ and $\varphi(k):=\theta_{R}^{k}\varphi_{0}(x)$.
2.3.3. Commutative algebraic theories
Medial racks are the largest commutative subtheory of racks. Recall that an algebraic theory $\mathcal{T}$ is called commutative if, for all $\mathcal{T}$-models $A$, every $n$-ary operation $\alpha:\mathcal{T}^{n}\to\mathcal{T}$ defines a homomorphism $A^{n}\to A$; see [alg-theories]*Theorem 3.10.3.
Let $\mathcal{T}_{\operatorname{ab}}$ be the largest commutative subtheory of an algebraic theory $\mathcal{T}$. If $\mathcal{C}$ is the category of $\mathcal{T}$-models in $\mathsf{Set}$, let $\mathcal{C}_{\operatorname{ab}}$ be the full subcategory of $\mathcal{C}$ whose objects are $\mathcal{T}_{\operatorname{ab}}$-models. The following strengthens part of [alg-theories]*Theorem 3.10.3; we could not find a reference for this precise statement.
Proposition 2.20.
For all $\mathcal{T}$-models $X$ in $\mathcal{C}$ and $\mathcal{T}_{\operatorname{ab}}$-models $Y$ in $\mathcal{C}_{\operatorname{ab}}$, the set $H:=\operatorname{Hom}_{\mathcal{C}}(X,Y)$ has a canonical $\mathcal{T}_{\operatorname{ab}}$-model structure defined by
$\alpha(f_{1},\dots,f_{n})(x):=\alpha(f_{1}(x),\dots,f_{n}(x))$
for all $n$-ary operations $\alpha:\mathcal{T}^{n}\to\mathcal{T}$, $\mathcal{T}$-model homomorphisms $f_{1},\dots,f_{n}\in H$, and elements $x\in X$.
Proof.
We refer the reader to the proof of the statement in [alg-theories]*Theorem 3.10.3 that (1) implies (3). This proof uses the commutativity of $n$-ary operations in $Y$ but not in $X$, so the implication still holds with the weakened assumption that $X$ is not necessarily a $\mathcal{T}_{\operatorname{ab}}$-model. ∎
Example 2.21.
Let $\mathcal{T}$ be the algebraic theory of groups. Then $\mathcal{T}_{\operatorname{ab}}$ is the algebraic theory of abelian groups because a group $A$ is abelian if and only if its group multiplication $\cdot:A\times A\to A$ and the inversion operation :${}^{-1}:A\to A$ are group homomorphisms.
Proposition 2.20 generalizes the well-known fact that the set of group homomorphisms from a possibly nonabelian group $G$ to an abelian group $A$ has an abelian group structure given by
$(\varphi+\psi)(g):=\varphi(g)+\psi(g)$
for all $g\in G$.
2.3.4. Mediality and Hom racks
If we take $\mathcal{T}$ to be the algebraic theory of racks (resp. quandles), then $\mathcal{T}_{\operatorname{ab}}$ is the algebraic theory of medial racks (resp. medial quandles). That is, if $\mathcal{C}=\mathsf{Rack}$, then $\mathcal{C}_{\operatorname{ab}}=\mathsf{Rack}_{\mathrm{med}}$; this follows directly from Definition 2.11 and Lemma 2.15.
Hence, Proposition 2.20 recovers the following result of Grøsfjeld [rack-roll]*Proposition 3.3 in 2021, which in turn generalized a result of Crans and Nelson [Hom]*Theorem 3 in 2014.
Corollary 2.22.
Let $R=(X,s)$ be a rack, and let $M=(Y,t)$ be a medial rack. Then the hom-set $H:=\operatorname{Hom}_{\mathsf{Rack}}(R,M)$ has a canonical medial rack structure $\tilde{t}:H\to S_{H}$ defined by
$\tilde{t}_{g}(f)(x):=t_{g(x)}f(x)$
for all $f,g\in H$ and $x\in X$. Moreover, if $R$ is a quandle or $M$ is a quandle, then $H$ is also a quandle.
Proof.
The only statement that does not directly follow from Proposition 2.20 is the final sentence, which is straightforwardly verified. ∎
3. Generalized Legendrian racks
3.1. GL-racks
In this subsection, we define generalized Legendrian racks (also called GL-racks), which Karmakar et al. [karmakar] and Kimura [bi] introduced independently in 2023. Once again, we express our definition in terms of permutations.
Definition 3.1.
Given a rack $R=(X,s)$, a generalized Legendrian structure or GL-structure on $R$ is a rack automorphism $\mathtt{u}\in\operatorname{Aut}R$ such that $\mathtt{u}s_{x}=s_{x}\mathtt{u}$ for all $x\in X$. We call the pair $(R,\mathtt{u})$ a generalized Legendrian rack or GL-rack. If in addition $R$ is a quandle or a medial rack, then we also call $(R,\mathtt{u})$ a GL-quandle or a medial GL-rack, respectively.
Definition 3.2.
A GL-rack homomorphism between two GL-racks $(R_{1},\mathtt{u}_{1})$ and $(R_{2},\mathtt{u}_{2})$ is a rack homomorphism $\varphi\in\operatorname{Hom}_{\mathsf{Rack}}(R_{1},R_{2})$ satisfying $\varphi\mathtt{u}_{1}=\mathtt{u}_{2}\varphi$. We denote the category of GL-racks and their homomorphisms by $\mathsf{GLR}$. Finally, let $\mathsf{GLQ}$, $\mathsf{GLR}_{\mathrm{med}}$, and $\mathsf{GLQ}_{\mathrm{med}}$ be the full subcategories of $\mathsf{GLR}$ whose objects are GL-quandles, medial GL-racks, and medial GL-quandles, respectively.
Remark 3.3.
Virtual racks are algebraic structures that can distinguish framed links in certain lens spaces and framed virtual links in thickened surfaces; see, for example, [virtual]*Section 3.2.
By Definition 3.1, GL-racks are precisely virtual racks in which all inner automorphisms $s_{x}$ are endomorphisms of virtual racks. Equivalently, a GL-rack $(X,s,\mathtt{u})$ is a virtual rack in which the operator group of $(X,s)$ identifies $x$ with $\mathtt{u}(x)$ for all $x\in X$; see [fenn]*Section 1.1.
[…]
Example 3.4.
For all racks $R=(X,s)$, the identity map $\operatorname{id}_{X}$ is a GL-structure on $R$. Consequently, every rack can be equipped with at least one GL-structure.
Example 3.5.
[bi]*Example 3.7 Given a permutation rack $P=(X,\sigma)_{\mathrm{perm}}$, a GL-structure on $P$ is precisely a permutation $\mathtt{u}\in S_{X}$ such that $\mathtt{u}\sigma=\sigma\mathtt{u}$. Given such a $\mathtt{u}$, we say that $(P,\mathtt{u})$ is a permutation GL-rack or constant action GL-rack, and we denote it by $(X,\sigma,\mathtt{u})_{\mathrm{perm}}$.
Example 3.6.
[bi]*Example 3.6 Let $G$ be a group, let $z\in Z(G)$ be a central element of $G$, and define $f:G\rightarrow G$ by $g\mapsto zg$. Then $(\operatorname{Conj}G,f)$ is a GL-quandle.
Example 3.7.
Let $n\geq 4$ be a multiple of $4$. Define four affine transformations ${f_{a,u}:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}}$ by $k\mapsto a+uk$ with $a\in\{0,n/2\}$ and $u\in\{1,1+n/2\}$. Each of these transformations is a GL-structure on the dihedral quandle $R_{n}$, and the translation defined by $k\mapsto k+1$ is a GL-rack isomorphism from $(R_{n},f_{0,1+n/2})$ to $(R_{n},f_{n/2,1+n/2})$. We later show that these are all the possible GL-structures on $R_{n}$; see Proposition 4.11.
Lemma 3.8.
If $\varphi$ is a GL-rack homomorphism from $(R_{1},\mathtt{u}_{1})$ to $(R_{2},\mathtt{u}_{2})$, then $\varphi\mathtt{u}_{1}^{-1}=\mathtt{u}_{2}^{-1}\varphi$.
Proof.
By hypothesis, $\varphi=\varphi\mathtt{u}_{1}\mathtt{u}_{1}^{-1}=\mathtt{u}_{2}\varphi\mathtt{u}_{1}^{-1}$. Applying $\mathtt{u}_{2}^{-1}$ on the left proves the claim. ∎
3.1.1. Bi-Legendrian racks
Next, we reproduce the definition of GL-racks given in the literature; see [karmakar]*Definition 3.1 or [bi]*Definition 3.2. Following Kimura [bi], we temporarily use the term bi-Legendrian racks to distinguish them from GL-racks in the sense of Definition 3.1. We also define Legendrian racks, which Ceniceros et al. [ceniceros] introduced in 2021.
Definition 3.9.
[karmakar]*Definition 3.1 Given a rack $R=(X,s)$, a bi-Legendrian structure on $R$ is a pair $(\mathtt{u},\mathtt{d})$ of maps $\mathtt{u},\mathtt{d}:X\to X$ satisfying the following axioms for all elements $x\in X$:
  1. (L1)
    $\mathtt{u}\mathtt{d}s_{x}(x)=x=\mathtt{d}\mathtt{u}s_{x}(x)$.
  2. (L2)
    $\mathtt{u}s_{x}=s_{x}\mathtt{u}$, and $\mathtt{d}s_{x}=s_{x}\mathtt{d}$.
  3. (L3)
    $s_{\mathtt{u}(x)}=s_{x}=s_{\mathtt{d}(x)}$.
We call the triple $(R,\mathtt{u},\mathtt{d})$ a bi-Legendrian rack; if $\mathtt{u}=\mathtt{d}$, then we also say that $(R,\mathtt{u},\mathtt{d})$ is a Legendrian rack. A bi-Legendrian rack homomorphism from $(R_{1},\mathtt{u}_{1},\mathtt{d}_{1})$ to $(R_{2},\mathtt{u}_{2},\mathtt{d}_{2})$ is a rack homomorphism $\varphi\in\operatorname{Hom}_{\mathsf{Rack}}(R_{1},R_{2})$ such that $\varphi\mathtt{u}_{1}=\mathtt{u}_{2}\varphi$ and $\varphi\mathtt{d}_{1}=\mathtt{d}_{2}\varphi$.
Remark 3.10.
Axiom 1 shows that, for all bi-Legendrian racks $(R,\mathtt{u},\mathtt{d})$, the underlying rack $R$ is a quandle if and only if $\mathtt{d}=\mathtt{u}^{-1}$.
3.2. Equivalence of definitions
In this subsection, we show the equivalence of Definitions 3.1 and 3.9. Specifically, we show that the category $\mathsf{BLR}$ of bi-Legendrian racks is isomorphic to $\mathsf{GLR}$.
Proposition 3.11.
If $(\mathtt{u},\mathtt{d})$ is a bi-Legendrian structure on a rack $R=(X,s)$, then $\mathtt{d}=\theta^{-1}\mathtt{u}^{-1}$, and $(R,\mathtt{u})$ is a GL-rack. In particular, there is a forgetful functor $\operatorname{For}:\mathsf{BLR}\to\mathsf{GLR}$ sending $(R,\mathtt{u},\mathtt{d})$ to $(R,\mathtt{u})$. Moreover, $\operatorname{For}$ is fully faithful.
Proof.
Axiom 1 states that
$\mathtt{u}\mathtt{d}=\mathtt{d}\mathtt{u}=\theta^{-1},$
which is bijective, so $\mathtt{u}$ is bijective. It follows that $\mathtt{d}=\theta^{-1}\mathtt{u}^{-1}$, as desired. Moreover, axioms 2 and 3 imply that $\mathtt{u}$ is a rack endomorphism that commutes with $s_{x}$ for all $x\in X$, so $(R,\mathtt{u})$ is a GL-rack. Thus, we have a forgetful functor $\operatorname{For}:\mathsf{BLR}\to\mathsf{GLR}$, which is clearly faithful.
Now, let $(R_{1},\mathtt{u}_{1},\mathtt{d}_{1})$ and $(R_{2},\mathtt{u}_{2},\mathtt{d}_{2})$ be bi-Legendrian racks. To show that $\operatorname{For}$ is full, we need to show that all GL-rack homomorphisms from $(R_{1},\mathtt{u}_{1})$ to $(R_{2},\mathtt{u}_{2})$ commute with the $\mathtt{d}_{i}$’s. But this follows from the formula $\mathtt{d}_{i}=\theta_{R_{i}}^{-1}\mathtt{u}_{i}^{-1}$, the inclusion $\Theta^{-1}\in Z(\mathsf{Rack})$, and Lemma 3.8. ∎
Proposition 3.12.
The forgetful functor $\operatorname{For}:\mathsf{BLR}\to\mathsf{GLR}$ is an isomorphism of categories. Hence, the data and axioms of GL-racks are equivalent to those of bi-Legendrian racks.
Proof.
Since $\operatorname{For}$ is fully faithful, it will suffice to show that $\operatorname{For}$ is bijective on objects. Injectivity follows from the first claim of Proposition 3.11.
To show surjectivity, let $(R,\mathtt{u})$ be a GL-rack with $R=(X,s)$, and define $\mathtt{d}:=\theta^{-1}\mathtt{u}^{-1}$ as in Proposition 3.11. It will suffice to show that $(R,\mathtt{u},\mathtt{d})$ is a bi-Legendrian rack. To that end, fix $x\in X$. Since $\Theta^{-1}\in Z(\mathsf{Rack})$,
$\mathtt{ud}s_{x}(x)=\mathtt{u}\theta^{-1}\mathtt{u}^{-1}\theta(x)=\mathtt{u}\mathtt{u}^{-1}\theta^{-1}\theta(x)=x$
and, similarly, $\mathtt{du}s_{x}(x)=x$. This verifies axiom 1. Lemma 3.8 and the inclusion $\Theta^{-1}\in Z(\mathsf{Rack})$ imply that $\mathtt{d}s_{x}=s_{x}\mathtt{d}$, which verifies axiom 2. Since $\mathtt{u}$ is a GL-structure,
$s_{\mathtt{u}(x)}\mathtt{u}=\mathtt{u}s_{x}=s_{x}\mathtt{u}.$
From the bijectivity of $\mathtt{u}$, we obtain $s_{\mathtt{u}(x)}=s_{x}$. By construction, $\mathtt{d}\in\operatorname{Aut}R$, so we similarly obtain $s_{\mathtt{d}(x)}=s_{x}$. This verifies axiom 3. Hence, $(R,\mathtt{u},\mathtt{d})$ is a bi-Legendrian rack, as desired. ∎
In light of these results, we will henceforth call bi-Legendrian racks GL-racks except when specifically citing the axioms in Definition 3.9 or denoting GL-racks as quadruples with $\mathtt{d}$. The previous two propositions make computing and classifying GL-racks significantly easier; see Section 4 and Appendix A.2. As a bonus, they also yield the following converse of [ceniceros]*Remark 2.
Corollary 3.13.
A GL-rack $(R,\mathtt{u})$ is a Legendrian rack if and only if $\theta_{R}=\mathtt{u}^{-2}$. In this case, $R$ is a quandle if and only if $\mathtt{u}$ is an involution.
3.3. GL-racks as an algebraic theory
In this subsection, we adapt the discussion in Section 2.3 to GL-racks, viewed as an algebraic theory with two binary operations $s_{-}(-)$ and $s^{-1}_{-}(-)$ and two unary operations $\mathtt{u}$ and $\mathtt{u}^{-1}$.
3.3.1. Examples of quotient GL-racks
First, we adapt the definitions of Section 2.3 from $\mathsf{Rack}$ to $\mathsf{GLR}$.
Definition 3.14.
Given a GL-rack $R=(X,s,\mathtt{u})$, the associated GL-quandle of $R$ is the GL-quandle $R_{\mathrm{qnd}}:=R/{\sim}$, where $\sim$ is the smallest congruence on $R$ such that $s_{x}(x)\sim x$ for all $x\in X$. Similarly, the medialization or abelianization of $R$ is the medial GL-rack $R_{\mathrm{med}}:=R/{\sim}$, where $\sim$ is the smallest congruence on $R$ such that equation (2) holds for all $x,y,z,a\in X$.
3.3.2. Free GL-racks
As with racks, the free GL-rack $\operatorname{Free}_{\mathsf{GLR}}(X)$ on a set $X$ exists and is uniquely (up to isomorphism) characterized by a universal property; see [karmakar]*Proposition 4.2 for a detailed derivation. Namely, there exists a set map $i:X\to\operatorname{Free}_{\mathsf{GLR}}(X)$ such that for all GL-racks $R=(Y,t,\mathtt{u})$ and all set maps $\varphi_{0}:X\to Y$, there exists a unique GL-rack homomorphism $\varphi\in\operatorname{Hom}_{\mathsf{GLR}}(\operatorname{Free}_{\mathsf{GLR}}(X),R)$ such that $\varphi i=\varphi_{0}$. That is, the following diagram commutes:
(3) ${X}$${{\operatorname{Free}_{\mathsf{GLR}}(X)}}$${Y}$$\scriptstyle{i}$$\scriptstyle{\varphi_{0}}$$\scriptstyle{\varphi}$
Karmakar et al. [karmakar]*Section 4 gave an explicit set-theoretic construction of $\operatorname{Free}_{\mathsf{GLR}}(X)$. Here, we simplify their construction using Definition 3.1 and and Proposition 3.12.
Definition 3.15.
[karmakar]*Section 4 Let $X$ be a set. We define the free GL-rack on $X$ as follows. If $X=\emptyset$, let $\operatorname{Free}_{\mathsf{GLR}}(X)$ be the empty GL-rack. Else, let the universe of words generated by $X$ be the set $W(X)$ such that $X\subset W(X)$ and $s_{y}(x),\,s_{y}^{-1}(x),\,\mathtt{u}(x),\,\mathtt{u}^{-1}(x)\in W(X)$ for all $x,y\in W(X)$. Let $V(X)$ be the set of equivalence classes of elements of $W(X)$ modulo the congruence generated by the following relations for all $x,y,z\in W(X)$:
  • $s_{y}^{-1}s_{y}(x)\sim s_{y}s_{y}^{-1}(x)\sim x\sim\mathtt{u}\mathtt{u}^{-1}(x)\sim\mathtt{u}^{-1}\mathtt{u}(x)$.
  • $s_{z}s_{y}(x)\sim s_{s_{z}(y)}s_{z}(x)$.
  • $s_{\mathtt{u}(y)}\mathtt{u}(x)\sim\mathtt{u}s_{y}(x)\sim s_{y}\mathtt{u}(x)$.
Thus, we have a rack structure $s:V(X)\to S_{V(X)}$ on $V(X)$ and a GL-structure $\mathtt{u}\in S_{V(X)}$ on $(V(X),s)$. So, we define $\operatorname{Free}_{\mathsf{GLR}}(X)$ to be the GL-rack $(V(X),s,\mathtt{u})$.
Remark 3.16.
By way of Proposition 3.11, for all $x\in V(X)$, we can consider $\mathtt{d}(x):=\theta^{-1}\mathtt{u}^{-1}(x)$ as an element of $V(X)$.
3.3.3. Free objects in subtheories of GL-racks
Since $\mathsf{GLR}$ is the category of models of the algebraic theory of GL-racks in $\mathsf{Set}$, the free GL-rack $L=\operatorname{Free}_{\mathsf{GLR}}(X)$ on the one-element set $X=\{x\}$ is a strong generator or separator for $\mathsf{GLR}$; see, for example, [alg-theories]*Proposition 3.3.3.
Since GL-quandles, Legendrian racks, and Legendrian quandles are subtheories of the algebraic theory of GL-racks, we can similarly consider free objects that strongly generate the corresponding subcategories of models by taking quotients of $L$ by congruences.
Definition 3.17.
Given a set $Y$, let $R$ be the free GL-rack on $Y$. We define the free GL-quandle on $Y$ to be $R_{\mathrm{qnd}}$.
Using Corollary 3.13, we similarly define the free Legendrian rack on $Y$ to be the quotient of $R$ by the smallest congruence $\sim$ such that $s_{y}(y)\sim\mathtt{u}^{-2}(y)$ for all $y\in V(Y)$, and we define the free Legendrian quandle on $Y$ to be $(R/{\sim})_{\mathrm{qnd}}$.
By the above discussion, we have the following.
Proposition 3.18.
Let $L$ be the free GL-rack on one element. Then $L_{\mathrm{qnd}}$, $L/{\sim}$, and $(L/{\sim})_{\mathrm{qnd}}$ are strong generators of $\mathsf{GLQ}$, the category of Legendrian racks, and the category of Legendrian quandles, respectively. Moreover, as sets,
$L_{\mathrm{qnd}}=L/{\sim}=\{\mathtt{u}^{k}(x)\mid k\in\mathbb{Z}\},\quad\text{and }(L/{\sim})_{\mathrm{qnd}}=\{x,\mathtt{u}(x)\}.$
3.3.4. The free GL-rack on one generator
We prove an analogue of Example 2.19 for GL-racks.
Proposition 3.19.
The free GL-rack on one generator is canonically isomorphic to the permutation GL-rack $L:=(\mathbb{Z}^{2},\sigma,\mathtt{u}_{0})_{\mathrm{perm}}$, where $\sigma(m,n)=(m+1,n)$ and $\mathtt{u}_{0}(m,n)=(m,n+1)$ for all $m,n\in\mathbb{Z}$.
Proof.
Let $X=\{x\}$. We will show that $L$ satisfies the universal property of $\operatorname{Free}_{\mathsf{GLR}}(X)$ with $i:X\to\mathbb{Z}^{2}$ defined by $x\mapsto(0,0)$. To that end, let $(R,\mathtt{u})$ be a GL-rack with $R=(Y,t)$, and let $\varphi_{0}:X\to Y$ be a set map. Define $\varphi:\mathbb{Z}^{2}\to Y$ by
$(m,n)\mapsto\mathtt{u}^{n}\theta_{R}^{m}\varphi_{0}(x).$
First, we show that $\varphi$ is a GL-rack homomorphism from $L$ to $(R,\mathtt{u})$. To that end, fix elements $(k,\ell),(m,n)\in\mathbb{Z}^{2}$. Denote the underlying rack structure on $L$ by $s:\mathbb{Z}^{2}\to S_{\mathbb{Z}^{2}}$ with $s_{(m,n)}:=\sigma$ for all $(m,n)\in\mathbb{Z}^{2}$. By part 2 of Proposition 2.17,
$\varphi s_{(k,\ell)}(m,n)=\varphi\sigma(m,n)=\varphi(m+1,n)=\mathtt{u}^{n}\theta_{R}^{m+1}\varphi_{0}(x)=\mathtt{u}^{n}t^{m+1}_{\varphi_{0}(x)}\varphi_{0}(x).$
On the other hand, we apply bi-Legendrian rack axiom 3, part 3 of Proposition 2.16, and part 2 of Proposition 2.17 to compute
$t_{\varphi(k,\ell)}\varphi(m,n)=t_{\mathtt{u}^{\ell}\theta_{R}^{k}\varphi_{0}(x)}\mathtt{u}^{n}\theta_{R}^{m}\varphi_{0}(x)=t_{\varphi_{0}(x)}\mathtt{u}^{n}t^{m}_{\varphi_{0}(x)}\varphi_{0}(x)=\mathtt{u}^{n}t^{m+1}_{\varphi_{0}(x)}\varphi_{0}(x)=\varphi s_{(k,\ell)}(m,n),$
so $\varphi$ is a rack homomorphism. Moreover,
$\varphi\mathtt{u}_{0}(m,n)=\varphi(m,n+1)=\mathtt{u}^{n+1}\theta_{R}^{m}\varphi_{0}(x)=\mathtt{u}\varphi(m,n),$
so $\varphi$ is a GL-rack homomorphism.
Next, we show that diagram (3) commutes and that $\varphi$ is unique. Indeed, we have commutativity because
$\varphi i(x)=\varphi(0,0)=\mathtt{u}^{0}\theta^{0}_{R}\varphi_{0}(x)=\varphi_{0}(x).$
Now, let $\psi\in\operatorname{Hom}_{\mathsf{GLR}}(L,(R,\mathtt{u}))$ be any GL-rack homomorphism such that $\psi i(x)=\varphi_{0}(x)$, and fix $(m,n)\in\mathbb{Z}^{2}$. By part 2 of Proposition 2.17,
$\mathtt{u}^{-n}\theta_{R}^{-m}\psi(m,n)=\mathtt{u}^{-n}t^{-m}_{\psi(m,n)}\psi(m,n)=\psi\mathtt{u}_{0}^{-n}s^{-m}_{(m,n)}(m,n)=\psi(0,0)=\psi i(x)=\varphi_{0}(x)$
and, similarly,
$\mathtt{u}^{-n}\theta_{R}^{-m}\varphi(m,n)=\varphi_{0}(x).$
Since $\mathtt{u}$ and $\theta_{R}$ are bijections, we obtain $\psi(m,n)=\varphi(m,n)$. Since the element $(m,n)\in\mathbb{Z}^{2}$ was arbitrary, this shows that $\varphi$ is unique. Hence, $L$ satisfies the universal property of $\operatorname{Free}_{\mathsf{GLR}}(X)$. ∎
3.3.5. Mediality and Hom GL-racks
[…]
Theorem 3.20.
Let $R_{1}:=(X,s,\mathtt{u}_{1})$ and $R_{2}:=(Y,t,\mathtt{u}_{2})$ be GL-racks, and suppose that $R_{2}$ is medial. Then $H:=\operatorname{Hom}_{\mathsf{GLR}}(R_{1},R_{2})$ is a subrack of $\operatorname{Hom}_{\mathsf{Rack}}((X,s),(Y,t))$ equipped with its medial rack structure $\tilde{t}$ from Corollary 2.22, and $(H,\tilde{t}|_{H})$ has a canonical GL-structure $\mathtt{u}:H\rightarrow H$ defined by $f\mapsto\mathtt{u}_{2}f$. In particular, if $R_{1}$ or $R_{2}$ is a GL-quandle, then so is $(H,\tilde{t}|_{H},\mathtt{u})$.
Proof.
By Lemmas 2.15 and 3.8, the binary operations $s_{-}(-)$ and $s^{-1}_{-}(-)$ and the unary operations $\mathtt{u}$ and $\mathtt{u}^{-1}$ are all homomorphisms in the algebraic theory of medial GL-racks. Therefore, medial GL-racks are a commutative algebraic theory, so the claim follows directly from Proposition 2.20 and Corollary 2.22. ∎
4. Classification of GL-structures, automorphisms, and centers
In 2024, Karmakar et al. [karmakar]*Section 3 posed the following question: what are all the possible GL-structures on a given rack? In this section, we answer this question and classify various infinite families of GL-racks. As applications, we discuss automorphism groups of GL-racks and compute these groups for all GL-racks whose underlying racks are dihedral quandles. We also compute the centers of $\mathsf{GLR}$, $\mathsf{GLQ}$, and the categories of Legendrian racks and Legendrian quandles.
4.1. Characterizing GL-structures
Given a rack $R=(X,s)$, let $U_{R}$ be the set of GL-structures $\mathtt{u}:X\to X$ on $R$, and define an equivalence relation $\sim$ on $U_{R}$ by identifying $\mathtt{u}_{1}\sim\mathtt{u}_{2}$ if and only if $(R,\mathtt{u}_{1})\cong(R,\mathtt{u}_{2})$. The following theorem completely characterizes $U_{R}$ and $U_{R}/{\sim}$. Given a group $G$ and a subset or element $H$ of $G$, we will denote the centralizer of $H$ in $G$ by $C_{G}(H)$.
Theorem 4.1.
Given a rack $R=(X,s)$, define $C:=C_{\operatorname{Aut}R}(\operatorname{Inn}R)$. Then $U_{R}=C$, and $U_{R}$ is a normal subgroup of $\operatorname{Aut}R$. Furthermore, $\mathtt{u}_{1}\sim\mathtt{u}_{2}$ if and only if $\mathtt{u}_{1}$ and $\mathtt{u}_{2}$ are conjugate in $\operatorname{Aut}R$. In particular, if $\operatorname{Aut}R$ is abelian, then $U_{R}=U_{R}/{\sim}=\operatorname{Aut}R$.
Proof.
The claim that $U_{R}=C$ is a restatement of Definition 3.1. It is straightforward to verify that $\operatorname{Inn}R$ is a normal subgroup of $\operatorname{Aut}R$, so $C$ is normal in $\operatorname{Aut}R$.
On the other hand, given two GL-structures $\mathtt{u}_{1},\mathtt{u}_{2}\in U_{R}=C$, a map $\varphi:X\to X$ is a GL-rack isomorphism from $(R,\mathtt{u}_{1})$ to $(R,\mathtt{u}_{2})$ if and only if $\varphi\in\operatorname{Aut}R$ and $\varphi\mathtt{u}_{1}=\mathtt{u}_{2}\varphi$. In other words, $(R,\mathtt{u}_{1})\cong(R,\mathtt{u}_{2})$ if and only if $\mathtt{u}_{1}$ and $\mathtt{u}_{2}$ are conjugate in $\operatorname{Aut}R$, as claimed.
Finally, suppose that $\operatorname{Aut}R$ is abelian. Then $\operatorname{Aut}R=C=U_{R}$, and each element of $\operatorname{Aut}R$ constitutes its own conjugacy class. Hence, $U_{R}=U_{R}/{\sim}$. ∎
One may ask whether $U_{R}$ is conjugacy-closed in $\operatorname{Aut}R$, that is, whether $\mathtt{u}_{1}\sim\mathtt{u}_{2}$ in $U_{R}/{\sim}$ if and only if $\mathtt{u}_{1}$ and $\mathtt{u}_{2}$ are conjugate in $U_{R}$. We will give a negative answer later; see Remark 4.12.
Corollary 4.2.
For all racks $R$, we have $U_{R}=U_{R^{\operatorname{op}}}$ and $U_{R}/{\sim}=U_{R^{\operatorname{op}}}/{\sim}$.
Proof.
This follows immediately from Proposition 2.10 and Theorem 4.1. ∎
4.2. Classification of GL-racks
In this subsection, we use Theorem 4.1 to classify GL-structures on various infinite families of racks.
4.2.1. Permutation GL-racks
First, we classify GL-structures on permutation racks.
Proposition 4.3.
Let $X$ be a set, let $\sigma\in S_{X}$, and let $P$ be the permutation rack $(X,\sigma)_{\mathrm{perm}}$. Then $U_{P}=C_{S_{X}}(\sigma)=\operatorname{Aut}P$, and $U_{P}/{\sim}$ is the set of conjugacy classes of $C_{S_{X}}(\sigma)$.
Proof.
An automorphism of $P$ is precisely a permutation $\varphi\in S_{X}$ such that $\varphi\sigma=\sigma\varphi$. Therefore, $\operatorname{Aut}P=C_{S_{X}}(\sigma)$. On the other hand, $\operatorname{Inn}P=\langle\sigma\rangle$, so Theorem 4.1 states that
$U_{P}=C_{C_{S_{X}}(\sigma)}(\langle\sigma\rangle)=C_{C_{S_{X}}(\sigma)}(\sigma)=C_{S_{X}}(\sigma)=\operatorname{Aut}P,$
as desired. Combined with Theorem 4.1, these equalities imply the last part of the claim. ∎
Example 4.4.
For all trivial quandles $P=(X,\operatorname{id}_{X})_{\mathrm{perm}}$, we have $C_{S_{X}}(\operatorname{id}_{X})=S_{X}$. So, Proposition 4.3 states that $U_{P}=S_{X}$, and $U_{P}/{\sim}$ is the set of conjugacy classes of $S_{X}$.
Example 4.5.
Let $X:=\{1,2,\dots,n\}$, let $\sigma\in S_{n}$ be an $n$-cycle, and let $P:=(X,\sigma)_{\mathrm{perm}}$. Then $C_{S_{n}}(\sigma)$ is the cyclic subgroup $\langle\sigma\rangle\cong\mathbb{Z}/n\mathbb{Z}$ of $S_{n}$; see, for example, [dummit]*p. 127. Since $\langle\sigma\rangle$ is abelian, Proposition 4.3 implies that $U_{P}/{\sim}=U_{P}=\langle\sigma\rangle$.
Example 4.6.
Let $F$ be the free rack on one element. For all $n\in\mathbb{Z}$, let $\tau_{n}:\mathbb{Z}\to\mathbb{Z}$ be the translation defined by $k\mapsto k+n$. Recall from Example 2.19 that $F$ is isomorphic to the permutation rack $(\mathbb{Z},\tau_{1})_{\mathrm{perm}}$. By Proposition 4.3,
$U_{F}=\operatorname{Aut}F=C_{S_{\mathbb{Z}}}(\tau_{1})=\{\tau_{n}\mid n\in\mathbb{Z}\}\cong\mathbb{Z}.$
In particular, $\operatorname{Aut}F$ is abelian, so $U_{F}=U_{F}/{\sim}$. In other words, there are infinitely many GL-structures on $F$, all of which are translations of $\mathbb{Z}$ and none of which yield isomorphic GL-racks.
4.2.2. Conjugation GL-quandles
Next, we use Theorem 4.1 to classify GL-structures on conjugation quandles of centerless or abelian groups. Given a group $G$, let $\operatorname{Aut}_{\mathsf{Grp}}G$ and $\operatorname{Inn}_{\mathsf{Grp}}G$ denote the automorphism group and inner automorphism group of $G$, respectively.
Proposition 4.7.
Let $G$ be a group, and let $Q:=\operatorname{Conj}G$. If $G$ is abelian, then $U_{Q}=S_{G}$, and $U_{Q}/{\sim}$ is the set of conjugacy classes of $S_{G}$. On the other hand, if $G$ is centerless, then $U_{Q}=\{\operatorname{id}_{G}\}$.
Proof.
If $G$ is abelian, then $Q$ is a trivial quandle, so the first claim follows from Example 4.4. Next, recall a result of Elhamdadi et al. [dihedral]*Theorem 2.3 that $\operatorname{Inn}Q=\operatorname{Inn}_{\mathsf{Grp}}G$ for all groups $G$. Also, recall a result of Bardakov et al. [centerless]*Corollary 2 that $\operatorname{Aut}Q=\operatorname{Aut}_{\mathsf{Grp}}G$ if and only if $G$ is centerless. In this case, Theorem 4.1 states that
$U_{Q}=C_{\operatorname{Aut}_{\mathsf{Grp}}G}(\operatorname{Inn}_{\mathsf{Grp}}G),$
which is the group of central automorphisms of $G$. However, this group is trivial when $G$ is centerless; see, for example, [robinson]*p. 410. ∎
Example 4.8.
Let $n\geq 3$ be an integer or $n=\infty$. Then the symmetric group $S_{n}$ is centerless, so Proposition 4.7 states that the only GL-structure on $\operatorname{Conj}S_{n}$ is $\operatorname{id}_{S_{n}}$. Similarly, if $n\geq 4$, then the alternating group $A_{n}$ is centerless, so the only GL-structure on $\operatorname{Conj}A_{n}$ is $\operatorname{id}_{A_{n}}$.
4.2.3. Takasaki GL-kei
Next, we classify GL-structures on a certain family of Takasaki kei and, as a consequence, all dihedral quandles of odd order.
We first recall a classification result of Bardakov et al. [takasaki-aut]*Theorem 4.2. For all abelian additive groups $A$ without $2$-torsion, $\operatorname{Aut}T(A)$ is isomorphic to the holomorph
$G:=A\rtimes\operatorname{Aut}_{\mathsf{Grp}}A$
of $A$. Under this identification, $\operatorname{Inn}T(A)$ is the semidirect product
$H:=2A\rtimes\{\pm\operatorname{id}_{A}\}\leq G,$
where $-\operatorname{id}_{A}$ denotes inversion. We prove the following result using these identifications.
Proposition 4.9.
If $A$ is an abelian additive group without $2$-torsion, then the only GL-structure on the Takasaki kei $T(A)$ is $\operatorname{id}_{A}$.
Proof.
First, note that for all automorphisms $\psi\in\operatorname{Aut}_{\mathsf{Grp}}A$ such that $\psi(2a)=2a$ for all $a\in A$, we have $02(\psi(a)-a)=$ 0. Since $A$ is $2$-torsion-free, it follows that $\psi(a)=a$, so $\psi=\operatorname{id}_{A}$. Therefore, by Theorem 4.1, it will suffice to show that
$C_{G}(H)\subseteq\{(0,\psi)\in G:\psi|_{2A}=\operatorname{id}_{2A}\}$
since, as we just observed, the right-hand side is the trivial subgroup of $G$. To that end, a direct computation shows that conjugation in $G$ is given by
$(a,\psi)(b,\varphi)(a,\psi)^{-1}=(a+\psi(b)-\psi\varphi\psi^{-1}(a),\psi\varphi\psi^{-1}).$
For all $(b,\varphi)\in H$, we have $\varphi=\pm\operatorname{id}_{A}$. It follows that, for all $(a,\psi)\in C_{G}(H)$ and $(b,\varphi)\in H$,
$(b,\varphi)=(a,\psi)(b,\pm\operatorname{id}_{A})(a,\psi)^{-1}=(a+\psi(b)\mp a,\pm\operatorname{id}_{A}).$
Taking $\varphi:=+\operatorname{id}_{A}$ yields $\psi(b)=b$; since this equality holds for all $(b,\varphi)\in H$ (and, hence, for all $b\in 2A$), we obtain ${\psi|_{2A}=\operatorname{id}_{2A}}$, as desired. Therefore, taking $\varphi:=-\operatorname{id}_{A}$ yields $b=2a+b$, so $02a=$ 0. Since $A$ has no $2$-torsion, it follows that $a=0$, as desired. ∎
Example 4.10.
For all odd integers $n\geq 3$, the cyclic group $\mathbb{Z}/n\mathbb{Z}$ is $2$-torsion-free. By Example 2.7 and Proposition 4.9, the only GL-structure on the dihedral quandle $R_{n}$ of order $n$ is $\operatorname{id}_{\mathbb{Z}/n\mathbb{Z}}$.
4.2.4. Dihedral GL-quandles
Without the assumption that $A$ is $2$-torsion-free, there are infinitely many counterexamples to Proposition 4.9. To show this, we will complete the classification of GL-structures on dihedral quandles $R_{n}$ using the following results of Elhamdadi et al. [dihedral]*Theorems 2.1 and 2.2. The automorphism group $\operatorname{Aut}R_{n}$ is the affine group of $\mathbb{Z}/n\mathbb{Z}$. Thus, $\operatorname{Aut}R_{n}$ is isomorphic to the holomorph
$G:=\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$
of $\mathbb{Z}/n\mathbb{Z}$. Under this identification, $\operatorname{Inn}R_{n}$ is the semidirect product
$H:=2\mathbb{Z}/n\mathbb{Z}\rtimes\{\pm 1\}\leq G,$
which is isomorphic to the dihedral group $D_{n/2}$ of order $n$.
The following result strengthens Example 3.7, and we state it in terms of the above identifications. Our proof uses the fact that conjugation in $G$ is given by
(4) $(a,u)(b,v)(a,u)^{-1}=(a+ub,uv)(-u^{-1}a,u^{-1})=(ub+(1-v)a,v).$
Proposition 4.11.
For all even integers $n\geq 2$, the GL-structures on the dihedral quandle $R_{n}$ are
(5) $U_{R_{n}}=\begin{cases}\{0,n/2\}\rtimes\{1\}&\text{if }4\nmid n,\\ \{0,n/2\}\rtimes\{1,1+n/2\}&\text{if }4\mid n.\end{cases}$
If $4\nmid n$, then $U_{R_{n}}=U_{R_{n}}/{\sim}$, so $|U_{R_{n}}/{\sim}|=2$. Otherwise, the only elements of $U_{R_{n}}$ that are identified in $U_{R_{n}}/{\sim}$ are $(0,1+n/2)$ and $(n/2,1+n/2)$, so $|U_{R_{n}}/{\sim}|=3$.
Proof.
Theorem 4.1 states that, to prove equation (5), it will suffice to show that the right-hand side equals $C_{G}(H)$. For all elements $(a,u)\in C_{G}(H)$ and $(b,v)\in H$, the right-hand side of equation (4) equals $(b,v)$. In particular,
$b=ub+(1-v)a$
for all $b\in 2\mathbb{Z}/n\mathbb{Z}$, so taking $v:=1$ yields $u\in\{1,1+n/2\}$. However, $1+n/2\in(\mathbb{Z}/n\mathbb{Z})^{\times}$ if and only if $4\mid n$, as desired. Since $ub=b$, taking $v:=-1$ yields $02a=$ 0. Hence, $a\in\{0,n/2\}$, as desired. This shows that $C_{G}(H)$ is a subset of the right-hand side of equation (5), and verifying the opposite containment is straightforward.
We now prove the second claim. Since $(0,1)$ is the identity element of $G$, it is not conjugate to any other element of $U_{R_{n}}$. If $4\nmid n$, then we are done.
Otherwise, let $(b,v),(c,w)\in U_{R_{n}}$. Then $(b,v)$ and $(c,w)$ are conjugate in $G$ if and only if there exists an element $(a,u)\in G$ such that $(c,w)$ equals the right-hand side of equation (4). In particular, $w=v$. It follows that neither $(0,1+n/2)$ nor $(n/2,1+n/2)$ is conjugate to $(n/2,1)$ in $G$. On the other hand, taking $(b,v):=(0,1+n/2)$ and $(a,u):=(1,1)$ in equation (4) shows that $(0,1+n/2)$ and $(n/2,1+n/2)$ are conjugate in $G$, so the proof is complete. ∎
Remark 4.12.
If $n\geq 4$ is a multiple of $4$, then equation (5) shows that $U_{R_{n}}\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$, so $U_{R_{n}}$ is abelian. It follows that $(0,1+n/2)$ and $(n/2,1+n/2)$ are conjugate in $G$ but not in $U_{R_{n}}$. Hence, the condition in Theorem 4.1 that $\mathtt{u}_{1}\sim\mathtt{u}_{2}$ in $U_{R}/{\sim}$ does not imply conjugacy in $U_{R}$.
4.3. Automorphism groups of GL-racks
Given a GL-rack $(R,\mathtt{u})$, let $\operatorname{Aut}_{\mathsf{GLR}}(R,\mathtt{u})$ denote its group of GL-rack automorphisms. The following characterization is simply a restatement of the definition of GL-rack automorphisms.
Proposition 4.13.
For all GL-racks $(R,\mathtt{u})$, we have $\operatorname{Aut}_{\mathsf{GLR}}(R,\mathtt{u})=C_{\operatorname{Aut}R}(\mathtt{u})$.
Example 4.14.
Let $L$ be the free GL-rack on one element, and identify $L=(\mathbb{Z}^{2},\sigma,\mathtt{u}_{0})_{\mathrm{perm}}$ as in Proposition 3.19. It is straightforward to show that GL-rack endomorphisms of $L$ are precisely translations of the form $(m,n)\mapsto(m+k,n+\ell)$ for some $(k,\ell)\in\mathbb{Z}^{2}$. Since all maps of this form are permutations of $\mathbb{Z}^{2}$, the mapping $(k,\ell)\mapsto\mathtt{u}_{0}^{\ell}\sigma^{k}$ is a group isomorphism from $\mathbb{Z}^{2}$ to $\operatorname{Aut}_{\mathsf{GLR}}(L)$.
4.3.1. Automorphisms of dihedral GL-quandles
We classify automorphism groups of GL-racks whose underlying racks are dihedral quandles $R_{n}$ of order $n$. Once again, we use a result of Elhamdadi et al. [dihedral]*Theorem 2.1 to identify $\operatorname{Aut}R_{n}\cong\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$.
Proposition 4.15.
Let $n\geq 2$ be an integer, let $R_{n}$ be the dihedral quandle of order $n$, and let $f_{b,v}:\mathbb{Z}/n\mathbb{Z}\to\mathbb{Z}/n\mathbb{Z}$ defined by $k\mapsto b+vk$ be a GL-structure on $R_{n}$; see Proposition 4.11. Let $G:=\operatorname{Aut}_{\mathsf{GLR}}(R_{n},f_{b,v})$. Then
$G\cong\begin{cases}2\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}&\text{if }4\mid n\text{ and }v=1+\frac{n}{2},\\ \mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}&\text{otherwise}.\end{cases}$
In the latter case, $G=\operatorname{Aut}R_{n}$.
Proof.
By Proposition 4.13, $G=C_{\operatorname{Aut}R_{n}}((b,v))$, so for all $(a,u)\in\operatorname{Aut}R_{n}$, we have $(a,u)\in G$ if and only if $(b,v)$ equals the right-hand side of equation (4). Certainly, if $(b,v)=(0,1)$, then $G=\operatorname{Aut}R_{n}$, as claimed. Otherwise, $n$ is even by Example 4.10. By Proposition 4.11, it suffices to only consider the cases that $(b,v)=(n/2,1)$ and $(b,v)=(0,1+n/2)$.
If $(b,v)=(n/2,1)$, then $(a,u)\in G$ if and only if $0=(u-1)(n/2)$ in $\mathbb{Z}/n\mathbb{Z}$. But $u-1$ is even for all elements $u\in(\mathbb{Z}/n\mathbb{Z})^{\times}$, so this equation always holds. In other words, every element $(a,u)\in\operatorname{Aut}R_{n}$ centralizes $(b,v)$, so $G=\operatorname{Aut}R_{n}$.
If ${(b,v)=(0,1+n/2)}$, then $(a,u)\in G$ if and only if $0=(-n/2)a$ in $\mathbb{Z}/n\mathbb{Z}$. This is true if and only if $a$ is even, and there are no restrictions placed on $u$. It follows that $G=2\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$, which completes the proof. ∎
4.4. The center of the category of GL-racks
Recall that the center of a category $\mathcal{C}$ is the commutative monoid $Z(\mathcal{C})$ of natural endomorphisms of the identity functor $\mathbf{1}_{\mathcal{C}}$. In 2018, Szymik [center]*Theorems 5.4 and 5.5 computed that $Z(\mathsf{Rack})=\langle\Theta\rangle\cong\mathbb{Z}$ and $Z(\mathsf{Qnd})\cong\{1\}$. In this subsection, we similarly compute the centers of $\mathsf{GLR}$ and various full subcategories of $\mathsf{GLR}$.
Theorem 4.16.
Let $\Theta$ be the collection of canonical automorphisms $\theta_{R}$ of racks $R$, and let $\mathtt{u}$ be the collection of all GL-structures on racks. Then we have the following:
  1. (Z1)
    The center $Z(\mathsf{GLR})$ is the free abelian group $\langle\Theta,\mathtt{u}\rangle\cong\mathbb{Z}^{2}$ generated by $\Theta$ and $\mathtt{u}$.
  2. (Z2)
    The centers of $\mathsf{GLQ}$ and the category of Legendrian racks are each the free group $\langle\mathtt{u}\rangle\cong\mathbb{Z}$.
  3. (Z3)
    The center of the category of Legendrian quandles is the group $\langle\mathtt{u}\mid\mathtt{u}^{2}=1\rangle\cong\mathbb{Z}/2\mathbb{Z}$.
Proof.
To prove claim 1, let $\eta$ be a natural endomorphism of $\mathbf{1}_{\mathsf{GLR}}$. By definition, $\eta$ is contained in the center $Z(\mathsf{GLR})$ if and only if, for all GL-racks $R_{1}=(X,s,\mathtt{u}_{1})$ and $R_{2}=(Y,t,\mathtt{u}_{2})$ and GL-rack homomorphisms $\varphi\in\operatorname{Hom}_{\mathsf{GLR}}(R_{1},R_{2})$, the following diagram commutes:
${X}$${X}$${Y}$${Y}$$\scriptstyle{\eta_{R_{1}}}$$\scriptstyle{\varphi}$$\scriptstyle{\varphi}$$\scriptstyle{\eta_{R_{2}}}$
To see that $\langle\Theta,\mathtt{u}\rangle\subseteq Z(\mathsf{GLR})$, note by the definition of a GL-rack homomorphism that taking $\eta:=\mathtt{u}^{n}$ with $n\in\mathbb{Z}$ makes the diagram commute. If we take $\eta:=\Theta^{m}$ with $m\in\mathbb{Z}$, then the diagram commutes because $\Theta$ generates $Z(\mathsf{Rack})$. In particular, taking $R_{2}:=R_{1}$ and $\varphi:=\mathtt{u}_{1}^{n}$ shows that $\Theta^{m}$ and $\mathtt{u}^{n}$ commute in $Z(\mathsf{GLR})$.
To see that $Z(\mathsf{GLR})\subseteq\langle\Theta,\mathtt{u}\rangle\cong\mathbb{Z}^{2}$, take $R_{1}$ to be the free GL-rack on one element $(0,0)$ with its identification $R_{1}=(\mathbb{Z}^{2},\sigma,\mathtt{u}_{0})_{\mathrm{perm}}$ from Proposition 3.19. Let $\varphi_{0}:\{(0,0)\}\to Y$ be a set map, let $\varphi:\mathbb{Z}^{2}\to Y$ be the induced GL-rack homomorphism from the universal property of free GL-racks, and fix $\eta\in Z(\mathsf{GLR})$. By Example 4.14, all GL-rack endomorphisms of $R_{1}$ have the form $(m,n)\mapsto(m+k,n+\ell)$ for some $(k,\ell)\in\mathbb{Z}^{2}$, and all of these endomorphisms are in fact automorphisms of $R_{1}$. In particular, $\eta_{R_{1}}$ is an automorphism of $R_{1}$, say
$\eta_{R_{1}}(m,n)=(m+k,n+\ell).$
By commutativity and the proof of Proposition 3.19, $\eta_{R_{2}}$ sends the image of the generator $(0,0)$ of $R_{1}$ under $\varphi_{0}$ to
$\mathtt{u}^{\ell}_{2}\theta_{(Y,t)}^{k}\varphi_{0}(0,0).$
Therefore, $\eta_{R_{2}}$ is completely determined by these powers of $\theta_{(Y,t)}$ and $\mathtt{u}_{2}$. Since $R_{2}$ was an arbitrary GL-rack, it follows that $Z(\mathsf{GLR})\subseteq\langle\Theta,\mathtt{u}\rangle$, as desired. It also follows from Example 4.14 and Proposition 3.18 that $Z(\mathsf{GLR})\cong\mathbb{Z}^{2}$, which proves claim 1.
To prove claim 2, observe that $\Theta$ fixes $\mathbf{1}_{\mathsf{GLQ}}$. It follows from Proposition 3.18 that $Z(\mathsf{GLQ})=\langle\mathtt{u}\rangle\cong\mathbb{Z}$, as desired. Now, consider the full subcategory of $\mathsf{GLR}$ whose objects are Legendrian racks. Using a similar argument as before, one can show using Corollary 3.13 and Proposition 3.18 that the center of this category is
$\langle\Theta,\mathtt{u}\mid\Theta\mathtt{u}=\mathtt{u}\Theta,\,\Theta=\mathtt{u}^{-2}\rangle=\langle\mathtt{u}\rangle\cong\mathbb{Z}.$
This proves claim 2. Similarly, claim 3 follows from Corollary 3.13 and Proposition 3.18. ∎
5. Categorical equivalence of racks and GL-quandles
In this section, we show that the categories of racks and GL-quandles are isomorphic in a way that preserves mediality. This surprising result generalizes the one-to-one correspondences observed in Appendix A.1 and induces isomorphisms of the respective algebraic theories.
5.1. Construction of $F$
We begin by defining a functor $F:\mathsf{Rack}\to\mathsf{GLQ}$. First, we define how $F$ acts on objects.
Proposition 5.1.
Given a rack $R=(X,s)$, define ${\widetilde{s}}:X\to S_{X}$ by
$x\mapsto{\widetilde{s}}_{x}:=\theta_{R}^{-1}s_{x}.$
Then $F(R):=(X,{\widetilde{s}},\theta_{R})$ is a GL-quandle.
Proof.
First, we show that $(X,{\widetilde{s}})$ is a quandle. Part 1 of Proposition 2.17 and part 3 of Proposition 2.16 imply that, for all elements $x,y\in X$,
${\widetilde{s}}_{x}{\widetilde{s}}_{y}=\theta^{-1}_{R}s_{x}\theta^{-1}_{R}s_{y}=\theta^{-2}_{R}s_{x}s_{y}=\theta^{-2}_{R}s_{s_{x}(y)}s_{x}=\theta^{-1}_{R}s_{s_{x}(y)}\theta^{-1}_{R}s_{x}=\theta^{-1}_{R}s_{\theta^{-1}_{R}s_{x}(y)}{\widetilde{s}}_{x}={\widetilde{s}}_{{\widetilde{s}}_{x}(y)}{\widetilde{s}}_{x},$
so $(X,{\widetilde{s}})$ is a rack. Moreover, part 1 of Proposition 2.16 and part 1 of Proposition 2.17 imply that
${\widetilde{s}}_{x}(x)=\theta^{-1}_{R}s_{x}(x)=s_{x}\theta^{-1}_{R}(x)=s_{x}s_{x}^{-1}(x)=x,$
so $(X,{\widetilde{s}})$ is a quandle.
Next, we show that $\theta_{R}$ is a GL-structure on $(X,{\widetilde{s}})$. Indeed, part 1 of Proposition 2.17 and part 3 of Proposition 2.16 imply that, for all $x\in X$,
(6) $\theta_{R}{\widetilde{s}}_{x}=\theta_{R}\theta_{R}^{-1}s_{x}=\theta_{R}^{-1}s_{x}\theta_{R}=\theta^{-1}_{R}s_{\theta_{R}(x)}\theta_{R}={\widetilde{s}}_{\theta_{R}(x)}\theta_{R},$
so $\theta_{R}$ is a rack endomorphism of $(X,{\widetilde{s}})$. Since $\theta_{R}$ is a bijection, we have $\theta_{R}\in\operatorname{Aut}(X,{\widetilde{s}})$, as desired. Moreover, the third expression of equation (6) equals ${\widetilde{s}}_{x}\theta_{R}$, so $\theta_{R}$ is a GL-structure. ∎
We now define how $F$ acts on morphisms.
Proposition 5.2.
For all racks $R=(X,s)$ and $S=(Y,t)$, and for all rack homomorphisms $f\in\operatorname{Hom}_{\mathsf{Rack}}(R,S)$, we have ${f\in\operatorname{Hom}_{\mathsf{GLQ}}(F(R),F(S))}$. So, if we define $F$ to fix $f$ as a set map, then $F$ is a covariant functor from $\mathsf{Rack}$ to $\mathsf{GLQ}$.
Proof.
Certainly, $F$ preserves the identity morphism and composition of morphisms, so we only need to verify that $f$ is a GL-rack homomorphism from $F(R)=(X,{\widetilde{s}},\theta_{R})$ to $F(S)=(Y,{\widetilde{t}},\theta_{S})$. Indeed, since $f\in\operatorname{Hom}_{\mathsf{Rack}}(R,S)$ and $\Theta^{-1}\in Z(\mathsf{Rack})$,
$f{\widetilde{s}}_{x}=f\theta_{R}^{-1}s_{x}=\theta_{S}^{-1}fs_{x}=\theta_{S}^{-1}t_{f(x)}f={\widetilde{t}}_{f(x)}f$
for all $x\in X$, so $f\in\operatorname{Hom}_{\mathsf{Rack}}((X,{\widetilde{s}}),(Y,{\widetilde{t}}))$. Moreover, $f\theta_{R}=\theta_{S}f$ since $\Theta\in Z(\mathsf{Rack})$, so $f$ is a GL-rack homomorphism. ∎
5.2. Construction of $G$
We now define a functor $G:\mathsf{GLQ}\to\mathsf{Rack}$ as the restriction of a functor $\widetilde{G}:\mathsf{GLR}\to\mathsf{Rack}$ to $\mathsf{GLQ}$. First, we define how $\widetilde{G}$ acts on objects.
Proposition 5.3.
Given a GL-rack $R=(X,s,\mathtt{u})$, define ${\hat{s}}:X\to S_{X}$ by
$x\mapsto{\hat{s}}_{x}:=\mathtt{u}s_{x}.$
Then $\widetilde{G}(R):=(X,{\hat{s}})$ is a rack.
Proof.
Fix $x,y\in X$. Since $\mathtt{u}$ is a GL-structure and $(X,s)$ is a rack,
${\hat{s}}_{{\hat{s}}_{x}(y)}{\hat{s}}_{x}=\mathtt{u}s_{\mathtt{u}s_{x}(y)}\mathtt{u}s_{x}=\mathtt{u}^{2}s_{s_{x}(y)}s_{x}=\mathtt{u}^{2}s_{x}s_{y}=\mathtt{u}s_{x}\mathtt{u}s_{y}={\hat{s}}_{x}{\hat{s}}_{y},$
so $(X,{\hat{s}})$ is a rack. ∎
Next, we define how $\widetilde{G}$ acts on morphisms.
Proposition 5.4.
For all GL-racks $R_{1}=(X,s,\mathtt{u}_{1})$ and $R_{2}=(Y,t,\mathtt{u}_{2})$ and GL-rack homomorphisms $g\in\operatorname{Hom}_{\mathsf{GLR}}(R_{1},R_{2})$, we have $g\in\operatorname{Hom}_{\mathsf{Rack}}(\widetilde{G}(R_{1}),\widetilde{G}(R_{2}))$. So, if we define $\widetilde{G}$ to fix $g$ as a set map, then $\widetilde{G}$ is a covariant functor from $\mathsf{GLR}$ to $\mathsf{Rack}$, and $G$ is a functor from $\mathsf{GLQ}$ to $\mathsf{Rack}$.
Proof.
Certainly, $\widetilde{G}$ preserves the identity morphism and composition of morphisms, so we only need to verify that $g\in\operatorname{Hom}_{\mathsf{Rack}}((X,{\hat{s}}),(Y,{\hat{t}}))$. Indeed, since $g$ is a GL-rack homomorphism,
$g{\hat{s}}_{x}=g\mathtt{u}_{1}s_{x}=\mathtt{u}_{2}gs_{x}=\mathtt{u}_{2}t_{g(x)}g={\hat{t}}_{g(x)}g$
for all $x\in X$, as desired. ∎
5.3. Isomorphism of categories
Having defined $F$ and $G$, we are now ready to prove the main results of this section.
Theorem 5.5.
The functors $F$ and $G$ are isomorphisms of categories $\mathsf{Rack}\cong\mathsf{GLQ}$, and they restrict to isomorphisms $\mathsf{Rack}_{\mathrm{med}}\cong\mathsf{GLQ}_{\mathrm{med}}$.
Proof.
To show that $F$ and $G$ are isomorphisms of categories, we only need to show that $GF$ and $FG$ fix the objects in the appropriate categories; $GF$ and $FG$ clearly fix the morphisms. To that end, let $R=(X,s)$ be a rack. To see that ${GF(R)=R}$, note that $GF(R)=(X,\hat{{\widetilde{s}}})$, where
$\hat{{\widetilde{s}}}_{x}=\theta_{R}{\widetilde{s}}_{x}=\theta_{R}\theta^{-1}_{R}s_{x}=s_{x}$
for all $x\in X$. That is, $\hat{{\widetilde{s}}}=s$, so $GF=\mathbf{1}_{\mathsf{Rack}}$, as desired.
Next, let $Q=(X,s,\mathtt{u})$ be a GL-quandle. We must show that $FG(Q)=Q$. Note that, for all $x\in X$, we have
$\hat{s}_{x}^{-1}=s_{x}^{-1}\mathtt{u}^{-1}=\mathtt{u}^{-1}s_{x}^{-1}.$
Since $(X,s)$ is a quandle, we also have $x=s_{x}(x)$, so $s_{x}^{-1}(x)=x$. Now, to see that $FG(Q)=Q$, write $FG(Q)=(X,\widetilde{{\hat{s}}},\theta_{G(Q)})$. Part 1 of Proposition 2.16 implies that, for all elements $x,y\in X$,
$\widetilde{{\hat{s}}}_{y}(x)=\theta^{-1}_{G(Q)}{\hat{s}}_{y}(x)={\hat{s}}_{y}\theta^{-1}_{G(Q)}(x)=\mathtt{u}s_{y}\hat{s}^{-1}_{x}(x)=s_{y}\mathtt{u}\mathtt{u}^{-1}s_{x}^{-1}(x)=s_{y}(x).$
Thus, $\widetilde{{\hat{s}}}_{y}=s_{y}$. Since $y\in X$ was arbitrary, this shows that $\widetilde{{\hat{s}}}=s$, as desired. Similarly,
$\theta_{G(Q)}(x)=\hat{s}_{x}(x)=\mathtt{u}s_{x}(x)=\mathtt{u}(x)$
for all $x\in X$, so $\theta_{G(Q)}=\mathtt{u}$. Hence, $FG=\mathbf{1}_{\mathsf{GLQ}}$, so $F$ and $G$ are isomorphisms of categories, as desired. Since $\theta^{-1}$ and $\mathtt{u}$ are always rack automorphisms, the final claim follows straightforwardly from the definition of mediality that uses homomorphisms. ∎
Corollary 5.6.
In the category of algebraic theories, the theory of racks and the theory of GL-quandles are isomorphic.
Proof.
Since $F$ and $G$ are left and right adjoints to each other, they both preserve limits and colimits. In particular, they preserve finite products and filtered colimits, so by [alg-theories]*Lemma 3.8.3, $F$ and $G$ are algebraic functors. Since $F$ and $G$ are equivalences of categories, the claim follows directly from [alg-theories]*Proposition 3.12.1. ∎
6. Tensor products of racks and GL-racks
In 2014, Crans and Nelson [Hom]*Section 8.1 categorified the results of Corollary 2.22 by considering universal-algebraic tensor products of medial quandles. However, these tensors remain unexplored in the literature.
With this motivation, we consider universal-algebraic tensor products of racks and GL-racks. We show that, unlike with groups, $\mathsf{Rack}$ and $\mathsf{GLR}$ have tensor units. This suggests that $\mathsf{GLR}$ and $\mathsf{GLR}_{\mathrm{med}}$ make natural settings for functorial invariants of Legendrian links.
6.1. Construction and universal property
We begin by considering tensor products of racks and GL-racks constructed via universal algebra. We discuss these tensors’ universal properties and show that the induced symmetric monoidal structures on $\mathsf{Rack}_{\mathrm{med}}$ and ${\mathsf{GLR}_{\mathrm{med}}}$ are closed.
Definition 6.1.
If $R_{1}=(X,r,\mathtt{u}_{1})$ and $R_{2}=(Y,t,\mathtt{u}_{2})$ are GL-racks, then we define their tensor product, denoted by $R_{1}\otimes R_{2}$, to be the free GL-rack $\operatorname{Free}_{\mathsf{GLR}}(X\times Y)$ modulo the smallest congruence such that the following hold for all $x,x_{1},x_{2}\in X$ and $y,y_{1},y_{2}\in Y$, writing $x\otimes y:=(x,y)$:
  1. (T1)
    $s_{x\otimes y_{2}}(x\otimes y_{1})\sim x\otimes t_{y_{2}}(y_{1})$, and  $s^{-1}_{x\otimes y_{2}}(x\otimes y_{1})\sim x\otimes t^{-1}_{y_{2}}(y_{1})$.
  2. (T2)
    $s_{x_{2}\otimes y}(x_{1}\otimes y)\sim r_{x_{2}}(x_{1})\otimes y$, and  $s^{-1}_{x_{2}\otimes y}(x_{1}\otimes y)\sim r^{-1}_{x_{2}}(x_{1})\otimes y$.
  3. (T3)
    $\mathtt{u}(x\otimes y)\sim\mathtt{u}_{1}(x)\otimes y\sim x\otimes\mathtt{u}_{2}(y).$
We also define the medial tensor product of $R_{1}$ and $R_{2}$ to be
$R_{1}\otimes_{\mathrm{med}}R_{2}:=(R_{1}\otimes R_{2})_{\mathrm{med}}.$
Similarly, we define the tensor product of two racks $R_{1}=(X,r)$ and $R_{2}=(Y,t)$ as $\operatorname{Free}_{\mathsf{Rack}}(X\times Y)$ modulo the smallest congruence generated by relations 1 and 2 above, and we define the medial tensor product of racks similarly.
Remark 6.2.
Relation 3 implies a similar relation involving $\mathtt{u}^{-1}$. Namely, in $(X,r,\mathtt{u}_{1})\otimes(Y,t,\mathtt{u}_{2})$,
$\mathtt{u}^{-1}(x\otimes y)=\mathtt{u}_{1}^{-1}(x)\otimes y=x\otimes\mathtt{u}^{-1}_{2}(y)$
for all $x\in X$ and $y\in Y$. To see this, compute
$\mathtt{u}_{1}^{-1}(x)\otimes y=\mathtt{u}^{-1}\mathtt{u}(\mathtt{u}_{1}^{-1}(x)\otimes y)=\mathtt{u}^{-1}(\mathtt{u}_{1}\mathtt{u}_{1}^{-1}(x)\otimes y)=\mathtt{u}^{-1}(x\otimes y),$
and similarly for $x\otimes\mathtt{u}^{-1}_{2}(y)$.
We note that the medialized associated quandles of the tensor products of racks in Definition 6.1 recover the tensor products of medial quandles that Crans and Nelson [Hom]*Section 8.1 introduced in 2014. On the other hand, these tensors are distinct from the tensor products that Kamada [tensors]*Definition 3.1 introduced in 2021, which are sets without canonical rack structures.
6.1.1. Bihomomorphisms
The discussion in [alg-theories]*p. 171 shows that the tensor products in Definition 6.1 are each characterized by a universal factorizing property. Before we can state this property, we will need the following definition from universal algebra. This is a special case of a more general definition for arbitrary algebraic theories; see, for example, [alg-theories]*Definition 3.10.2.
Definition 6.3.
Let $(X,r)$, $(Y,s),$ and $(Z,t)$ be racks. We say that a map $\beta_{0}:X\times Y\to Z$ is a rack bihomomorphism if, for all $x\in X$ and $y\in Y$, the restricted maps $\beta_{0}(-,y):X\to Z$ and $\beta_{0}(x,-):Y\to Z$ are rack homomorphisms. GL-rack bihomomorphisms are defined similarly.
Example 6.4.
Let $R=(X,r)$ be a rack, and let $S=(Y,s)$ and $T=(Z,t)$ be medial racks. Recall from Corollary 2.22 that $H_{1}:=\operatorname{Hom}_{\mathsf{Rack}}(R,S)$, $H_{2}:=\operatorname{Hom}_{\mathsf{Rack}}(S,T)$, and $H_{3}:=\operatorname{Hom}_{\mathsf{Rack}}(R,T)$ each have a canonical rack structure.
In analogy with the composition of $A$-linear maps in the category of modules over a ring $A$, the composition map $\beta_{0}:H_{1}\times H_{2}\to H_{3}$ is a rack bihomomorphism. To see this, fix a homomorphism $g\in H_{2}$. For all homomorphisms $\varphi,\psi\in H_{1}$ and elements $x\in X$, we have
$\beta_{0}({\widetilde{s}}_{\psi}(\varphi),g)(x)=g{\widetilde{s}}_{\psi}(\varphi)(x)=gs_{\psi(x)}\varphi(x)=t_{g\psi(x)}g\varphi(x)=\widetilde{t}_{g\psi}g\varphi(x)=\widetilde{t}_{\beta_{0}(\psi,g)}\beta_{0}(\varphi,g)(x),$
so the restriction $\beta_{0}(-,g):H_{1}\to H_{3}$ is a rack homomorphism. Similarly, for all homomorphisms $f\in H_{1}$, the restriction $\beta_{0}(f,-):H_{2}\to H_{3}$ is a homomorphism, so $\beta_{0}$ is a bihomomorphism.
It is straightforward to verify using Theorem 3.20 that if $R$, $S$, and $T$ also have GL-structures, then the restriction of $\beta_{0}$ to the respective hom-sets in $\mathsf{GLR}$ is a GL-rack bihomomorphism.
6.1.2. Universal property of tensor products
As suggested in Example 6.4, the definition of a rack bihomomorphism analogizes the definition of a bilinear map in the category of modules over a ring. The universal property of tensor products, which we state below, extends this analogy.
Proposition 6.5.
Let $R_{1}$ and $R_{2}$ be racks. Then $R_{1}\otimes R_{2}$ is characterized up to isomorphism by a universal property. Namely, there exists a rack bihomomorphism $\psi$ from $R_{1}\times R_{2}$ to $R_{1}\otimes R_{2}$ such that, for all racks $R_{3}$ and rack bihomomorphisms $\beta_{0}:R_{1}\times R_{2}\to R_{3}$, there exists a unique rack homomorphism $\beta:R_{1}\otimes R_{2}\to R_{3}$ such that $\beta_{0}=\beta\psi$. In particular, the following diagram commutes:
[…]
A similar result holds for GL-racks.
Proof.
This is a consequence of universal algebra. Regard racks and GL-racks as algebraic theories with binary operations $s_{-}(-)$ and $s^{-1}_{-}(-)$, along with unary operations $\mathtt{u}$ and $\mathtt{u}^{-1}$ for GL-racks. Due to Remark 6.2, the tensor product in Definition 6.1 is precisely the tensor product constructed in the proof of [alg-theories]*Theorem 3.10.3, and $\psi(x,y)=x\otimes y$ for all $(x,y)\in X\times Y$. Thus, the claim follows from the discussion in [alg-theories]*p. 171. ∎
Example 6.6.
A rack $(X,s)$ is called left-distributive if
(8) $s_{s_{a}(b)}(x)=s_{s_{a}(x)}(s_{b}(x))$
for all $a,b,x\in X$. The name comes from the fact that, in terms of the right-distributive binary operation $\triangleright$ often used in the literature, equation (8) states that
$x\triangleright(b\triangleright a)=(x\triangleright b)\triangleright(x\triangleright a).$
For example, it is straightforward to verify using equation (1) that medial quandles are left-distributive. Moreover, self-distributive quasigroups are precisely left-distributive quandles satisfying an axiom called the Latin condition; see, for example, [quandlebook]*p. 143.
Evidently, a rack $R=(X,s)$ is left-distributive if and only if, for all $x\in X$, the map $X\to X$ defined by $y\mapsto s_{y}(x)$ is an endomorphism of $R$. Equivalently, the map $\beta_{0}:X\times X\to X$ defined by $(x,y)\mapsto s_{y}(x)$ is a bihomomorphism from $R\times R$ to $R$. In this case, the universal property of $R\otimes R$ implies the existence of a unique homomorphism $\beta\in\operatorname{Hom}_{\mathsf{Rack}}(R\otimes R,R)$ such that $\beta(x\otimes y)=s_{y}(x)$ for all $x,y\in X$.
6.1.3. Internal hom-tensor adjunctions
Mirroring the work of Crans and Nelson [Hom]*Theorem 12 on the category of medial quandles, the next result continues the analogy with the category of modules over a ring by describing internal hom-tensor adjunctions in $\mathsf{Rack}_{\mathrm{med}}$ and $\mathsf{GLR}_{\mathrm{med}}$.
Theorem 6.7.
The categories $\mathsf{Rack}_{\mathrm{med}}$ and ${\mathsf{GLR}_{\mathrm{med}}}$ are closed symmetric monoidal with respect to the medial tensor product $\otimes_{\mathrm{med}}$ in each category and the closed structures $\operatorname{Hom}_{\mathsf{Rack}_{\mathrm{med}}}(-,-)$ and $\operatorname{Hom}_{{\mathsf{GLR}_{\mathrm{med}}}}(-,-)$ from Corollary 2.22 and Theorem 3.20, respectively.
Proof.
Recall that the algebraic theories of medial racks and medial GL-racks are commutative. Therefore, both claims are special cases of a general result for commutative algebraic theories; see, for example, [alg-theories]*Theorem 3.10.3.
In particular, $\otimes_{\mathrm{med}}$ is precisely the tensor product constructed in the proof of [alg-theories]*Theorem 3.10.3, and the tensor unit is the free rack (resp. GL-rack) $L$ on one element. Indeed, combining Example 2.19 (resp. Proposition 3.19) with Example 2.12 shows that $L$ is medial. ∎
6.2. Nonmedial tensor products
One may ask how much of the structure in Theorem 6.7 remains if we drop the mediality assumption. By [alg-theories]*Theorem 3.10.3, we lose the closed structure because racks and GL-racks are noncommutative algebraic theories. Nevertheless, we show that tensor products in $\mathsf{Rack}$ and $\mathsf{GLR}$ surprisingly retain tensor units.
Although universal-algebraic tensor products with the appropriate universal properties exist for all algebraic theories, they are often not well-behaved in noncommutative algebraic theories. For example, the universal-algebraic tensor product of groups $G\otimes H$ is isomorphic to the usual $\mathbb{Z}$-module tensor product $G^{\operatorname{ab}}\otimes_{\mathbb{Z}}H^{\operatorname{ab}}$ of the abelianizations of $G$ and $H$; see, for example, [alg-theories]*p. 171. In particular, no tensor unit exists.
In this subsection, we show that the pathologies in the previous paragraph do not apply to tensor products of racks or GL-racks, even though racks and GL-racks are noncommutative algebraic theories. In particular, while the universal-algebraic tensor product of groups is always abelian, tensor products of racks and GL-racks are not necessarily medial.
Lemma 6.8.
Let $R_{1}=(X,r)$, $R_{2}=(Y,s)$, and $R_{3}=(Z,t)$ be racks, and let $\beta_{0}:X\times Y\to Z$ be a rack bihomomorphism from $R_{1}\times R_{2}$ to $R_{3}$. Then for all integers $k\in\mathbb{Z}$, elements $x\in X$, and elements $y\in Y$,
$\beta_{0}(\theta_{R_{1}}^{k}(x),y)=\beta_{0}(x,\theta_{R_{2}}^{k}(y)).$
Proof.
For all $y\in Y$, the restriction $\beta_{0}(-,y):X\to Z$ is a homomorphism from $R_{1}$ to $R_{3}$. For all integers $k\in\mathbb{Z}$, the inclusion $\Theta^{k}\in Z(\mathsf{Rack})$ implies that
$\beta_{0}(\theta_{R_{1}}^{k}(x),y)=\theta^{k}_{R_{3}}\beta_{0}(x,y)$
for all $x\in X$. A similar argument using the restriction $\beta_{0}:(x,-):Y\to Z$ yields the claim. ∎
Theorem 6.9.
Let $F$ be the free rack on one element. Then for all racks $R$, we have natural isomorphisms $R\otimes F\cong R\cong F\otimes R$.
Similarly, let $L$ be the free GL-rack on one element. Then for all GL-racks $R$, we have natural isomorphisms $R\otimes L\cong R\cong L\otimes R$.
Proof.
First, let $R=(X,s)$ be a rack. We will show that $R\otimes F\cong R$; the proof that $F\otimes R\cong R$ is similar, and naturality is straightfoward to verify from there. Identify $F=(\mathbb{Z},\sigma)_{\mathrm{perm}}$ as in Example 2.19.
By Proposition 6.5, it will suffice to show that $R$ satisfies the universal property of $R\otimes F$ with $\psi:X\times\mathbb{Z}\to X$ defined by
$(x,k)\mapsto\theta^{k}_{R}(x).$
First, we show that $\psi$ is a rack bihomomorphism. Fix $x\in X$, and let $s^{\mathbb{Z}}:\mathbb{Z}\to\{\sigma\}$ denote the rack structure of $F$. For all $k,n\in\mathbb{Z}$,
$\psi(x,s^{\mathbb{Z}}_{k}(n))=\psi(x,\sigma(n))=\psi(x,n+1)=\theta^{n+1}_{R}(x)=\theta^{n}_{R}s_{x}(x)=s_{\theta_{R}^{k}(x)}\theta_{R}^{n}(x)=s_{\psi(x,k)}\psi(x,n).$
In the fifth equality, we have used part 1 of Proposition 2.17 and part 3 of Proposition 2.16. So, ${\psi(x,-):\mathbb{Z}\to Y}$ is a rack homomorphism from $F$ to $R^{\prime}$. Next, fix $k\in\mathbb{Z}$. For all $x,y\in X$,
$\psi(s_{x}(y),k)=\theta^{k}_{R}s_{x}(y)=s_{x}\theta^{k}_{R}(y)=s_{\theta_{R}^{k}(x)}\theta^{k}_{R}(y)=s_{\psi(x,k)}\psi(y,k)$
by part 1 of Proposition 2.17 and part 3 of Proposition 2.16, so $\psi(-,k):X\to Y$ is a rack homomorphism from $R$ to $R^{\prime}$. Therefore, $\psi$ is a rack bihomomorphism, as desired.
Now, let $R^{\prime}=(Y,t)$ be a rack, and let $\beta_{0}:X\times\mathbb{Z}\to Y$ be a rack bihomomorphism from $R\times F$ to $R^{\prime}$. Define $\beta:X\to Y$ by $x\mapsto\beta_{0}(x,0)$. Then $\beta$ is precisely the restriction $\beta_{0}(-,0):X\to Y$, so $\beta$ is a rack homomorphism, as desired. Since $\theta_{F}=\sigma$, Lemma 6.8 implies that
$\beta\psi(x,k)=\beta\theta^{k}_{R}(x)=\beta_{0}(\theta^{k}_{R}(x),0)=\beta_{0}(x,\sigma^{k}(0))=\beta_{0}(x,k)$
for all $(x,k)\in X\times\mathbb{Z}$, so diagram (7) commutes. Finally, uniqueness follows from the surjectivity of $\psi$. Hence, $R$ satisfies the universal property of $R\otimes F$, so $R\cong R\otimes F$, as claimed.
For the second part of the claim, let $R=(X,s,\mathtt{u})$ be a GL-rack. Identify $L=(\mathbb{Z}^{2},\sigma,\mathtt{u}_{0})_{\mathrm{perm}}$ as in Proposition 3.19. Again, it suffices to show that $R$ satisfies the universal property of $R\otimes L$ with $\psi:X\times\mathbb{Z}^{2}$ defined by
$(x,m,n)\mapsto\mathtt{u}^{n}\theta^{m}_{R}(x).$
The proof that $\psi$ is a GL-rack bihomomorphism is similar to the argument given above; we leave the details to the reader.
Given a GL-rack $R^{\prime}=(Y,t,\mathtt{u}_{2})$ and a GL-rack bihomomorphism $\beta_{0}:X\times\mathbb{Z}^{2}\to Y$ from $R\times L$ to $R^{\prime}$, define ${\beta:X\to Y}$ by $x\mapsto\beta_{0}(x,0,0)$. Once again, $\beta$ is a GL-rack homomorphism since it is the restriction $\beta_{0}(-,0,0):X\to Y$. The commutativity of diagram (7) and uniqueness of $\beta$ are shown in a similar way as before, so $R\cong R\otimes L$. ∎
Recall that the free rack on one element is medial, and similarly for the free GL-rack on one element. Therefore, Theorem 6.9 implies the following result—a surprising statement considering that the universal-algebraic tensor product of groups is necessarily abelian.
Corollary 6.10.
Even if one of the tensor factors is medial, tensor products of racks are not necessarily medial, and similarly for tensor products of GL-racks.
It would be interesting to study more properties and applications of the tensor products in Definition 6.1. In Section 8, we propose future work in this direction.
7. Distinguishing Legendrian knots
In this section, we use GL-racks to distinguish Legendrian knots whose classical invariants are identical. This answers a question of Kimura [bi]*Section 4 and completes the classification of Legendrian $8_{13}$ knots in the extended Legendrian knot atlases of Bhattacharyya et al. [atlas-site] and Petkova and Schwartz [new-atlas]. By convention, we take all Legendrian links to be oriented.
7.1. Legendrian knots and links
In this subsection, we briefly summarize the Legendrian isotopy problem. Although we establish the relevant concepts here, we also refer the reader to [what] for an accessible introduction to Legendrian knot theory. For a more formal contact-geometric treatment, we refer the reader to [etnyre].
7.1.1. Definitions
Recall that a smooth link $\Lambda\subset\mathbb{R}^{3}$ is called Legendrian if it lies everywhere tangent to the standard contact structure $\ker(dz-y\,dx)$ on $\mathbb{R}^{3}$ depicted in Figure 1.
An assignment of a plane to each point in 3D Euclidean space. The planes’ slopes are determined by the y-value at each point.
Figure 1. The standard contact structure on $\mathbb{R}^{3}$. Reprinted from [us]*Figure 1.
The front projection of a Legendrian link to the $xz$-plane viewed from the negative $y$-direction has cusps instead of vertical tangencies. Also, the overstrand of every crossing has a lower slope than the understrand. For example, Figure 2 depicts front projections of two Legendrian unknots.
Two oriented knot diagrams representing Legendrian unknots. The knots have no crossings but have different numbers of cusps.
Figure 2. Front projections of nonequivalent Legendrian unknots.
One major problem in contact geometry is the classification of Legendrian knots up to Legendrian isotopy or, equivalently, up to finite sequences of Legendrian Reidemeister moves. For example, the Legendrian unknots in Figure 2 are nonequivalent; that is, they are not related by any Legendrian Reidemeister moves.
7.1.2. Legendrian nonsimplicity
To address the Legendrian isotopy problem, contact topologists have developed various invariants of Legendrian links. To distinguish Legendrian links, it suffices to distinguish their invariants.
[…]
One of the main challenges of the Legendrian isotopy problem is the existence of Legendrian nonsimple smooth knot types, which have nonequivalent Legendrian representatives whose classical invariants are identical. Oftentimes, other invariants also fail to distinguish representatives of Legendrian nonsimple knot types, especially for knot types of high arc index. This phenomenon opens up many conjectures in the Legendrian knot atlases of Chongchitmate and Ng [atlas], Bhattacharyya et al. [atlas-site], and Petkova and Schwartz [new-atlas], several of which we will address using GL-racks.
7.2. GL-racks of Legendrian links
In this subsection, we briefly recall how to assign a canonical GL-rack to a Legendrian link. We also discuss GL-rack coloring invariants.
7.2.1. Invariance of GL-racks
In 2023, Karmakar et al. [karmakar] and Kimura [bi] introduced GL-rack invariants of Legendrian links. These generalize similar invariants introduced by Kulkarni and Prathamesh [original] and Ceniceros et al. [ceniceros].
The Legendrian Reidemeister moves are encoded in the bi-Legendrian rack axioms; see [bi]*Figures 6–8. Consequently, every Legendrian link $\Lambda$ is assigned a canonical GL-rack $\mathcal{G}(\Lambda)$ using any of its front projections, and the isomorphism type of $\mathcal{G}(\Lambda)$ is an invariant of $\Lambda$; see [karmakar]*Theorem 4.3. That said, there exist nonequivalent Legendrian knots with isomorphic GL-racks (see [leg-cocycle]*Examples 21–24), so $\mathcal{G}(\Lambda)$ is not a complete invariant of Legendrian links.
7.2.2. Assignment of GL-racks to Legendrian links
We summarize the construction of $\mathcal{G}(\Lambda)$ from [karmakar]*Section 4. During a traversal of any front projection of $\Lambda$, label the strands (i.e., connected components) of the projection by $x_{1},\dots,x_{k}$. Partition each strand into substrands each containing no cusps, labeling the substrands as stipulated in (U) and (D) of Figure 3.
Negative and positive crossings and up- and down-cusps as they appear in Legendrian knots labeled with relations on GL-racks.
Figure 3. Crossing relations and cusp labels used to obtain a presentation of $\mathcal{G}(\Lambda)$.
Using Remark 3.16, define $\mathcal{G}(\Lambda)$ to be the free GL-rack on the set $\{x_{1},\dots,x_{k}\}$ modulo the congruence generated by the relations specified in (N) and (P) of Figure 3 at all crossings between uncusped substrands. Relations (N) and (P) respectively correspond to negative and positive orientations of crossings in the skein-theoretic sense; see [quandlebook]*p. 87 for details.
Note the similarity between the definition of $\mathcal{G}(\Lambda)$ and the combinatorial construction of fundamental quandles of smooth links; see, for example, [joyce]*Section 15 and cf. [leg-cocycle]*Remark 23.
Example 7.1.
Let $q\geq 3$ be an odd integer, let $L$ be a $(2,-q)$-torus knot, and let $\Lambda$ be the Legendrian representative of $L$ whose Thurston–Bennequin and rotation numbers are maximal. [^1] This choice is well-defined because torus knots are Legendrian simple; see [etnyre]*Subsection 5.2. In this example, we compute $\mathcal{G}(\Lambda)$ using the front projection in Figure 4.
An oriented Legendrian negative torus knot with its cusped strands and uncusped substrands labeled.
Figure 4. Front projection and cusp labels of the Legendrian $(2,-q)$-torus knot $\Lambda$ whose Thurston–Bennequin and rotation numbers are maximal.
Starting at any crossing (which, in Figure 4, we arbitrarily choose to be the bottommost crossing), traverse $\Lambda$ along its depicted orientation. All crossings in this particular front projection are negative, so by imposing relation (N) from Figure 3, we compute that $\mathcal{G}(\Lambda)$ is the free GL-rack on the set $\{x_{1},\dots,x_{q}\}$ modulo the congruence generated by the relations
$s_{\mathtt{u}(x_{1})}(x_{q})=\mathtt{u}\mathtt{d}(x_{2}),\quad s_{\mathtt{d}(x_{q})}(x_{q-1})=\mathtt{d}\mathtt{u}(x_{1}),\quad\text{and }s_{\mathtt{d}(x_{i-1})}(x_{i-2})=\mathtt{d}^{2}(x_{i})\quad\text{for all }3\leq i\leq q.$
By bi-Legendrian rack axiom 3, these crossing relations simplify to
$s_{x_{1}}(x_{q})=\mathtt{u}\mathtt{d}(x_{2}),\quad s_{x_{q}}(x_{q-1})=\mathtt{d}\mathtt{u}(x_{1}),\quad\text{and }s_{x_{i-1}}(x_{i-2})=\mathtt{d}^{2}(x_{i})\quad\text{for all }3\leq i\leq q.$
7.2.3. Legendrian coloring invariants
At the time of writing, the easiest known way to distinguish GL-racks of Legendrian links is by using the GL-rack coloring number $\operatorname{Col}(\Lambda,R)$ of each link $\Lambda$ with respect to a given GL-rack $R$. Defined to be the cardinality of the hom-set $\operatorname{Hom}_{\mathsf{GLR}}(\Lambda,R)$, $\operatorname{Col}(\Lambda,R)$ is an invariant of $\Lambda$; see [bi]*Proposition 3.9. Explicitly, a sufficient condition for two Legendrian links $\Lambda_{1},\Lambda_{2}$ to be nonequivalent is the existence of a GL-rack $R$ such that $\operatorname{Col}(\Lambda_{1},R)\neq\operatorname{Col}(\Lambda_{2},R)$.
[…]
Proposition 7.2.
Let $\Lambda$ be a Legendrian link. Then for all medial GL-racks $M$, the isomorphism type of $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda),M)$ as a medial GL-rack is an invariant of $\Lambda$.
7.3. Distinguishing results
In this subsection, we use GL-rack coloring numbers to distinguish various Legendrian knots whose classical invariants are identical. Specifically, we consider conjecturally distinct pairs of Legendrian $6_{2}$, $8_{10}$, and $8_{13}$ knots listed in the atlases of Chongchitmate and Ng [atlas], Bhattacharyya et al. [atlas-site], and Petkova and Schwartz [new-atlas].
7.3.1. Discussion
Prior to this work, various authors have used coloring invariants to distinguish Legendrian representatives of Legendrian simple knot types. Kulkarni and Prathamesh [original]*Main Theorem 2, Kimura [bi]*Theorem 4.1, and Karmakar et al. [karmakar]*Theorem 4.6 each used coloring invariants to distinguish infinitely many Legendrian unknots. Karmakar et al. [karmakar]*Theorem 4.7 also used GL-rack coloring numbers to distinguish infinitely many Legendrian trefoils, and Ceniceros et al. [ceniceros]*Example 16 used them to distinguish connected sums of Legendrian trefoils.
However, to our knowledge at the time of writing, the following is the first application of GL-racks to distinguish Legendrian knots whose classical invariants are identical.
7.3.2. Finding coloring maps
We employ a well-known characterization of homomorphisms of models of an algebraic theory in terms of generators and relations. Similarly to group homomorphisms, GL-rack homomorphisms from $\mathcal{G}(\Lambda)$ to $R$ are determined by the images of the generators of $\mathcal{G}(\Lambda)$ and whether these images satisfy the relations of $\mathcal{G}(\Lambda)$ as equations in $R$; see [leg-cocycle]*Remark 12.
In Appendix A.3, we describe algorithms that compute GL-rack coloring numbers. These algorithms helped us select the GL-racks and coloring homomorphisms used in this subsection. We provide an implementation of our search algorithms in a GitHub repository [my-code].
[…]
7.3.3. Legendrian $6_{2}$ knots
In 2021, Dynnikov and Prasolov [6-2]*Proposition 2.3 used impressive topological and combinatorial machinery to distinguish the two Legendrian $6_{2}$ knots with classical invariants $(\operatorname{tb},\operatorname{rot})=(-7,2)$ in Figure 5, settling a conjecture of Chongchitmate and Ng [atlas]. At the time of writing, we are unaware of any other proofs of this conjecture. Indeed, linearized contact homology and the ruling invariant fail to distinguish these Legendrian knots; see [atlas].
Two oriented Legendrian 6 2 knots with equal classical invariants. All cusped strands and uncusped substrands are labeled.
Figure 5. Front projections of the two Legendrian representatives of the topological knot $6_{2}$ with $(\operatorname{tb},\operatorname{rot})=(-7,2)$ given in [atlas].
Using GL-rack coloring numbers, we offer a simpler proof that these Legendrian knots are nonequivalent. Let $\Lambda_{1}$ and $\Lambda_{2}$ be the Legendrian knots on the left and right of Figure 5, respectively. First, we compute presentations for $\mathcal{G}(\Lambda_{1})$ and $\mathcal{G}(\Lambda_{2})$. Using Figure 3, we find that $\mathcal{G}(\Lambda_{1})$ is the free GL-rack on the set $\{x_{1},\dots,x_{6}\}$ modulo the congruence generated by the following relations:
(9) $\mathcal{G}(\Lambda_{1})\begin{cases}s_{\mathtt{u}(x_{1})}\mathtt{u}(x_{4})=x_{5}\iff s_{x_{1}}\mathtt{u}(x_{4})=x_{5},\\ s_{x_{4}}\mathtt{d}\mathtt{u}(x_{1})=x_{2},\\ s_{\mathtt{d}(x_{2})}(x_{1})=\mathtt{d}^{2}(x_{6})\iff s_{x_{2}}(x_{1})=\mathtt{d}^{2}(x_{6}),\\ s_{x_{5}}(x_{3})=\mathtt{u}\mathtt{d}(x_{2}),\\ s_{\mathtt{d}(x_{3})}(x_{6})=\mathtt{d}(x_{5})\iff s_{x_{3}}(x_{6})=\mathtt{d}(x_{5}),\\ s_{\mathtt{d}(x_{6})}(x_{4})=\mathtt{d}^{2}(x_{3})\iff s_{x_{6}}(x_{4})=\mathtt{d}^{2}(x_{3}).\end{cases}$
We have simplified the first, third, fifth, and sixth relations using bi-Legendrian rack axiom 3.
Similarly, we compute $\mathcal{G}(\Lambda_{2})$ to be the free GL-rack on the set $\{x_{1},\dots,x_{7}\}$ modulo the congruence generated by the following relations:
(10) $\mathcal{G}(\Lambda_{2})\begin{cases}s_{x_{1}}\mathtt{u}\mathtt{d}^{2}(x_{3})=x_{4},&s_{x_{5}}(x_{3})=\mathtt{d}(x_{2}),\\ s_{x_{1}}(x_{6})=x_{7},&s_{x_{3}}(x_{6})=\mathtt{d}^{2}(x_{5}),\\ s_{x_{6}}(x_{2})=\mathtt{u}(x_{1}),&s_{x_{3}}(x_{7})=x_{1}.\\ s_{x_{2}}(x_{5})=\mathtt{d}(x_{4}),\\ \end{cases}$
Theorem 7.3.
The two Legendrian $6_{2}$ knots with $(\operatorname{tb},\operatorname{rot})=(-7,2)$ in Figure 5 are nonequivalent; they are distinguishable using coloring numbers with respect to a permutation GL-rack of order 3.
Proof.
Let $Y:=\{1,2,3\}$. In cycle notation, let $\sigma\in S_{3}$ be the permutation $(123)$. In the notation of Example 3.5, let $R:=(Y,\sigma,\sigma^{-1})_{\mathrm{perm}}$, so that $R$ is the $11$th GL-rack in Table B.2. By Proposition 3.11, we have $\mathtt{d}=\operatorname{id}_{Y}$ in $R$.
We will show that $\operatorname{Col}(\Lambda_{2},R)>\operatorname{Col}(\Lambda_{1},R)$. To that end, let $X$ denote the underlying set of $\mathcal{G}(\Lambda_{2})$ as presented in (10), and define $\psi:X\to Y$ by
$\psi(x_{i}):=\begin{cases}1&\text{if }i\in\{1,3,4\},\\ 2&\text{if }i\in\{2,6\},\\ 3&\text{if }i\in\{5,7\}.\end{cases}$
Using the relations in system (10), it is straightforward to verify that $\psi$, $\sigma\psi$, and $\sigma^{2}\psi$ are GL-rack homomorphisms from $\mathcal{G}(\Lambda_{2})$ to $R$. Hence, $\operatorname{Col}(\Lambda_{2},R)\geq 3$. (In fact, using a similar method as in the remainder of this proof, one can show that this bound is actually an equality.)
On the other hand, we claim that $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda_{1}),R)=\emptyset$. Using the presentation of $\mathcal{G}(\Lambda_{1})$ in system (9), let $\varphi$ be any GL-rack homomorphism from $\mathcal{G}(\Lambda_{1})$ to $R$. Apply $\varphi$ to system (9). Then the images $y_{i}:=\varphi(x_{i})\in Y$ satisfy the following system of equations in $R$:
(11) $R\begin{cases}\sigma\sigma^{-1}(y_{4})=y_{5}\iff y_{4}=y_{5},\\ \sigma\sigma^{-1}(y_{1})=y_{2}\iff y_{1}=y_{2},\\ \sigma(y_{1})=y_{6},\\ \sigma(y_{3})=\sigma^{-1}(y_{2})\iff y_{3}=\sigma(y_{2}),\\ \sigma(y_{6})=y_{5},\\ \sigma(y_{4})=y_{3}.\end{cases}$
Here, we have used the fact that $\sigma^{3}=\operatorname{id}_{Y}$ to rewrite the fourth equality. We deduce that
$\sigma(y_{2})=y_{3}=\sigma(y_{4})=\sigma(y_{5})=\sigma^{2}(y_{6})=\sigma^{3}(y_{1})=y_{1}=y_{2},$
but $\sigma$ has no fixed points in $Y$. Hence, system (11) has no solutions in $R$, so $\varphi$ cannot exist. ∎
Incidentally, Theorem 7.3 gives a positive answer to a question that Kimura [bi]*Section 4 posed in 2023, as we state below; cf. [bi]*Theorem 4.3.
Corollary 7.4.
GL-rack coloring numbers by nonquandle GL-racks are not generally unable to distinguish nonequivalent Legendrian knots whose classical invariants are identical, even when linearized contact homology and the ruling invariant fail to do so.
In Appendix A.3, we show a version (Proposition A.2) of Corollary 7.4 for Legendrian racks.
7.3.4. Legendrian $8_{10}$ knots
Next, we distinguish the Legendrian $8_{10}$ knots with classical invariants $(\operatorname{tb},\operatorname{rot})=(-8,3)$ in Figure 6. This confirms a conjecture of Bhattacharyya et al. [atlas-site] and Petkova and Schwartz [new-atlas]. [^2] While we and Bhattacharyya et al. [atlas-site] refer to the $8_{10}$ knot by its name in the Rolfsen table, Petkova and Schwartz [new-atlas] call it the $m8a10$ knot. At the time of writing, we are unaware of any other proofs of this conjecture.
Two oriented Legendrian 8 10 knots with equal classical invariants. All cusped strands and uncusped substrands are labeled.
Figure 6. Front projections of the two Legendrian representatives of the topological knot $8_{10}$ with $(\operatorname{tb},\operatorname{rot})=(-8,3)$ given in [atlas-site].
Let $\Lambda_{1}$ and $\Lambda_{2}$ be the Legendrian knots on the left and right of Figure 6, respectively. Using Figure 3 and bi-Legendrian rack axiom 3, we compute that $\mathcal{G}(\Lambda_{1})$ and $\mathcal{G}(\Lambda_{2})$ are the free GL-racks on the set $\{x_{1},\dots,x_{10}\}$ modulo the congruences generated by the following relations:
(12) $\mathcal{G}(\Lambda_{1})\begin{cases}s_{x_{1}}\mathtt{u}(x_{3})=x_{4},&s_{x_{4}}\mathtt{d}\mathtt{u}(x_{1})=x_{2},\\ s_{x_{7}}(x_{5})=\mathtt{d}(x_{4}),&s_{x_{5}}(x_{8})=\mathtt{d}(x_{7}),\\ s_{x_{8}}(x_{6})=\mathtt{d}^{2}(x_{5}),&s_{x_{6}}(x_{9})=\mathtt{d}(x_{6}),\\ s_{x_{6}}(x_{3})=x_{2},&s_{x_{6}}(x_{10})=x_{1},\\ s_{x_{3}}(x_{7})=\mathtt{u}\mathtt{d}^{2}(x_{6}),&s_{x_{3}}(x_{9})=x_{10}.\end{cases}\;\mathcal{G}(\Lambda_{2})\begin{cases}s_{x_{1}}(x_{7})=\mathtt{u}\mathtt{d}(x_{6}),&s_{x_{5}}(x_{2})=x_{3},\\ s_{x_{7}}(x_{2})=\mathtt{d}\mathtt{u}(x_{1}),&s_{x_{3}}\mathtt{u}(x_{5})=x_{6},\\ s_{x_{10}}(x_{4})=\mathtt{d}^{2}(x_{3}),&s_{x_{3}}(x_{10})=x_{9},\\ s_{x_{4}}(x_{1})=\mathtt{d}^{2}(x_{10}),&s_{x_{6}}\mathtt{d}(x_{4})=x_{5},\\ s_{x_{5}}\mathtt{d}^{2}(x_{7})=x_{8},&s_{x_{6}}(x_{9})=x_{8}.\end{cases}$
Theorem 7.5.
The two Legendrian $8_{10}$ knots with $(\operatorname{tb},\operatorname{rot})=(-8,3)$ in Figure 6 are nonequivalent; they are distinguishable using coloring numbers with respect to a permutation GL-rack of order 2.
Proof.
Let $Y:=\{1,2\}$. In cycle notation, let $\sigma\in S_{2}$ be the permutation $(12)$. Let $R$ be the permutation GL-rack $(Y,\sigma,\sigma)_{\mathrm{perm}}$, so that $R$ is the fourth GL-rack in Table B.1. By Proposition 3.11, we have $\mathtt{d}=\operatorname{id}_{Y}$ in $R$. We will show that $\operatorname{Col}(\Lambda_{1},R)>\operatorname{Col}(\Lambda_{2},R)$. To that end, let $X$ denote the underlying set of $\mathcal{G}(\Lambda_{1})$ as presented in system (12), and define $\psi:X\to Y$ by
$\psi(x_{i}):=\begin{cases}1&\text{if }i\in\{1,2,5,8,9\},\\ 2&\text{if }i\in\{3,4,6,7,10\}.\end{cases}$
Using the relations of $\mathcal{G}(\Lambda_{1})$ in (12), it is straightforward to verify that $\psi$ and $\sigma\psi$ are GL-rack homomorphisms from $\mathcal{G}(\Lambda_{1})$ to $R$. Hence, $\operatorname{Col}(\Lambda_{1},R)\geq 2$ (which is actually an equality).
On the other hand, we claim that $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda_{2}),R)=\emptyset$. Using the presentation of $\mathcal{G}(\Lambda_{2})$ in system (12), let $\varphi$ be any GL-rack homomorphism from $\mathcal{G}(\Lambda_{2})$ to $R$. Once again, the elements $y_{i}:=\varphi(x_{i})\in Y$ must satisfy the following system of equations in $R$:
$R\begin{cases}\sigma(y_{7})=\sigma(y_{6}),&\sigma(y_{2})=\sigma(x_{1}),\\ \sigma(y_{2})=y_{3},&\sigma^{2}(y_{5})=y_{6},\\ \sigma(y_{10})=y_{9},&\sigma(y_{4})=y_{3},\\ \sigma(y_{1})=y_{10},&\sigma(y_{4})=y_{5},\\ \sigma(y_{7})=y_{8},&\sigma(y_{9})=y_{8}.\end{cases}$
Since $\sigma^{2}=\operatorname{id}_{Y}$, we can also rewrite the equalities $\sigma(y_{7})=\sigma(x_{6})$, $\sigma^{2}(y_{5})=y_{6}$, and $\sigma(y_{10})=y_{9}$ as $y_{7}=y_{6}$, $y_{5}=y_{6}$, and $y_{10}=\sigma(y_{9})$, respectively. Therefore,
$y_{7}=y_{6}=y_{5}=\sigma(y_{4})=y_{3}=\sigma(y_{2})=\sigma(y_{1})=y_{10}=\sigma(y_{9})=y_{8}=\sigma(y_{7}),$
but $\sigma$ has no fixed points in $Y$. Hence, $\varphi$ cannot exist. ∎
7.3.5. Legendrian $8_{13}$ knots
Finally, we distinguish the Legendrian $8_{13}$ knots with classical invariants $(\operatorname{tb},\operatorname{rot})=(-6,1)$ in Figure 7. This completes the classification of Legendrian $8_{13}$ knots in the extended Legendrian knot atlases of Bhattacharyya et al. [atlas-site] and Petkova and Schwartz [new-atlas]. [^3] Petkova and Schwartz [new-atlas] call the $8_{13}$ knot the $m8a13$ knot. At the time of writing, we are unaware of any other proofs that these Legendrian knots are nonequivalent.
Two oriented Legendrian 8 13 knots with equal classical invariants. All cusped strands and uncusped substrands are labeled.
Figure 7. Front projections of the two Legendrian representatives of the topological knot $8_{13}$ with $(\operatorname{tb},\operatorname{rot})=(-6,1)$ given in [atlas-site].
Let $\Lambda_{1}$ and $\Lambda_{2}$ be the Legendrian knots on the left and right of Figure 6, respectively. Similarly to before, $\mathcal{G}(\Lambda_{1})$ and $\mathcal{G}(\Lambda_{2})$ are respectively the free GL-racks on the sets $\{x_{1},\dots,x_{10}\}$ and $\{x_{1},\dots,x_{9}\}$ modulo the congruences generated by the following relations:
(13) $\mathcal{G}(\Lambda_{1})\begin{cases}s_{x_{1}}(x_{5})=\mathtt{u}\mathtt{d}(x_{4}),&s_{x_{4}}(x_{9})=\mathtt{d}(x_{8}),\\ s_{x_{7}}(x_{2})=\mathtt{u}(x_{1}),&s_{x_{4}}(x_{1})=\mathtt{d}^{2}(x_{10}),\\ s_{x_{2}}(x_{7})=x_{6},&s_{x_{5}}(x_{8})=\mathtt{u}(x_{7}),\\ s_{x_{6}}(x_{3})=x_{2},&s_{x_{5}}(x_{10})=x_{9},\\ s_{x_{5}}(x_{4})=x_{3},&s_{x_{10}}(x_{6})=\mathtt{d}(x_{5}).\end{cases}\;\mathcal{G}(\Lambda_{2})\begin{cases}s_{x_{1}}(x_{7})=\mathtt{u}(x_{6}),&s_{x_{3}}(x_{8})=\mathtt{d}(x_{7}),\\ s_{x_{1}}(x_{5})=x_{4},&s_{x_{9}}(x_{4})=\mathtt{d}(x_{3}),\\ s_{x_{4}}(x_{2})=\mathtt{u}(x_{1}),&s_{x_{9}}(x_{6})=x_{5},\\ s_{x_{2}}(x_{9})=x_{8},&s_{x_{6}}(x_{1})=\mathtt{d}^{2}(x_{9}).\\ s_{x_{8}}(x_{3})=x_{2},\end{cases}$
Theorem 7.6.
The two Legendrian $8_{13}$ knots with $(\operatorname{tb},\operatorname{rot})=(-6,1)$ in Figure 7 are nonequivalent; they are distinguishable using coloring numbers with respect to a permutation GL-rack of order 3.
Proof.
Let $Y=\{1,2,3\}$. In cycle notation, let $\sigma\in S_{3}$ be the permutation $(123)$, and let $R$ be the permutation rack $(Y,\sigma,\operatorname{id}_{Y})_{\mathrm{perm}}$, so that $R$ is the tenth GL-rack listed in Table B.2. By Proposition 3.11, we have $\mathtt{d}=\sigma^{-1}=(132)$ in $R$. Let $X$ be the underlying set of $\mathcal{G}(\Lambda_{1})$ as presented in system (13), and define $\psi:X\to Y$ by
$\psi(x_{i}):=\begin{cases}1&\text{if }i\in\{1,4,8,10\},\\ 2&\text{if }i\in\{3,5,6,7\},\\ 3&\text{if }i\in\{2,6\}.\end{cases}$
It is straightforward to verify that $\varphi$, $\sigma\psi$, and $\sigma^{2}\psi$ are GL-rack homomorphisms from $\mathcal{G}(\Lambda_{1})$ to $R$, so $\operatorname{Col}(\Lambda_{1},R)\geq 3$ (which is actually an equality). On the other hand, an argument similar to the ones in Theorems 7.3 and 7.5 shows that $\varphi(x_{6})=\sigma\varphi(x_{6})$ for all GL-rack homomorphisms $\varphi\in\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda_{2}),R)$; we leave the details to the reader. Since $\sigma$ has no fixed points in $Y$, we conclude that $\operatorname{Col}(\Lambda_{2},R)=0<\operatorname{Col}(\Lambda_{1},R)$, so $\Lambda_{1}$ and $\Lambda_{2}$ are nonequivalent. ∎
8. Directions for future work
In light of our results, we propose questions for further research.
  1. (1)
[…]
  2. (2)
[…]
  3. (3)
    Classify more families of GL-racks using Theorem 4.1.
  4. (4)
    Compute more GL-rack automorphism groups using Proposition 4.13.
  5. (5)
    In light of Proposition 7.2, do there exist Legendrian knots $\Lambda_{1}$ and $\Lambda_{2}$ and a medial GL-rack $M$ such that $\operatorname{Col}(\Lambda_{1},M)=\operatorname{Col}(\Lambda_{2},M)$ but $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda_{1}),M)\not\cong\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda_{2}),M)$? In the spirit of Corollary 7.4, do there exist such $\Lambda_{1}$ and $\Lambda_{2}$ whose classical invariants are identical? A positive answer would show that $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda),M)$ is a proper enhancement of $\operatorname{Col}(\Lambda,M)$ as an invariant of $\Lambda$.
  6. (6)
    Let $\mathbb{F}$ be a field, and let $M$ be a medial GL-rack. In 2023, Elhamdadi et al. [enhancements]*Theorems 4.2 and 5.1 properly enhanced medial quandle-valued invariants of smooth links using $\mathbb{F}$-algebra homomorphisms between quandle rings and colorings of smooth links by idempotents of quandle rings. Do similar proper enhancements of $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda),M)$ exist?
  7. (7)
    In light of Example 6.6, do tensor products of left-distributive racks, medial quandles (cf. [Hom]*Section 8.1), or Latin quandles have any interesting properties?
  8. (8)
    Theorem 6.9 shows that universal-algebraic tensor products make $\mathsf{Rack}$ and $\mathsf{GLR}$ into symmetric magmoidal categories with units, leading us to ask what their unital nuclei are; see [magmoidal]*Section 2.5. In particular, is $\otimes$ associative? If so, then $\otimes$ induces symmetric monoidal structures on $\mathsf{Rack}$ and $\mathsf{GLR}$.
  9. (9)
    Can tensor products of racks or GL-racks be used to define new invariants of smooth links or Legendrian links?
  10. (10)
    Theorem 6.7 implies that ${\mathsf{GLR}_{\mathrm{med}}}$ enriches over itself. In this light, what are the applications of enriched category theory to medial GL-racks and invariants of Legendrian links?
  11. (11)
    Extend the results of Sections 36 and the algorithms in Appendix A to 4-Legendrian racks and 4-Legendrian biracks, which Kimura [leg-cocycle]*Section 4 introduced in 2024.
  12. (12)
    In light of Remark 3.3, can our results about GL-racks be generalized to virtual racks or virtual biracks?
  13. (13)
    In 2015, Cahn and Levi [virtual-leg] introduced virtual Legendrian knots, which are Legendrian knots in the spherical cotangent bundle of a surface equipped with the natural contact structure. Can GL-racks be used to define invariants of virtual Legendrian links (cf. Remark 3.3)?
  14. (14)
    Transverse knots are knots that lie everywhere transverse to the standard contact structure on $\mathbb{R}^{3}$; see [etnyre]*Section 2.4. Can one define rack-theoretic invariants of transverse knots?
  15. (15)
    The fundamental quandle of a smooth link is interpreted topologically as the set of homotopy classes of paths from a basepoint in the link complement to the boundary of the link complement with several restrictions; see, for example, [quandlebook]*p. 125. Is there a similar contact-topological interpretation of the GL-rack of a Legendrian link?
  16. (16)
    Are there formulas for the number of GL-racks and medial GL-racks of a given finite order?
Appendix A Exhaustive search algorithms
The remainder of this article focuses on computational results and approaches to studying and applying GL-racks. In this appendix, we enumerate GL-racks, medial GL-racks, GL-quandles, and medial GL-quandles of orders $n\leq 8$ up to isomorphism and describe the algorithms we used to do so. An implementation of these algorithms in GAP [GAP4] and the raw data we collected are available in the GitHub repository at [my-code].
A.1. Enumeration of small GL-racks
In Table A.1, we enumerate isomorphism classes of GL-racks, medial GL-racks, GL-quandles, and medial GL-quandles up to order $8$. For comparison, we also list the corresponding numbers for classical racks and quandles.
We obtained the numbers $g(n)$ from Algorithm A.1 and $g^{m}(n)$, $g_{q}(n)$, and $g^{m}_{q}(n)$ from Algorithm A.2. Meanwhile, the numbers $r(n)$, $r^{m}(n)$, $r_{q}(n)$, and $r^{m}_{q}(n)$ were originally computed by McCarron [rack-count] in 2010, Vojtěchovský and Yang [library] in 2019, Henderson et al. [quandle-count] in 2006, and Jedlička et al. [medial-quandles] in 2015, respectively. It appears that each of $g(n)$, $g^{m}(n)$, $g_{q}(n)$, and $g^{m}_{q}(n)$ in Table A.1 grows exponentially and at a much faster rate than its counterpart for classical racks.
$n$ 00 $1$ $2$ $3$ $4$ $5$ $6$ $7$ $8$
$g(n)$ $1$ $1$ $4$ $13$ $62$ $308$ $2132$ $17268$ $189373$
$g^{m}(n)$ $1$ $1$ $4$ $13$ $61$ $298$ $2087$ $16941$ $187160$
$g_{q}(n)$ $1$ $1$ $2$ $6$ $19$ $74$ $353$ $2080$ $16023$
$g^{m}_{q}(n)$ $1$ $1$ $2$ $6$ $18$ $68$ $329$ $1965$ $15455$
$r(n)$ $1$ $1$ $2$ $6$ $19$ $74$ $353$ $2080$ $16023$
$r^{m}(n)$ $1$ $1$ $2$ $6$ $18$ $68$ $329$ $1965$ $15455$
$r_{q}(n)$ $1$ $1$ $1$ $3$ $7$ $22$ $73$ $298$ $1581$
$r^{m}_{q}(n)$ $1$ $1$ $1$ $3$ $6$ $18$ $58$ $251$ $1410$
Table A.1. The numbers of GL-racks $g(n)$, medial GL-racks $g^{m}(n)$, GL-quandles $g_{q}(n)$, and medial GL-quandles $g^{m}_{q}(n)$ of orders $0\leq n\leq 8$ up to isomorphism, compared against the corresponding numbers of racks $r(n)$, medial racks $r^{m}(n)$, quandles $r_{q}(n)$, and medial quandles $r^{m}_{q}(n)$.
For explicit representatives of each GL-rack isomorphism class counted in Table A.1, see Appendix B for those of orders $2\leq n\leq 4$ and the GitHub repository in [my-code] for those of orders $5\leq n\leq 8$. The unique GL-rack isomorphism classes of orders 00 and $1$ correspond to the initial and terminal objects in $\mathsf{GLR}$, respectively.
Note in Table A.1 that $g_{q}(n)=r(n)$ and $g^{m}_{q}(n)=r^{m}(n)$ for all $n\leq 8$. This observation was the original motivation for Theorem 5.5, which generalizes these one-to-one correspondences in a natural way.
A.2. Classification of small GL-racks
We now distinguish the exhaustive search algorithms in GAP [GAP4] that we used to compute these isomorphism classes. We build upon the work of Vojtěchovský and Yang [library] in 2019, who classified racks up to order $11$ [library-site].
In what follows, let $\mathcal{R}_{n}$ denote Vojtěchovský and Yang’s list of racks of order $n$. Whenever the underlying set $X=\{1,\dots,n\}$ is clearly established, we write GL-racks $(X,s,\mathtt{u})$ as lists $[s,\mathtt{u},\mathtt{d}]$ containing the elements $s$, $\mathtt{u}$, and $\mathtt{d}=\theta^{-1}\mathtt{u}^{-1}$ as in Proposition 3.11. This is also how we encode GL-racks in our GAP implementation.
Data: List $\mathcal{R}_{n}$ of racks with underlying set $X=\{1,\dots,n\}$ from the library of Vojtěchovský and Yang [library-site] with $1\leq n\leq 11$
Result: List isoClasses of all isomorphism classes of GL-racks of order $n$ with no repeats
begin
[…]
       foreach rack structure $s$ in $\mathcal{R}_{n}$ do
[…]
             foreach permutation $\mathtt{u}_{0}$ in $S_{n}$ do
                   if $\mathtt{u}_{0}$ is a GL-structure on $R:=(X,s)$ then
                         seen $\leftarrow$ false;
                         foreach list $[t,\mathtt{u},\mathtt{d}]$ in isoClasses such that $t=s$ do
                               if $\mathtt{u}_{0}$ and $\mathtt{u}$ are not conjugate in $S_{n}$ then continue;
                               foreach permutation $\varphi$ in $S_{n}\setminus$notHoms do
                                     if $\varphi\mathtt{u}_{0}=\mathtt{u}\varphi$ then
                                           if $\varphi\notin\operatorname{Aut}R$ then Add(notHoms, $\varphi$);
                                           else
                                                 seen $\leftarrow$ true;
                                                 break;
                              if seen then break;
                        if seen $=$ false then Add(isoClasses, $[s,\mathtt{u}_{0},\theta_{R}^{-1}\mathtt{u}_{0}^{-1}]$);
Algorithm A.1 Classification of all GL-racks of a given order $1\leq n\leq 11$ up to isomorphism.
A.2.1. Tabulation of GL-racks
Algorithm A.1 uses $\mathcal{R}_{n}$ to create a list $\mathsf{IsoClasses}$ with exactly one representative of each GL-rack isomorphism class with underlying set $X=\{1,\dots,n\}$. For each rack $R$ obtained from $\mathcal{R}_{n}$, the program searches for GL-structures $\mathtt{u}_{0}$ on $R$ using Definition 3.1.
After finding a GL-structure $\mathtt{u}_{0}$, to ensure that $\mathsf{IsoClasses}$ contains no isomorphic elements, the algorithm uses Definition 3.2 to search for a permutation $\varphi\in S_{n}$ that defines a GL-rack homomorphism (hence an isomorphism) from $(R,\mathtt{u}_{0})$ to any previously encountered GL-rack of the form $(R,\mathtt{u})$ obtained from $\mathsf{IsoClasses}$. By Proposition 3.12 and Theorem 4.1, it suffices to only consider GL-structures $\mathtt{u}$ on $R$ such that $\mathtt{u}_{0}$ and $\mathtt{u}$ are conjugate in $S_{n}$. The last line of the above pseudocode uses Proposition 3.11.
On the author’s personal computer, running our implementation of Algorithm A.1 in GAP with $n=7$ took just under two minutes, while the computation for $n=8$ took seven days. By contrast, a similar program using Definition 3.9 took three hours for the $n=7$ computation. This shows that verifying Definition 3.1 is significantly easier than verifying the bi-Legendrian rack axioms.
Data: List isoClasses of isomorphism classes of GL-racks with underlying set $X=\{1,\dots,n\}$ returned by Algorithm A.1
Result: Lists $\mathcal{Q}_{n}$, $\mathcal{M}_{n}$, and $\mathcal{B}_{n}$ of isomorphism classes of GL-quandles, medial GL-racks, and medial GL-quandles, respectively, whose underlying set is $X$
begin
       $\mathcal{Q}_{n},\mathcal{M}_{n},\mathcal{B}_{n},\textnormal{{seen}}\leftarrow\emptyset$;
       foreach list $R=[s,\mathtt{u},\mathtt{d}]$ in isoClasses do
             if seen contains a list $[t,\textnormal{{isQuandle}},\textnormal{{isMedial}}]$ such that $t=s$ then
                   if isQuandle then Add($\mathcal{Q}_{n}$, $R$);
                   if isMedial then
                         Add($\mathcal{M}_{n}$, $R$);
                         if isQuandle then Add($\mathcal{B}_{n}$, $R$);
             else
                   if $\mathtt{d}=\mathtt{u}^{-1}$ then
                         isQuandle $\leftarrow$ true;
                         Add($\mathcal{Q}_{n}$, $R$);
                   else isQuandle $\leftarrow$ false;
                   isMedial $\leftarrow$ true;
                   foreach ordered triple $(x,y,z)$ in $X^{3}$ do
                         if $s_{s_{x}(z)}s_{y}\neq s_{s_{x}(y)}s_{z}$ then
                               isMedial $\leftarrow$ false;
                               break;
                  if isMedial then
                         Add($\mathcal{M}_{n}$, $R$);
                         if isQuandle then Add($\mathcal{B}_{n}$, $R$);
                  Add(seen,$[s,\textnormal{{isQuandle}},\textnormal{{isMedial}}]$);
Algorithm A.2 Classification of medial GL-racks, all GL-quandles, and medial GL-quandles of order $n$ up to isomorphism, given a classification of GL-racks of order $n$.
A.2.2. Tabulation of medial GL-racks and GL-quandles
Algorithm A.2 tests whether or not each GL-rack in the output of Algorithm A.1 is medial or a GL-quandle. To test for mediality, the algorithm simply verifies equation (1) for all elements $x,y,z\in X$. By Remark 3.10, to test whether a GL-rack $[s,\mathtt{u},\mathtt{d}]$ is a GL-quandle, it suffices to verify whether $\mathtt{d}=\mathtt{u}^{-1}$.
On the author’s personal computer, running our implementation of Algorithm A.2 in GAP with $n=7$ took just under an hour, while the $n=8$ computation took eight days.
A.3. Exhaustive searches for GL-rack coloring numbers
We now describe Algorithm A.3, which computes all colorings of the GL-rack of an oriented Legendrian link $\Lambda$ by each GL-rack in the list $\mathsf{isoClasses}$ computed by Algorithm A.1. Before running Algorithm A.3, the user must input a presentation of $\mathcal{G}(\Lambda)$ in terms of crossing relations between generators of $\mathcal{G}(\Lambda)$; see Section 7.3 for examples of such presentations. By the discussion in Section 7.3, it suffices for the algorithm to search for all valid solutions in $R$ to the inputted crossing relations.
Data: List isoClasses of isomorphism classes of GL-racks with underlying set $Y=\{1,\dots,n\}$ from Algorithm A.1 and a presentation of $\mathcal{G}(\Lambda)=[X,s^{\Lambda},\mathtt{u}^{\Lambda},\mathtt{d}^{\Lambda})]$
Result: List solutions whose elements are lists $[s,\mathtt{u},\mathtt{d},\mathbf{y}]$ such that the mapping $x_{i}\mapsto y_{i}$ defines a GL-rack homomorphism $\mathcal{G}(\Lambda)\to[Y,s,\mathtt{u},\mathtt{d}]$
begin
       $m\leftarrow|X_{\Lambda}|$;
[…]
       foreach GL-rack $[Y,s,\mathtt{u},\mathtt{d}]$ in isoClasses do
             foreach ordered $m$-tuple $\mathbf{y}\leftarrow(y_{1},\dots,y_{m})$ in $Y^{m}$ do
                   if all crossing relations are satisfied after replacing each $x_{i}\in X_{\Lambda}$, $s^{\Lambda}$, $\mathtt{u}^{\Lambda}$, and $\mathtt{d}^{\Lambda}$ with $y_{i}$, $s$, $\mathtt{u}$, and $\mathtt{d}$, respectively then Add(solutions, $[s,\mathtt{u},\mathtt{d},\mathbf{y}]$);
Algorithm A.3 Computation of colorings of an oriented Legendrian link $\Lambda$ by GL-racks of a given order $1\leq n\leq 11$ as computed by Algorithm A.1.
In particular, if $R=(Y,s,\mathtt{u},\mathtt{d})$ is a bi-Legendrian rack of order $n\leq 11$, then $\operatorname{Col}(\Lambda,R)$ is simply the number of lists in $\mathsf{solutions}$ produced by Algorithm A.3 whose first three list elements are $s$, $\mathtt{u}$, and $\mathtt{d}$. To distinguish two oriented Legendrian links $\Lambda_{1}$ and $\Lambda_{2}$, it suffices to run Algorithm A.3 twice, once inputting $\mathcal{G}(\Lambda_{1})$ and again inputting $\mathcal{G}(\Lambda_{2})$, and find a GL-rack $R$ in $\mathsf{isoClasses}$ such that $\operatorname{Col}(\Lambda_{1},R)\neq\operatorname{Col}(\Lambda_{2},R)$. For example, running Algorithm A.3 with $n=2,3$ is how we determined which GL-racks and homomorphisms to employ in Section 7.3. Running the algorithm with $n=5$ also gave us the following example.
Example A.1.
In this example, we use Algorithm A.3 to once again distinguish the Legendrian $6_{2}$ knots $\Lambda_{1}$ and $\Lambda_{2}$ on the left and right of Figure 5, respectively. This time, we use the $222$nd GL-rack $R$ of order $5$ listed in the data linked above, which is a Legendrian rack. Let $Y:=\{1,2,3,4,5\}$. In cycle notation, let $\sigma,\tau\in S_{5}$ be the $5$-cycles $\sigma:=(12345)$ and $\tau:=(13524)$. Let $R:=(Y,\sigma,\tau)_{\mathrm{perm}}$. By Corollary 3.13, $R$ is a Legendrian rack.
We input the relations of $\mathcal{G}(\Lambda_{1})$ in (9) and then those of $\mathcal{G}(\Lambda_{2})$ in (10) into our GAP implentation of Algorithm A.3. After running the program with $n=5$, the program outputs the text in Figure A.1 upon reaching $\mathsf{isoClasses}[222]=R$. The output states that $\operatorname{Col}(\Lambda_{1},R)=0\neq 5=\operatorname{Col}(\Lambda_{2},R)$, and the images of the generators $(x_{1},\dots,x_{7})$ of $\mathcal{G}(\Lambda_{2})$ under each element of $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda_{2}),R)$ are given by the orbit of $(1,2,3,5,1,4,5)\in Y^{7}$ under the action of the subgroup $\langle\sigma\rangle\leq S_{5}$ on $Y^{7}$.
In particular, Example A.1 yields an analogue of Corollary 7.4 for Legendrian racks.
Proposition A.2.
There exist Legendrian knots sharing the same classical invariants that are distinguished by the Legendrian rack coloring numbers originally defined in [ceniceros]*Proposition 1.
Finding all colorings of knot 1 by GL-rack 222 of 308…
Finding all colorings of knot 2 by GL-rack 222 of 308…
[ [(1,2,3,4,5), … (1,2,3,4,5)], (1,3,5,2,4), (1,3,5,2,4), 1, 2, 3, 5, 1, 4, 5 ]
[ [(1,2,3,4,5), … (1,2,3,4,5)], (1,3,5,2,4), (1,3,5,2,4), 2, 3, 4, 1, 2, 5, 1 ]
[ [(1,2,3,4,5), … (1,2,3,4,5)], (1,3,5,2,4), (1,3,5,2,4), 3, 4, 5, 2, 3, 1, 2 ]
[ [(1,2,3,4,5), … (1,2,3,4,5)], (1,3,5,2,4), (1,3,5,2,4), 4, 5, 1, 3, 4, 2, 3 ]
[ [(1,2,3,4,5), … (1,2,3,4,5)], (1,3,5,2,4), (1,3,5,2,4), 5, 1, 2, 4, 5, 3, 4 ]
Number of colorings of knot 1 by GL-rack 222 of 308: 0
Number of colorings of knot 2 by GL-rack 222 of 308: 5
Since their GL-rack coloring numbers are distinct, these knots are not Legendrian isotopic.
Figure A.1. Excerpt from the output of our GAP implementation of Algorithm A.3 with $n=5$. Here, knots $1$ and $2$ are the Legendrian knots in Figure 5, while GL-rack $222$ of $308$ is the Legendrian rack $R$ described in Example A.1.
Appendix B Tabulation of GL-racks of orders 2, 3, and 4
Tables B.1, B.2, and LABEL:tab:tab4 respectively tabulate GL-racks of orders $n=2,3,4$ up to isomorphism, computed using our implementation of Algorithm A.1 in [my-code].
In each table, we write the permutations $s_{i},\mathtt{u},\mathtt{d}$ as either the identity map $\operatorname{id}$ or elements of $S_{n}$ in cycle notation, with permutations composed from right to left.
The number of GL-racks of each order is given by the number of entries in the second column of each table. These entries denote all valid bi-Legendrian structures $[\mathtt{u},\mathtt{d}]$ up to isomorphism on the rack ($(\{1,\dots,n\}\},s)$, where $s$ is given by the corresponding entry in the first column.
For example, the $11$th entry in Table B.2 is the permutation GL-rack of order 3 with $\sigma=(123)$ and bi-Legendrian structure $[\mathtt{u},\mathtt{d}]=[(132),\operatorname{id}]$, which we used to prove Theorem 7.3.
$[s_{1},s_{2}]$ $[\mathtt{u},\mathtt{d}]$ GL-quandle? Medial?
$[\operatorname{id},\operatorname{id}]$
$[\operatorname{id},\operatorname{id}]$,
$[(12),(12)]$
Yes Yes
$[(12),(12)]$
$[\operatorname{id},(12)]$,
$[(12),\operatorname{id}]$
No Yes
Table B.1. The four isomorphism classes of GL-racks of order 2.
$[s_{1},s_{2},s_{3}]$ $[\mathtt{u},\mathtt{d}]$ GL-quandle? Medial?
$[\operatorname{id},\operatorname{id},\operatorname{id}]$
$[\operatorname{id},\operatorname{id}]$,
$[(23),(23)]$,
$[(132),(123)]$
Yes Yes
$[\operatorname{id},(23),(23)]$
$[\operatorname{id},(23)]$,
$[(23),\operatorname{id}]$
No Yes
$[(23),\operatorname{id},\operatorname{id}]$
$[\operatorname{id},\operatorname{id}]$,
$[(23),(23)]$
Yes Yes
$[(23),(23),(23)]$
$[\operatorname{id},(23)]$,
$[(23),\operatorname{id}]$
No Yes
$[(123),(123),(123)]$
$[\operatorname{id},(132)]$,
$[(132),\operatorname{id}]$,
$[(123),(123)]$
No Yes
$[(23),(13),(12)]$ $[\operatorname{id},\operatorname{id}]$ Yes Yes
Table B.2. The 13 isomorphism classes of GL-racks of order 3.
Table B.3. The 62 isomorphism classes of GL-racks of order 4.
$[s_{1},s_{2},s_{3},s_{4}]$ $[\mathtt{u},\mathtt{d}]$ GL-quandle? Medial?
$[\operatorname{id},\operatorname{id},\operatorname{id},\operatorname{id}]$
$[\operatorname{id},\operatorname{id}]$,
$[(34),(34)]$,
$[(243),(234)]$,
$[(1432),(1234)]$,
$[(14)(23),(14)(23)]$
Yes Yes
$[\operatorname{id},(13)(24),\operatorname{id},(13)(24)]$
$[\operatorname{id},(24)]$,
$[(24),\operatorname{id}]$,
$[(13),(13)(24)]$,
$[(13)(24),(13)]$
No Yes
$[(13)(24),(13)(24),(13)(24),(13)(24)]$
$[\operatorname{id},(13)(24)]$,
$[(24),(13)]$,
$[(1432)(1432)]$,
$[(14)(23),(12)(34)]$,
$[(13)(24),\operatorname{id}]$
No Yes
$[\operatorname{id},\operatorname{id},(34),(34)]$
$[\operatorname{id},(34)]$,
$[(34),\operatorname{id}]$,
[(12),(12)(24)[(12),(12)(24)[ ( 12 ) , ( 12 ) ( 24 ),
$[(12)(34),(12)]$
No Yes
$[\operatorname{id},(34),\operatorname{id},\operatorname{id}]$
$[\operatorname{id},\operatorname{id}]$,
$[(34),(34)]$
Yes Yes
$[\operatorname{id},(34),(34),(34)]$
$[\operatorname{id},(34)]$,
$[(34),\operatorname{id}]$
No Yes
$[(34),(34),\operatorname{id},\operatorname{id}]$
$[\operatorname{id},\operatorname{id}]$,
$[(34),(34)]$,
$[(12),(12)]$,
$[(12)(34),(12)(34)]$
Yes Yes
$[(34),(34),(34),(34)]$
$[\operatorname{id},(34)]$,
$[(34),\operatorname{id}]$,
$[(12),(12)(34)]$,
$[(12)(34),(12)]$
No Yes
$[\operatorname{id},(234),(234),(234)]$
$[\operatorname{id},(243)]$,
$[(243),\operatorname{id}]$,
$[(234),(234)]$
No Yes
$[(234),\operatorname{id},\operatorname{id},\operatorname{id}]$
$[\operatorname{id},\operatorname{id}]$,
$[(243),(234)]$,
$[(234),(243)]$
Yes Yes
$[(234),(234),(234),(234)]$
$[\operatorname{id},(243)]$,
$[(243),\operatorname{id}]$,
$[(234),(234)]$
No Yes
$[(234),(243),(243),(243)]$
$[\operatorname{id},(234)]$,
$[(243),(243)]$,
$[(234),\operatorname{id}]$
No Yes
$[(34),(34),(12),(12)]$
$[\operatorname{id},\operatorname{id}]$,
$[(34),(34)]$,
$[(12)(34),(12)(34)]$
Yes Yes
$[(34),(34),(12)(34),(12)(34)]$
$[\operatorname{id},(34)]$,
$[(34),\operatorname{id}]$,
$[(12),(12)(34)]$,
$[(12)(34),(12)]$
No Yes
$[(12),(12),(34),(34)]$
$[\operatorname{id},(12)(34)]$,
$[(34),(12)]$,
$[(12)(34),\operatorname{id}]$
No Yes
$[(12),(12),(12)(34),(12)(34)]$
$[\operatorname{id},(12)(34)]$,
$[(34),(12)]$,
$[(12),(34)]$,
$[(12)(34),\operatorname{id}]$
No Yes
$[(1324),(1324),(1324),(1324)]$
$[\operatorname{id},(1423)]$,
$[(1423),\operatorname{id}]$,
$[(12)(34),(1324)]$,
$[(1324),(12)(34)]$
No Yes
$[\operatorname{id},(34),(24),(23)]$ $[\operatorname{id},\operatorname{id}]$ Yes No
$[(234),(143),(124),(132)]$ $[\operatorname{id},\operatorname{id}]$ Yes Yes

### v3

@@
-and knot coloring invariants
+Classification and structure of generalized Legendrian racks
 Abstract.
-Generalized Legendrian racks are nonassociative algebraic structures based on the Legendrian Reidemeister moves. We study algebraic aspects of GL-racks and coloring invariants of Legendrian links.
+We study algebraic aspects of generalized Legendrian racks, which are nonassociative structures based on the Legendrian Reidemeister moves.
 We answer an open question characterizing the group of GL-structures on a given rack.
 As applications, we classify several infinite families of GL-racks.
-We also compute automorphism groups of dihedral GL-quandles and the categorical center of GL-racks.
-Then we construct an equivalence of categories between racks and GL-quandles.
+We also compute automorphism groups of dihedral GL-quandles.
+Then we compute the centers of the category of GL-racks and several of its full subcategories. We also construct an equivalence of categories between racks and GL-quandles.
 We also study tensor products of racks and GL-racks coming from universal algebra.
@@
 The induced symmetric monoidal structure on medial racks is closed, and similarly for medial GL-racks.
-Answering another open question, we use GL-racks to distinguish Legendrian knots whose classical invariants are identical. In particular, we complete the classification of Legendrian $8_{13}$ knots.
 Finally, we use exhaustive search algorithms to classify GL-racks up to order 8.
 Key words and phrases:
-Algebraic theory, center, coloring invariant, Legendrian knot, medial quandle, rack, tensor product
+Algebraic theory, center, equivalence of categories, Legendrian knot invariant, medial, quandle, rack, tensor product
 2020 Mathematics Subject Classification:
-Primary 08A35, 57K12; Secondary 18B99, 20B25, 57K10, 57K33
+Primary 20N02; Secondary 08A35, 18B99, 20B25, 57K12
 1. Introduction
 Generalized Legendrian racks, also called GL-racks or bi-Legendrian racks, are a nonassociative algebraic structure used to distinguish Legendrian links in $\mathbb{R}^{3}$ or $S^{3}$.
-GL-racks can be traced back to algebraic structures called kei, which Takasaki [takasaki] introduced in 1942 to study symmetric spaces; quandles, which Joyce [joyce] introduced in 1982 to study conjugation in groups and links in $\mathbb{R}^{3}$ and $S^{3}$; and racks, which Fenn and Rourke [fenn] introduced in 1992 to study framed links in $3$-manifolds.
+GL-racks can be traced back to algebraic structures called kei, which Takasaki [takasaki] introduced in 1942 to study symmetric spaces; quandles, which Joyce [joyce] and Matveev [matveev] independently introduced in 1982 to study links in $\mathbb{R}^{3}$ and $S^{3}$; and racks, which Fenn and Rourke [fenn] introduced in 1992 to study framed links in $3$-manifolds.
 Kei, quandles, and racks have enjoyed significant study as link invariants in geometric topology and in their own rights in quantum algebra and group theory.
@@
 [karmakar] and Kimura [bi] independently strengthened these constructions by introducing GL-racks, which are racks equipped with additional structure.
-In this paper, we further develop the theory and applications of GL-racks, answer open questions about GL-racks and Legendrian knot classifications, and classify GL-racks up to order $8$.
-In particular, we provide the first examples of GL-racks that distinguish Legendrian knots that are not distinguishable by their classical invariants, answering a question of Kimura [bi]*Section 4.
+After completing the initial preprint of this paper, the author [ta] also studied involutory GL-racks in relation to similar algebraic structures used to study surface-knots.
+In this paper, we further develop the theory and classification of GL-racks. In particular, we study categorical and group-theoretic aspects of GL-racks, and we classify GL-racks up to order $8$.
 1.1. The structure of this paper
@@
 In Section 3, we give a simplified definition of GL-racks, show its equivalence to the definition in the literature, and discuss examples and universal-algebraic aspects of GL-racks.
-Sections 46 study GL-racks from group-theoretic, categorical, and universal-algebraic perspectives, respectively; readers only interested in applications to Legendrian knot theory may safely skip these sections.
-In Section 4, we determine all GL-structures on a given rack, answering a question of Karmakar et al.
-[karmakar]*Section 3.
-As applications, we classify all GL-structures on permutation racks, conjugation quandles of abelian and centerless groups, Takasaki kei of abelian $2$-torsion-free groups, and dihedral quandles.
-We also characterize GL-rack automorphism groups and compute them for dihedral quandles.
-Finally, we compute the centers of the categories of GL-racks, GL-quandles, Legendrian racks, and Legendrian quandles.
-In Section 5, we construct an equivalence of categories between the categories of racks and GL-quandles.
+In Section 4, we study group-theoretic aspects of GL-racks.
+We determine all GL-structures on a given rack, answering a question of Karmakar et al.
+[karmakar]*Sec.
+3.
+As applications, we classify all GL-structures on permutation racks, certain conjugation quandles, and certain Takasaki kei, including all dihedral quandles.
+We also characterize GL-rack automorphism groups and compute them for dihedral quandles equipped with GL-structures.
+In Section 5, we study categorical aspects of GL-racks.
+First, we compute the centers of the categories of GL-racks, GL-quandles, Legendrian racks, and Legendrian quandles.
+Then we construct an equivalence of categories between the categories of racks and GL-quandles.
 This equivalence restricts to an equivalence between the full subcategories of medial racks and medial GL-quandles.
@@
 This is unusual for noncommutative algebraic theories; for example, the tensor product in the category of groups does not have a tensor unit.
-We show that the induced symmetric monoidal structure on medial racks is closed, and similarly for medial GL-racks.
-In Section 7, we discuss the applications of GL-racks to Legendrian knot theory.
-We use GL-rack coloring invariants to distinguish several conjecturally distinct Legendrian knots with the same classical invariants.
-This gives a positive answer to a question of Kimura [bi]*Section 4; we also show in Appendix A.3 that Legendrian racks give a positive answer.
-In particular, we add to the classification of Legendrian $8_{10}$ knots and complete the classification of Legendrian $8_{13}$ knots in the extended Legendrian knot atlases of Bhattacharyya et al.
-[atlas-site] and Petkova and Schwartz [new-atlas].
-In Section 8, we propose questions for further research based on our results.
+The induced symmetric monoidal structure on medial racks is closed, and similarly for medial GL-racks.
+In Section 7, we propose questions for further research.
 In Appendix A, we describe algorithms that can classify finite GL-racks of a given order up to isomorphism.
@@
 We provide implementations of our algorithms in GAP [GAP4] and the data we were able to compute and enumerate for all $n\leq 8$.
-We also provide an algorithm that computes GL-rack coloring invariants of Legendrian links.
-As an application, we show that Legendrian racks positively answer the question of Kimura [bi]*Section 4 answered in Section 7 for GL-racks.
 In Appendix B, we use the algorithms in Appendix A.2 to tabulate all GL-racks of orders $2\leq n\leq 4$ up to isomorphism.
@@
 Acknowledgments.
-I am indebted to Sam Raskin, my undergraduate thesis adviser, for his mentorship, support, and many insights throughout the research and writing processes.
-I thank Samantha Pezzimenti and Wing Hong Tony Wong for introducing me to Legendrian knot theory and mathematics research, Jose Ceniceros for informing me of the categorical literature on racks, and Peyton Wood for introducing me to GL-racks.
-I also thank Noam Scully for helpful discussions, Qiaochu Yuan for a terminological correction, and an anonymous Stack Exchange user for referencing me to [alg-theories]*Theorem 3.10.3.
-This paper was written in partial fulfillment of the requirements for the bachelor’s degree in mathematics at Yale.
-I thank Patrick Devlin, Miki Havlíčková, Kati Hubley, Matthew King, Andrew Neitzke, and Carol Rutschman for inspiring me to pursue and complete this degree.
+I thank Sam Raskin for providing mentorship, support, and insights throughout the research and writing processes.
+I thank Samantha Pezzimenti, Tony W.
+H.
+Wong, Jose Ceniceros, and Peyton Wood for introducing me to various aspects of the theory, and I thank Noam Scully for helpful discussions.
+I also thank Zhiyi He for identifying errors in an earlier draft; these errors were entirely my own.
+The affected section has been omitted.
+I thank Qiaochu Yuan for a terminological correction and an anonymous Stack Exchange user for referencing me to [alg-theories].
+This research was conducted in partial fulfillment of the senior requirements for the bachelor’s degree in mathematics at Yale University.
+I thank Patrick Devlin, Miki Havlíčková, Kati Hubley, Matthew King, Andrew Neitzke, and Carol Rutschman for inspiring me to complete this degree.
 2. Racks
@@
 Finally, if $Y\subseteq X$ and $s_{y}^{\pm 1}(z)\in Y$ for all $y,z\in Y$, then we say that $(Y,s|_{Y})$ is a subrack of $(X,s)$.
-Example 2.2.
-[quandlebook]*Example 99 Let $X$ be a set, and fix $\sigma\in S_{X}$.
+Example 2.2 ([quandlebook]*Ex. 99).
+Let $X$ be a set, and fix $\sigma\in S_{X}$.
 Define $s:X\rightarrow S_{X}$ by $x\mapsto\sigma$, so that $s_{x}(y)=\sigma(y)$ for all $x,y\in X$.
@@
 (Our notation $(X,\sigma)_{\mathrm{perm}}$ is nonstandard, unlike the notation in the next three examples.) Note that $(X,\sigma)_{\mathrm{perm}}$ is a quandle if and only if $\sigma=\operatorname{id}_{X}$, in which case we call $(X,\operatorname{id}_{X})_{\mathrm{perm}}$ a trivial quandle.
-Example 2.3.
-[quandlebook]*Example 54 Let $A$ be an abelian additive group.
+Example 2.3 ([quandlebook]*Ex. 54).
+Let $A$ be an abelian additive group.
 Define $s:A\to S_{A}$ by $b\mapsto s_{b}$ with $s_{b}(a):=2b-a$ for all elements $a,b\in A$.
@@
 Finally, the inner automorphism group of $R$ is the subgroup $\operatorname{Inn}R:=\langle s_{x}\mid x\in X\rangle$ of $\operatorname{Aut}R$.
-Example 2.5.
-[book]*Example 2.13 Let $X$ be a union of conjugacy classes in a group $G$, and define $c^{G}:X\rightarrow S_{X}$ by sending $x$ to the conjugation map $c^{G}_{x}:=[y\mapsto xyx^{-1}]$.
+Example 2.5 ([book]*Ex. 2.13).
+Let $X$ be a union of conjugacy classes in a group $G$, and define $c^{G}:X\rightarrow S_{X}$ by sending $x$ to the conjugation map $c^{G}_{x}:=[y\mapsto xyx^{-1}]$.
 Then $\operatorname{Conj}X:=(X,c^{G})$ is a quandle called a conjugation quandle or conjugacy quandle.
@@
 for all $x,y\in X$.
-Example 2.7.
-[quandlebook]*Example 66 Let $n\geq 3$ be an integer.
+Example 2.7 ([quandlebook]*Ex. 66).
+Let $n\geq 3$ be an integer.
 We call the Takasaki kei $R_{n}:=T(\mathbb{Z}/n\mathbb{Z})$ a dihedral quandle due to the following observation: if $\Sigma=\{s,rs,\dots,r^{n-1}s\}$ is the set of all reflections in the dihedral group $D_{n}=\langle r,s\mid r^{n}=s^{2}=1,\,srs=r^{-1}\rangle$ of order $2n$, then the map $\varphi:\mathbb{Z}/n\mathbb{Z}\to\Sigma$ defined by $k\mapsto r^{k}s$ is a rack isomorphism from $R_{n}$ to $\operatorname{Conj}\Sigma$.
@@
 Every rack has a canonical dual rack, which we define as follows.
-Definition 2.8.
-[dual]*Definition 2.6 Given a rack $R=(X,s)$, define $s^{\prime}:X\to S_{X}$ by $x\mapsto s_{x}^{-1}$.
+Definition 2.8 ([dual]*Def. 2.6).
+Given a rack $R=(X,s)$, define $s^{\prime}:X\to S_{X}$ by $x\mapsto s_{x}^{-1}$.
 Then $R^{\operatorname{op}}:=(X,s^{\prime})$ is a rack called the dual rack of $R$.
@@
 From the perspective of universal algebra, racks are an algebraic theory with two binary operations $s_{-}(-)$ and $s^{-1}_{-}(-)$, and $\mathsf{Rack}$ is the category of models of this algebraic theory in $\mathsf{Set}$, the category of sets and set maps.
-Thus, $\mathsf{Rack}$ is complete and cocomplete (see [alg-theories]*Theorem 3.4.5), as we will need later.
+Thus, $\mathsf{Rack}$ is complete and cocomplete (see [alg-theories]*Thm.
+3.4.5), as we will need later.
 In particular, the Cartesian product of racks has a natural rack structure, so we can define the following.
-Definition 2.11.
-[rack-roll]*Section 3 A rack $(X,s)$ is called medial or abelian if the map $X\times X\to X$ defined by $(x,y)\mapsto s_{y}(x)$ is a rack homomorphism.
-Equivalently, the subgroup $\langle s_{x}s_{y}^{-1}\mid x,y\in X\rangle$ of $\operatorname{Inn}X$ is abelian; see, for example, [medial-quandles]*Proposition 2.1.
+Definition 2.11 ([rack-roll]*Sec. 3).
+A rack $(X,s)$ is called medial or abelian if the map $X\times X\to X$ defined by $(x,y)\mapsto s_{y}(x)$ is a rack homomorphism.
+Equivalently, the subgroup $\langle s_{x}s_{y}^{-1}\mid x,y\in X\rangle$ of $\operatorname{Inn}X$ is abelian; see, for example, [medial-quandles]*Prop.
+2.1.
 This subgroup is called the transvection group or displacement group of $(X,s)$.
@@
 2.2.1. Definition of $\theta_{R}$
-Given a rack $R=(X,s)$, define $\theta_{R}:X\to X$ by ${x\mapsto s_{x}(x)}$; see [center]*Proposition 2.5.
+Given a rack $R=(X,s)$, define $\theta_{R}:X\to X$ by ${x\mapsto s_{x}(x)}$; see [center]*Prop. 2.5.
 Evidently, $R$ is a quandle if and only if $\theta_{R}=\operatorname{id}_{X}$.
@@
 so $\theta$ is a rack endomorphism. Combined with claims 1 and 2, this proves claim 4. ∎
-Hence, $\Theta$ generates a cyclic subgroup of $Z(\mathsf{Rack})$. In fact, this subgroup is $Z(\mathsf{Rack})$; see [center]*Theorem 5.4.
+Hence, $\Theta$ generates a cyclic subgroup of $Z(\mathsf{Rack})$. In fact, this subgroup is $Z(\mathsf{Rack})$; see [center]*Thm. 5.4.
 The following identities will be useful later on.
@@
 Recall that $\mathsf{Rack}$ is the category of models of the algebraic theory of racks in $\mathsf{Set}$, so $\mathsf{Rack}$ is complete and cocomplete.
-Thus, we can consider quotients of racks by congruences; see [alg-theories]*Lemma 3.5.1 and cf.
-[dual]*Proposition 3.6.
+Thus, we can consider quotients of racks by congruences; see [alg-theories]*Lem.
+3.5.1 and cf.
+[dual]*Prop.
+3.6.
 This allows us to state the following definitions, which are actually left adjoints of the inclusion functors from $\mathsf{Qnd}$ and $\mathsf{Rack}_{\mathrm{med}}$ into $\mathsf{Rack}$.
-Definition 2.18.
-[fenn]*Section 2, [joyce]*Section 10 Given a rack $R=(X,s)$, the associated quandle of $R$ is the quandle $R_{\mathrm{qnd}}:=R/{\sim}$, where $\sim$ is the smallest congruence on $X$ such that $s_{x}(x)=x$ for all $x\in X$.
+Definition 2.18 ([fenn]*Sec. 2, [joyce]*Sec. 10).
+Given a rack $R=(X,s)$, the associated quandle of $R$ is the quandle $R_{\mathrm{qnd}}:=R/{\sim}$, where $\sim$ is the smallest congruence on $X$ such that $s_{x}(x)=x$ for all $x\in X$.
 Similarly, the medialization or abelianization of $R$ is the medial rack $R_{\mathrm{med}}:=R/{\sim}$, where $\sim$ is the smallest congruence on $X$ such that
@@
 2.3.2. Free racks
-In analogy to free groups or free modules, the free rack $\operatorname{Free}_{\mathsf{Rack}}(X)$ on a set $X$ exists and is uniquely (up to isomorphism) characterized by a universal property; see, for example, [alg-theories]*Corollary 3.7.8.
+In analogy to free groups or free modules, the free rack $\operatorname{Free}_{\mathsf{Rack}}(X)$ on a set $X$ exists and is uniquely (up to isomorphism) characterized by a universal property; see, for example, [alg-theories]*Cor.
+3.7.8.
 Namely, there exists a set map $i:X\to\operatorname{Free}_{\mathsf{Rack}}(X)$ such that for all racks $R=(Y,t)$ and all set maps $\varphi_{0}:X\to Y$, there exists a unique rack homomorphism ${\varphi\in\operatorname{Hom}_{\mathsf{Rack}}(\operatorname{Free}_{\mathsf{Rack}}(X),R)}$ such that $\varphi i=\varphi_{0}$.
@@
 ${X}$${{\operatorname{Free}_{\mathsf{Rack}}(X)}}$${Y}$$\scriptstyle{i}$$\scriptstyle{\varphi_{0}}$$\scriptstyle{\varphi}$
-For an explicit set-theoretic construction of free racks, see [free-z]*Proposition 1.3.
+For an explicit set-theoretic construction of free racks, see [free-z]*Prop.
+1.3.
 We can also define the free quandle on $X$ to be $(\operatorname{Free}_{\mathsf{Rack}}(X))_{\mathrm{qnd}}$; cf.
-[fenn]*Section 7.1.
-Example 2.19.
-[free-z]*Example 1.6 The free rack $F$ on one generator is canonically isomorphic to the permutation rack $(\mathbb{Z},\sigma)_{\mathrm{perm}}$, where $\sigma(k)=k+1$ for all $k\in\mathbb{Z}$.
+[fenn]*Sec.
+7.1.
+Example 2.19 ([free-z]*Ex. 1.6).
+The free rack $F$ on one generator is canonically isomorphic to the permutation rack $(\mathbb{Z},\sigma)_{\mathrm{perm}}$, where $\sigma(k)=k+1$ for all $k\in\mathbb{Z}$.
 Under this identification, $\theta_{F}=\sigma$.
@@
 Medial racks are the largest commutative subtheory of racks.
-Recall that an algebraic theory $\mathcal{T}$ is called commutative if, for all $\mathcal{T}$-models $A$, every $n$-ary operation $\alpha:\mathcal{T}^{n}\to\mathcal{T}$ defines a homomorphism $A^{n}\to A$; see [alg-theories]*Theorem 3.10.3.
+Recall that an algebraic theory $\mathcal{T}$ is called commutative if, for all $\mathcal{T}$-models $A$, every $n$-ary operation $\alpha:\mathcal{T}^{n}\to\mathcal{T}$ defines a homomorphism $A^{n}\to A$; see [alg-theories]*Thm.
+3.10.3.
 Let $\mathcal{T}_{\operatorname{ab}}$ be the largest commutative subtheory of an algebraic theory $\mathcal{T}$.
 If $\mathcal{C}$ is the category of $\mathcal{T}$-models in $\mathsf{Set}$, let $\mathcal{C}_{\operatorname{ab}}$ be the full subcategory of $\mathcal{C}$ whose objects are $\mathcal{T}_{\operatorname{ab}}$-models.
-The following strengthens part of [alg-theories]*Theorem 3.10.3; we could not find a reference for this precise statement.
+The following strengthens part of [alg-theories]*Thm.
+3.10.3; we could not find a reference for this precise statement.
 Proposition 2.20.
@@
 Proof.
-We refer the reader to the proof of the statement in [alg-theories]*Theorem 3.10.3 that (1) implies (3).
+We refer the reader to the proof of the statement in [alg-theories]*Thm.
+3.10.3 that (1) implies (3).
 This proof uses the commutativity of $n$-ary operations in $Y$ but not in $X$, so the implication still holds with the weakened assumption that $X$ is not necessarily a $\mathcal{T}_{\operatorname{ab}}$-model.
@@
 That is, if $\mathcal{C}=\mathsf{Rack}$, then $\mathcal{C}_{\operatorname{ab}}=\mathsf{Rack}_{\mathrm{med}}$; this follows directly from Definition 2.11 and Lemma 2.15.
-Hence, Proposition 2.20 recovers the following result of Grøsfjeld [rack-roll]*Proposition 3.3 in 2021, which in turn generalized a result of Crans and Nelson [Hom]*Theorem 3 in 2014.
+Hence, Proposition 2.20 recovers the following result of Grøsfjeld [rack-roll]*Prop. 3.3 in 2021, which in turn generalized a result of Crans and Nelson [Hom]*Thm. 3 in 2014.
 Corollary 2.22.
@@
 Remark 3.3.
-Virtual racks are algebraic structures that can distinguish framed links in certain lens spaces and framed virtual links in thickened surfaces; see, for example, [virtual]*Section 3.2.
+Virtual racks are algebraic structures that can distinguish framed links in certain lens spaces and framed virtual links in thickened surfaces; see, for example, [virtual]*Sec. 3.2.
 By Definition 3.1, GL-racks are precisely virtual racks in which all inner automorphisms $s_{x}$ are endomorphisms of virtual racks.
-Equivalently, a GL-rack $(X,s,\mathtt{u})$ is a virtual rack in which the operator group of $(X,s)$ identifies $x$ with $\mathtt{u}(x)$ for all $x\in X$; see [fenn]*Section 1.1.
+Equivalently, a GL-rack $(X,s,\mathtt{u})$ is a virtual rack in which the operator group of $(X,s)$ identifies $x$ with $\mathtt{u}(x)$ for all $x\in X$; see [fenn]*Sec.
+1.1.
 […]
@@
 For all racks $R=(X,s)$, the identity map $\operatorname{id}_{X}$ is a GL-structure on $R$. Consequently, every rack can be equipped with at least one GL-structure.
-Example 3.5.
-[bi]*Example 3.7 Given a permutation rack $P=(X,\sigma)_{\mathrm{perm}}$, a GL-structure on $P$ is precisely a permutation $\mathtt{u}\in S_{X}$ such that $\mathtt{u}\sigma=\sigma\mathtt{u}$.
+Example 3.5 ([bi]*Ex. 3.7).
+Given a permutation rack $P=(X,\sigma)_{\mathrm{perm}}$, a GL-structure on $P$ is precisely a permutation $\mathtt{u}\in S_{X}$ such that $\mathtt{u}\sigma=\sigma\mathtt{u}$.
 Given such a $\mathtt{u}$, we say that $(P,\mathtt{u})$ is a permutation GL-rack or constant action GL-rack, and we denote it by $(X,\sigma,\mathtt{u})_{\mathrm{perm}}$.
-Example 3.6.
-[bi]*Example 3.6 Let $G$ be a group, let $z\in Z(G)$ be a central element of $G$, and define $f:G\rightarrow G$ by $g\mapsto zg$. Then $(\operatorname{Conj}G,f)$ is a GL-quandle.
+Example 3.6 ([bi]*Ex. 3.6).
+Let $G$ be a group, let $z\in Z(G)$ be a central element of $G$, and define $f:G\rightarrow G$ by $g\mapsto zg$. Then $(\operatorname{Conj}G,f)$ is a GL-quandle.
 Example 3.7.
@@
 3.1.1. Bi-Legendrian racks
-Next, we reproduce the definition of GL-racks given in the literature; see [karmakar]*Definition 3.1 or [bi]*Definition 3.2.
+Next, we reproduce the definition of GL-racks given in the literature; see [karmakar]*Def.
+3.1 or [bi]*Def.
+3.2.
 Following Kimura [bi], we temporarily use the term bi-Legendrian racks to distinguish them from GL-racks in the sense of Definition 3.1.
@@
 [ceniceros] introduced in 2021.
-Definition 3.9.
-[karmakar]*Definition 3.1 Given a rack $R=(X,s)$, a bi-Legendrian structure on $R$ is a pair $(\mathtt{u},\mathtt{d})$ of maps $\mathtt{u},\mathtt{d}:X\to X$ satisfying the following axioms for all elements $x\in X$:
+Definition 3.9 ([karmakar]*Def. 3.1).
+Given a rack $R=(X,s)$, a bi-Legendrian structure on $R$ is a pair $(\mathtt{u},\mathtt{d})$ of maps $\mathtt{u},\mathtt{d}:X\to X$ satisfying the following axioms for all elements $x\in X$:
   1. (L1)
@@
 The previous two propositions make computing and classifying GL-racks significantly easier; see Section 4 and Appendix A.2.
-As a bonus, they also yield the following converse of [ceniceros]*Remark 2.
+As a bonus, they also yield the following converse of [ceniceros]*Rem.
+2.
 Corollary 3.13.
@@
 3.3.2. Free GL-racks
-As with racks, the free GL-rack $\operatorname{Free}_{\mathsf{GLR}}(X)$ on a set $X$ exists and is uniquely (up to isomorphism) characterized by a universal property; see [karmakar]*Proposition 4.2 for a detailed derivation.
+As with racks, the free GL-rack $\operatorname{Free}_{\mathsf{GLR}}(X)$ on a set $X$ exists and is uniquely (up to isomorphism) characterized by a universal property; see [karmakar]*Prop.
+4.2 for a detailed derivation.
 Namely, there exists a set map $i:X\to\operatorname{Free}_{\mathsf{GLR}}(X)$ such that for all GL-racks $R=(Y,t,\mathtt{u})$ and all set maps $\varphi_{0}:X\to Y$, there exists a unique GL-rack homomorphism $\varphi\in\operatorname{Hom}_{\mathsf{GLR}}(\operatorname{Free}_{\mathsf{GLR}}(X),R)$ such that $\varphi i=\varphi_{0}$.
@@
 Karmakar et al.
-[karmakar]*Section 4 gave an explicit set-theoretic construction of $\operatorname{Free}_{\mathsf{GLR}}(X)$.
+[karmakar]*Sec.
+4 gave an explicit set-theoretic construction of $\operatorname{Free}_{\mathsf{GLR}}(X)$.
 Here, we simplify their construction using Definition 3.1 and and Proposition 3.12.
-Definition 3.15.
-[karmakar]*Section 4 Let $X$ be a set.
+Definition 3.15 ([karmakar]*Sec. 4).
+Let $X$ be a set.
 We define the free GL-rack on $X$ as follows.
@@
 3.3.3. Free objects in subtheories of GL-racks
-Since $\mathsf{GLR}$ is the category of models of the algebraic theory of GL-racks in $\mathsf{Set}$, the free GL-rack $L=\operatorname{Free}_{\mathsf{GLR}}(X)$ on the one-element set $X=\{x\}$ is a strong generator or separator for $\mathsf{GLR}$; see, for example, [alg-theories]*Proposition 3.3.3.
+Since $\mathsf{GLR}$ is the category of models of the algebraic theory of GL-racks in $\mathsf{Set}$, the free GL-rack $L=\operatorname{Free}_{\mathsf{GLR}}(X)$ on the one-element set $X=\{x\}$ is a strong generator or separator for $\mathsf{GLR}$; see, for example, [alg-theories]*Prop.
+3.3.3.
 Since GL-quandles, Legendrian racks, and Legendrian quandles are subtheories of the algebraic theory of GL-racks, we can similarly consider free objects that strongly generate the corresponding subcategories of models by taking quotients of $L$ by congruences.
@@
 ∎
-4. Classification of GL-structures, automorphisms, and centers
+[…]
+Corollary 3.21.
+Let $\mathcal{G}(\Lambda)$ be the fundamental GL-rack of a Legendrian link $\Lambda$ (see [karmakar]).
+Then for all medial GL-racks $M$, the isomorphism type of $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda),M)$ as a medial GL-rack is an invariant of $\Lambda$.
+4. Group-theoretic aspects of GL-racks
 In 2024, Karmakar et al.
-[karmakar]*Section 3 posed the following question: what are all the possible GL-structures on a given rack?
+[karmakar]*Sec.
+3 posed the following question: what are all the possible GL-structures on a given rack?
 In this section, we answer this question and classify various infinite families of GL-racks.
 As applications, we discuss automorphism groups of GL-racks and compute these groups for all GL-racks whose underlying racks are dihedral quandles.
-We also compute the centers of $\mathsf{GLR}$, $\mathsf{GLQ}$, and the categories of Legendrian racks and Legendrian quandles.
 4.1. Characterizing GL-structures
@@
 Next, recall a result of Elhamdadi et al.
-[dihedral]*Theorem 2.3 that $\operatorname{Inn}Q=\operatorname{Inn}_{\mathsf{Grp}}G$ for all groups $G$.
+[dihedral]*Thm.
+2.3 that $\operatorname{Inn}Q=\operatorname{Inn}_{\mathsf{Grp}}G$ for all groups $G$.
 Also, recall a result of Bardakov et al.
-[centerless]*Corollary 2 that $\operatorname{Aut}Q=\operatorname{Aut}_{\mathsf{Grp}}G$ if and only if $G$ is centerless.
+[centerless]*Cor.
+2 that $\operatorname{Aut}Q=\operatorname{Aut}_{\mathsf{Grp}}G$ if and only if $G$ is centerless.
 In this case, Theorem 4.1 states that
@@
 Next, we classify GL-structures on a certain family of Takasaki kei and, as a consequence, all dihedral quandles of odd order.
-We first recall a classification result of Bardakov et al. [takasaki-aut]*Theorem 4.2. For all abelian additive groups $A$ without $2$-torsion, $\operatorname{Aut}T(A)$ is isomorphic to the holomorph
+We first recall a classification result of Bardakov et al. [takasaki-aut]*Thm. 4.2. For all abelian additive groups $A$ without $2$-torsion, $\operatorname{Aut}T(A)$ is isomorphic to the holomorph
 $G:=A\rtimes\operatorname{Aut}_{\mathsf{Grp}}A$
@@
 To show this, we will complete the classification of GL-structures on dihedral quandles $R_{n}$ using the following results of Elhamdadi et al.
-[dihedral]*Theorems 2.1 and 2.2.
+[dihedral]*Thms.
+2.1 and 2.2.
 The automorphism group $\operatorname{Aut}R_{n}$ is the affine group of $\mathbb{Z}/n\mathbb{Z}$.
@@
 Once again, we use a result of Elhamdadi et al.
-[dihedral]*Theorem 2.1 to identify $\operatorname{Aut}R_{n}\cong\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$.
+[dihedral]*Thm.
+2.1 to identify $\operatorname{Aut}R_{n}\cong\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$.
 Proposition 4.15.
@@
 ∎
-4.4. The center of the category of GL-racks
+5. Categorical aspects of GL-racks
+In this section, we compute the centers of $\mathsf{GLR}$ and several of its full subcategories. We also prove a surprising equivalence of categories between $\mathsf{Rack}$ and $\mathsf{GLQ}$.
+5.1. Categorical centers
 Recall that the center of a category $\mathcal{C}$ is the commutative monoid $Z(\mathcal{C})$ of natural endomorphisms of the identity functor $\mathbf{1}_{\mathcal{C}}$.
-In 2018, Szymik [center]*Theorems 5.4 and 5.5 computed that $Z(\mathsf{Rack})=\langle\Theta\rangle\cong\mathbb{Z}$ and $Z(\mathsf{Qnd})\cong\{1\}$.
-In this subsection, we similarly compute the centers of $\mathsf{GLR}$ and various full subcategories of $\mathsf{GLR}$.
+In 2018, Szymik [center]*Thms.
+5.4 and 5.5 computed that $Z(\mathsf{Rack})=\langle\Theta\rangle\cong\mathbb{Z}$ and $Z(\mathsf{Qnd})\cong\{1\}$.
+In this section, we similarly compute the centers of $\mathsf{GLR}$, $\mathsf{GLQ}$, and the categories of Legendrian racks and Legendrian quandles.
 Theorem 5.1.
@@
 This proves claim 2. Similarly, claim 3 follows from Corollary 3.13 and Proposition 3.18. ∎
-5. Categorical equivalence of racks and GL-quandles
-In this section, we show that the categories of racks and GL-quandles are isomorphic in a way that preserves mediality.
-This surprising result generalizes the one-to-one correspondences observed in Appendix A.1 and induces isomorphisms of the respective algebraic theories.
-5.1. Construction of $F$
+5.2. Categorical equivalence of racks and GL-quandles
+Next, we show that the categories of racks and GL-quandles are isomorphic in a way that preserves mediality.
+This surprising result generalizes the one-to-one correspondences observed in Appendix A.1 and corresponds to isomorphisms of the respective algebraic theories.
+5.2.1. Construction of $F$
 We begin by defining a functor $F:\mathsf{Rack}\to\mathsf{GLQ}$. First, we define how $F$ acts on objects.
@@
 ∎
-5.2. Construction of $G$
+5.2.2. Construction of $G$
 We now define a functor $G:\mathsf{GLQ}\to\mathsf{Rack}$ as the restriction of a functor $\widetilde{G}:\mathsf{GLR}\to\mathsf{Rack}$ to $\mathsf{GLQ}$.
@@
 for all $x\in X$, as desired. ∎
-5.3. Isomorphism of categories
+5.2.3. Isomorphism of categories
 Having defined $F$ and $G$, we are now ready to prove the main results of this section.
@@
 Since $F$ and $G$ are left and right adjoints to each other, they both preserve limits and colimits.
-In particular, they preserve finite products and filtered colimits, so by [alg-theories]*Lemma 3.8.3, $F$ and $G$ are algebraic functors.
-Since $F$ and $G$ are equivalences of categories, the claim follows directly from [alg-theories]*Proposition 3.12.1.
+In particular, they preserve finite products and filtered colimits, so by [alg-theories]*Lem.
+3.8.3, $F$ and $G$ are algebraic functors.
+Since $F$ and $G$ are equivalences of categories, the claim follows directly from [alg-theories]*Prop.
+3.12.1.
 ∎
 6. Tensor products of racks and GL-racks
-In 2014, Crans and Nelson [Hom]*Section 8.1 categorified the results of Corollary 2.22 by considering universal-algebraic tensor products of medial quandles.
+In 2014, Crans and Nelson [Hom]*Sec.
+8.1 categorified the results of Corollary 2.22 by considering universal-algebraic tensor products of medial quandles.
 However, these tensors remain unexplored in the literature.
@@
 and similarly for $x\otimes\mathtt{u}^{-1}_{2}(y)$.
-We note that the medialized associated quandles of the tensor products of racks in Definition 6.1 recover the tensor products of medial quandles that Crans and Nelson [Hom]*Section 8.1 introduced in 2014.
-On the other hand, these tensors are distinct from the tensor products that Kamada [tensors]*Definition 3.1 introduced in 2021, which are sets without canonical rack structures.
+We note that the medialized associated quandles of the tensor products of racks in Definition 6.1 recover the tensor products of medial quandles that Crans and Nelson [Hom]*Sec.
+8.1 introduced in 2014.
+On the other hand, these tensors are distinct from the tensor products that Kamada [tensors]*Def.
+3.1 introduced in 2021, which are sets without canonical rack structures.
 6.1.1. Bihomomorphisms
@@
 Before we can state this property, we will need the following definition from universal algebra.
-This is a special case of a more general definition for arbitrary algebraic theories; see, for example, [alg-theories]*Definition 3.10.2.
+This is a special case of a more general definition for arbitrary algebraic theories; see, for example, [alg-theories]*Def.
+3.10.2.
 Definition 6.3.
@@
 Regard racks and GL-racks as algebraic theories with binary operations $s_{-}(-)$ and $s^{-1}_{-}(-)$, along with unary operations $\mathtt{u}$ and $\mathtt{u}^{-1}$ for GL-racks.
-Due to Remark 6.2, the tensor product in Definition 6.1 is precisely the tensor product constructed in the proof of [alg-theories]*Theorem 3.10.3, and $\psi(x,y)=x\otimes y$ for all $(x,y)\in X\times Y$.
+Due to Remark 6.2, the tensor product in Definition 6.1 is precisely the tensor product constructed in the proof of [alg-theories]*Thm.
+3.10.3, and $\psi(x,y)=x\otimes y$ for all $(x,y)\in X\times Y$.
 Thus, the claim follows from the discussion in [alg-theories]*p.
@@
 6.1.3. Internal hom-tensor adjunctions
-Mirroring the work of Crans and Nelson [Hom]*Theorem 12 on the category of medial quandles, the next result continues the analogy with the category of modules over a ring by describing internal hom-tensor adjunctions in $\mathsf{Rack}_{\mathrm{med}}$ and $\mathsf{GLR}_{\mathrm{med}}$.
+Mirroring the work of Crans and Nelson [Hom]*Thm.
+12 on the category of medial quandles, the next result continues the analogy with the category of modules over a ring by describing internal hom-tensor adjunctions in $\mathsf{Rack}_{\mathrm{med}}$ and $\mathsf{GLR}_{\mathrm{med}}$.
 Theorem 6.7.
@@
 Recall that the algebraic theories of medial racks and medial GL-racks are commutative.
-Therefore, both claims are special cases of a general result for commutative algebraic theories; see, for example, [alg-theories]*Theorem 3.10.3.
-In particular, $\otimes_{\mathrm{med}}$ is precisely the tensor product constructed in the proof of [alg-theories]*Theorem 3.10.3, and the tensor unit is the free rack (resp.
+Therefore, both claims are special cases of a general result for commutative algebraic theories; see, for example, [alg-theories]*Thm.
+3.10.3.
+In particular, $\otimes_{\mathrm{med}}$ is precisely the tensor product constructed in the proof of [alg-theories]*Thm.
+3.10.3, and the tensor unit is the free rack (resp.
 GL-rack) $L$ on one element.
@@
 One may ask how much of the structure in Theorem 6.7 remains if we drop the mediality assumption.
-By [alg-theories]*Theorem 3.10.3, we lose the closed structure because racks and GL-racks are noncommutative algebraic theories.
+By [alg-theories]*Thm.
+3.10.3, we lose the closed structure because racks and GL-racks are noncommutative algebraic theories.
 Nevertheless, we show that tensor products in $\mathsf{Rack}$ and $\mathsf{GLR}$ surprisingly retain tensor units.
@@
 It would be interesting to study more properties and applications of the tensor products in Definition 6.1. In Section 7, we propose future work in this direction.
-7. Distinguishing Legendrian knots
-In this section, we use GL-racks to distinguish Legendrian knots whose classical invariants are identical.
-This answers a question of Kimura [bi]*Section 4 and completes the classification of Legendrian $8_{13}$ knots in the extended Legendrian knot atlases of Bhattacharyya et al.
-[atlas-site] and Petkova and Schwartz [new-atlas].
-By convention, we take all Legendrian links to be oriented.
-7.1. Legendrian knots and links
-In this subsection, we briefly summarize the Legendrian isotopy problem.
-Although we establish the relevant concepts here, we also refer the reader to [what] for an accessible introduction to Legendrian knot theory.
-For a more formal contact-geometric treatment, we refer the reader to [etnyre].
-7.1.1. Definitions
-Recall that a smooth link $\Lambda\subset\mathbb{R}^{3}$ is called Legendrian if it lies everywhere tangent to the standard contact structure $\ker(dz-y\,dx)$ on $\mathbb{R}^{3}$ depicted in Figure 1.
-An assignment of a plane to each point in 3D Euclidean space. The planes’ slopes are determined by the y-value at each point.
-Figure 1. The standard contact structure on $\mathbb{R}^{3}$. Reprinted from [us]*Figure 1.
-The front projection of a Legendrian link to the $xz$-plane viewed from the negative $y$-direction has cusps instead of vertical tangencies.
-Also, the overstrand of every crossing has a lower slope than the understrand.
-For example, Figure 2 depicts front projections of two Legendrian unknots.
-Two oriented knot diagrams representing Legendrian unknots. The knots have no crossings but have different numbers of cusps.
-Figure 2. Front projections of nonequivalent Legendrian unknots.
-One major problem in contact geometry is the classification of Legendrian knots up to Legendrian isotopy or, equivalently, up to finite sequences of Legendrian Reidemeister moves.
-For example, the Legendrian unknots in Figure 2 are nonequivalent; that is, they are not related by any Legendrian Reidemeister moves.
-7.1.2. Legendrian nonsimplicity
-To address the Legendrian isotopy problem, contact topologists have developed various invariants of Legendrian links. To distinguish Legendrian links, it suffices to distinguish their invariants.
-[…]
-One of the main challenges of the Legendrian isotopy problem is the existence of Legendrian nonsimple smooth knot types, which have nonequivalent Legendrian representatives whose classical invariants are identical.
-Oftentimes, other invariants also fail to distinguish representatives of Legendrian nonsimple knot types, especially for knot types of high arc index.
-This phenomenon opens up many conjectures in the Legendrian knot atlases of Chongchitmate and Ng [atlas], Bhattacharyya et al.
-[atlas-site], and Petkova and Schwartz [new-atlas], several of which we will address using GL-racks.
-7.2. GL-racks of Legendrian links
-In this subsection, we briefly recall how to assign a canonical GL-rack to a Legendrian link. We also discuss GL-rack coloring invariants.
-7.2.1. Invariance of GL-racks
-In 2023, Karmakar et al.
-[karmakar] and Kimura [bi] introduced GL-rack invariants of Legendrian links.
-These generalize similar invariants introduced by Kulkarni and Prathamesh [original] and Ceniceros et al.
-[ceniceros].
-The Legendrian Reidemeister moves are encoded in the bi-Legendrian rack axioms; see [bi]*Figures 6–8.
-Consequently, every Legendrian link $\Lambda$ is assigned a canonical GL-rack $\mathcal{G}(\Lambda)$ using any of its front projections, and the isomorphism type of $\mathcal{G}(\Lambda)$ is an invariant of $\Lambda$; see [karmakar]*Theorem 4.3.
-That said, there exist nonequivalent Legendrian knots with isomorphic GL-racks (see [leg-cocycle]*Examples 21–24), so $\mathcal{G}(\Lambda)$ is not a complete invariant of Legendrian links.
-7.2.2. Assignment of GL-racks to Legendrian links
-We summarize the construction of $\mathcal{G}(\Lambda)$ from [karmakar]*Section 4.
-During a traversal of any front projection of $\Lambda$, label the strands (i.e., connected components) of the projection by $x_{1},\dots,x_{k}$.
-Partition each strand into substrands each containing no cusps, labeling the substrands as stipulated in (U) and (D) of Figure 3.
-Negative and positive crossings and up- and down-cusps as they appear in Legendrian knots labeled with relations on GL-racks.
-Figure 3. Crossing relations and cusp labels used to obtain a presentation of $\mathcal{G}(\Lambda)$.
-Using Remark 3.16, define $\mathcal{G}(\Lambda)$ to be the free GL-rack on the set $\{x_{1},\dots,x_{k}\}$ modulo the congruence generated by the relations specified in (N) and (P) of Figure 3 at all crossings between uncusped substrands.
-Relations (N) and (P) respectively correspond to negative and positive orientations of crossings in the skein-theoretic sense; see [quandlebook]*p.
-87 for details.
-Note the similarity between the definition of $\mathcal{G}(\Lambda)$ and the combinatorial construction of fundamental quandles of smooth links; see, for example, [joyce]*Section 15 and cf.
-[leg-cocycle]*Remark 23.
-Example 7.1.
-Let $q\geq 3$ be an odd integer, let $L$ be a $(2,-q)$-torus knot, and let $\Lambda$ be the Legendrian representative of $L$ whose Thurston–Bennequin and rotation numbers are maximal.
-[^1] This choice is well-defined because torus knots are Legendrian simple; see [etnyre]*Subsection 5.2.
-In this example, we compute $\mathcal{G}(\Lambda)$ using the front projection in Figure 4.
-An oriented Legendrian negative torus knot with its cusped strands and uncusped substrands labeled.
-Figure 4. Front projection and cusp labels of the Legendrian $(2,-q)$-torus knot $\Lambda$ whose Thurston–Bennequin and rotation numbers are maximal.
-Starting at any crossing (which, in Figure 4, we arbitrarily choose to be the bottommost crossing), traverse $\Lambda$ along its depicted orientation.
-All crossings in this particular front projection are negative, so by imposing relation (N) from Figure 3, we compute that $\mathcal{G}(\Lambda)$ is the free GL-rack on the set $\{x_{1},\dots,x_{q}\}$ modulo the congruence generated by the relations
-$s_{\mathtt{u}(x_{1})}(x_{q})=\mathtt{u}\mathtt{d}(x_{2}),\quad s_{\mathtt{d}(x_{q})}(x_{q-1})=\mathtt{d}\mathtt{u}(x_{1}),\quad\text{and }s_{\mathtt{d}(x_{i-1})}(x_{i-2})=\mathtt{d}^{2}(x_{i})\quad\text{for all }3\leq i\leq q.$
-By bi-Legendrian rack axiom 3, these crossing relations simplify to
-$s_{x_{1}}(x_{q})=\mathtt{u}\mathtt{d}(x_{2}),\quad s_{x_{q}}(x_{q-1})=\mathtt{d}\mathtt{u}(x_{1}),\quad\text{and }s_{x_{i-1}}(x_{i-2})=\mathtt{d}^{2}(x_{i})\quad\text{for all }3\leq i\leq q.$
-7.2.3. Legendrian coloring invariants
-At the time of writing, the easiest known way to distinguish GL-racks of Legendrian links is by using the GL-rack coloring number $\operatorname{Col}(\Lambda,R)$ of each link $\Lambda$ with respect to a given GL-rack $R$.
-Defined to be the cardinality of the hom-set $\operatorname{Hom}_{\mathsf{GLR}}(\Lambda,R)$, $\operatorname{Col}(\Lambda,R)$ is an invariant of $\Lambda$; see [bi]*Proposition 3.9.
-Explicitly, a sufficient condition for two Legendrian links $\Lambda_{1},\Lambda_{2}$ to be nonequivalent is the existence of a GL-rack $R$ such that $\operatorname{Col}(\Lambda_{1},R)\neq\operatorname{Col}(\Lambda_{2},R)$.
-[…]
-Proposition 7.2.
-Let $\Lambda$ be a Legendrian link.
-Then for all medial GL-racks $M$, the isomorphism type of $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda),M)$ as a medial GL-rack is an invariant of $\Lambda$.
-7.3. Distinguishing results
-In this subsection, we use GL-rack coloring numbers to distinguish various Legendrian knots whose classical invariants are identical.
-Specifically, we consider conjecturally distinct pairs of Legendrian $6_{2}$, $8_{10}$, and $8_{13}$ knots listed in the atlases of Chongchitmate and Ng [atlas], Bhattacharyya et al.
-[atlas-site], and Petkova and Schwartz [new-atlas].
-7.3.1. Discussion
-Prior to this work, various authors have used coloring invariants to distinguish Legendrian representatives of Legendrian simple knot types.
-Kulkarni and Prathamesh [original]*Main Theorem 2, Kimura [bi]*Theorem 4.1, and Karmakar et al.
-[karmakar]*Theorem 4.6 each used coloring invariants to distinguish infinitely many Legendrian unknots.
-Karmakar et al.
-[karmakar]*Theorem 4.7 also used GL-rack coloring numbers to distinguish infinitely many Legendrian trefoils, and Ceniceros et al.
-[ceniceros]*Example 16 used them to distinguish connected sums of Legendrian trefoils.
-However, to our knowledge at the time of writing, the following is the first application of GL-racks to distinguish Legendrian knots whose classical invariants are identical.
-7.3.2. Finding coloring maps
-We employ a well-known characterization of homomorphisms of models of an algebraic theory in terms of generators and relations.
-Similarly to group homomorphisms, GL-rack homomorphisms from $\mathcal{G}(\Lambda)$ to $R$ are determined by the images of the generators of $\mathcal{G}(\Lambda)$ and whether these images satisfy the relations of $\mathcal{G}(\Lambda)$ as equations in $R$; see [leg-cocycle]*Remark 12.
-In Appendix A.3, we describe algorithms that compute GL-rack coloring numbers.
-These algorithms helped us select the GL-racks and coloring homomorphisms used in this subsection.
-We provide an implementation of our search algorithms in a GitHub repository [my-code].
-[…]
-7.3.3. Legendrian $6_{2}$ knots
-In 2021, Dynnikov and Prasolov [6-2]*Proposition 2.3 used impressive topological and combinatorial machinery to distinguish the two Legendrian $6_{2}$ knots with classical invariants $(\operatorname{tb},\operatorname{rot})=(-7,2)$ in Figure 5, settling a conjecture of Chongchitmate and Ng [atlas].
-At the time of writing, we are unaware of any other proofs of this conjecture.
-Indeed, linearized contact homology and the ruling invariant fail to distinguish these Legendrian knots; see [atlas].
-Two oriented Legendrian 6 2 knots with equal classical invariants. All cusped strands and uncusped substrands are labeled.
-Figure 5. Front projections of the two Legendrian representatives of the topological knot $6_{2}$ with $(\operatorname{tb},\operatorname{rot})=(-7,2)$ given in [atlas].
-Using GL-rack coloring numbers, we offer a simpler proof that these Legendrian knots are nonequivalent.
-Let $\Lambda_{1}$ and $\Lambda_{2}$ be the Legendrian knots on the left and right of Figure 5, respectively.
-First, we compute presentations for $\mathcal{G}(\Lambda_{1})$ and $\mathcal{G}(\Lambda_{2})$.
-Using Figure 3, we find that $\mathcal{G}(\Lambda_{1})$ is the free GL-rack on the set $\{x_{1},\dots,x_{6}\}$ modulo the congruence generated by the following relations:
-(9) $\mathcal{G}(\Lambda_{1})\begin{cases}s_{\mathtt{u}(x_{1})}\mathtt{u}(x_{4})=x_{5}\iff s_{x_{1}}\mathtt{u}(x_{4})=x_{5},\\ s_{x_{4}}\mathtt{d}\mathtt{u}(x_{1})=x_{2},\\ s_{\mathtt{d}(x_{2})}(x_{1})=\mathtt{d}^{2}(x_{6})\iff s_{x_{2}}(x_{1})=\mathtt{d}^{2}(x_{6}),\\ s_{x_{5}}(x_{3})=\mathtt{u}\mathtt{d}(x_{2}),\\ s_{\mathtt{d}(x_{3})}(x_{6})=\mathtt{d}(x_{5})\iff s_{x_{3}}(x_{6})=\mathtt{d}(x_{5}),\\ s_{\mathtt{d}(x_{6})}(x_{4})=\mathtt{d}^{2}(x_{3})\iff s_{x_{6}}(x_{4})=\mathtt{d}^{2}(x_{3}).\end{cases}$
-We have simplified the first, third, fifth, and sixth relations using bi-Legendrian rack axiom 3.
-Similarly, we compute $\mathcal{G}(\Lambda_{2})$ to be the free GL-rack on the set $\{x_{1},\dots,x_{7}\}$ modulo the congruence generated by the following relations:
-(10) $\mathcal{G}(\Lambda_{2})\begin{cases}s_{x_{1}}\mathtt{u}\mathtt{d}^{2}(x_{3})=x_{4},&s_{x_{5}}(x_{3})=\mathtt{d}(x_{2}),\\ s_{x_{1}}(x_{6})=x_{7},&s_{x_{3}}(x_{6})=\mathtt{d}^{2}(x_{5}),\\ s_{x_{6}}(x_{2})=\mathtt{u}(x_{1}),&s_{x_{3}}(x_{7})=x_{1}.\\ s_{x_{2}}(x_{5})=\mathtt{d}(x_{4}),\\ \end{cases}$
-Theorem 7.3.
-The two Legendrian $6_{2}$ knots with $(\operatorname{tb},\operatorname{rot})=(-7,2)$ in Figure 5 are nonequivalent; they are distinguishable using coloring numbers with respect to a permutation GL-rack of order 3.
-Proof.
-Let $Y:=\{1,2,3\}$.
-In cycle notation, let $\sigma\in S_{3}$ be the permutation $(123)$.
-In the notation of Example 3.5, let $R:=(Y,\sigma,\sigma^{-1})_{\mathrm{perm}}$, so that $R$ is the $11$th GL-rack in Table B.2.
-By Proposition 3.11, we have $\mathtt{d}=\operatorname{id}_{Y}$ in $R$.
-We will show that $\operatorname{Col}(\Lambda_{2},R)>\operatorname{Col}(\Lambda_{1},R)$.
-To that end, let $X$ denote the underlying set of $\mathcal{G}(\Lambda_{2})$ as presented in (10), and define $\psi:X\to Y$ by
-$\psi(x_{i}):=\begin{cases}1&\text{if }i\in\{1,3,4\},\\ 2&\text{if }i\in\{2,6\},\\ 3&\text{if }i\in\{5,7\}.\end{cases}$
-Using the relations in system (10), it is straightforward to verify that $\psi$, $\sigma\psi$, and $\sigma^{2}\psi$ are GL-rack homomorphisms from $\mathcal{G}(\Lambda_{2})$ to $R$.
-Hence, $\operatorname{Col}(\Lambda_{2},R)\geq 3$.
-(In fact, using a similar method as in the remainder of this proof, one can show that this bound is actually an equality.)
-On the other hand, we claim that $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda_{1}),R)=\emptyset$.
-Using the presentation of $\mathcal{G}(\Lambda_{1})$ in system (9), let $\varphi$ be any GL-rack homomorphism from $\mathcal{G}(\Lambda_{1})$ to $R$.
-Apply $\varphi$ to system (9).
-Then the images $y_{i}:=\varphi(x_{i})\in Y$ satisfy the following system of equations in $R$:
-(11) $R\begin{cases}\sigma\sigma^{-1}(y_{4})=y_{5}\iff y_{4}=y_{5},\\ \sigma\sigma^{-1}(y_{1})=y_{2}\iff y_{1}=y_{2},\\ \sigma(y_{1})=y_{6},\\ \sigma(y_{3})=\sigma^{-1}(y_{2})\iff y_{3}=\sigma(y_{2}),\\ \sigma(y_{6})=y_{5},\\ \sigma(y_{4})=y_{3}.\end{cases}$
-Here, we have used the fact that $\sigma^{3}=\operatorname{id}_{Y}$ to rewrite the fourth equality. We deduce that
-$\sigma(y_{2})=y_{3}=\sigma(y_{4})=\sigma(y_{5})=\sigma^{2}(y_{6})=\sigma^{3}(y_{1})=y_{1}=y_{2},$
-but $\sigma$ has no fixed points in $Y$. Hence, system (11) has no solutions in $R$, so $\varphi$ cannot exist. ∎
-Incidentally, Theorem 7.3 gives a positive answer to a question that Kimura [bi]*Section 4 posed in 2023, as we state below; cf. [bi]*Theorem 4.3.
-Corollary 7.4.
-GL-rack coloring numbers by nonquandle GL-racks are not generally unable to distinguish nonequivalent Legendrian knots whose classical invariants are identical, even when linearized contact homology and the ruling invariant fail to do so.
-In Appendix A.3, we show a version (Proposition A.2) of Corollary 7.4 for Legendrian racks.
-7.3.4. Legendrian $8_{10}$ knots
-Next, we distinguish the Legendrian $8_{10}$ knots with classical invariants $(\operatorname{tb},\operatorname{rot})=(-8,3)$ in Figure 6.
-This confirms a conjecture of Bhattacharyya et al.
-[atlas-site] and Petkova and Schwartz [new-atlas].
-[^2] While we and Bhattacharyya et al.
-[atlas-site] refer to the $8_{10}$ knot by its name in the Rolfsen table, Petkova and Schwartz [new-atlas] call it the $m8a10$ knot.
-At the time of writing, we are unaware of any other proofs of this conjecture.
-Two oriented Legendrian 8 10 knots with equal classical invariants. All cusped strands and uncusped substrands are labeled.
-Figure 6. Front projections of the two Legendrian representatives of the topological knot $8_{10}$ with $(\operatorname{tb},\operatorname{rot})=(-8,3)$ given in [atlas-site].
-Let $\Lambda_{1}$ and $\Lambda_{2}$ be the Legendrian knots on the left and right of Figure 6, respectively.
-Using Figure 3 and bi-Legendrian rack axiom 3, we compute that $\mathcal{G}(\Lambda_{1})$ and $\mathcal{G}(\Lambda_{2})$ are the free GL-racks on the set $\{x_{1},\dots,x_{10}\}$ modulo the congruences generated by the following relations:
-(12) $\mathcal{G}(\Lambda_{1})\begin{cases}s_{x_{1}}\mathtt{u}(x_{3})=x_{4},&s_{x_{4}}\mathtt{d}\mathtt{u}(x_{1})=x_{2},\\ s_{x_{7}}(x_{5})=\mathtt{d}(x_{4}),&s_{x_{5}}(x_{8})=\mathtt{d}(x_{7}),\\ s_{x_{8}}(x_{6})=\mathtt{d}^{2}(x_{5}),&s_{x_{6}}(x_{9})=\mathtt{d}(x_{6}),\\ s_{x_{6}}(x_{3})=x_{2},&s_{x_{6}}(x_{10})=x_{1},\\ s_{x_{3}}(x_{7})=\mathtt{u}\mathtt{d}^{2}(x_{6}),&s_{x_{3}}(x_{9})=x_{10}.\end{cases}\;\mathcal{G}(\Lambda_{2})\begin{cases}s_{x_{1}}(x_{7})=\mathtt{u}\mathtt{d}(x_{6}),&s_{x_{5}}(x_{2})=x_{3},\\ s_{x_{7}}(x_{2})=\mathtt{d}\mathtt{u}(x_{1}),&s_{x_{3}}\mathtt{u}(x_{5})=x_{6},\\ s_{x_{10}}(x_{4})=\mathtt{d}^{2}(x_{3}),&s_{x_{3}}(x_{10})=x_{9},\\ s_{x_{4}}(x_{1})=\mathtt{d}^{2}(x_{10}),&s_{x_{6}}\mathtt{d}(x_{4})=x_{5},\\ s_{x_{5}}\mathtt{d}^{2}(x_{7})=x_{8},&s_{x_{6}}(x_{9})=x_{8}.\end{cases}$
-Theorem 7.5.
-The two Legendrian $8_{10}$ knots with $(\operatorname{tb},\operatorname{rot})=(-8,3)$ in Figure 6 are nonequivalent; they are distinguishable using coloring numbers with respect to a permutation GL-rack of order 2.
-Proof.
-Let $Y:=\{1,2\}$.
-In cycle notation, let $\sigma\in S_{2}$ be the permutation $(12)$.
-Let $R$ be the permutation GL-rack $(Y,\sigma,\sigma)_{\mathrm{perm}}$, so that $R$ is the fourth GL-rack in Table B.1.
-By Proposition 3.11, we have $\mathtt{d}=\operatorname{id}_{Y}$ in $R$.
-We will show that $\operatorname{Col}(\Lambda_{1},R)>\operatorname{Col}(\Lambda_{2},R)$.
-To that end, let $X$ denote the underlying set of $\mathcal{G}(\Lambda_{1})$ as presented in system (12), and define $\psi:X\to Y$ by
-$\psi(x_{i}):=\begin{cases}1&\text{if }i\in\{1,2,5,8,9\},\\ 2&\text{if }i\in\{3,4,6,7,10\}.\end{cases}$
-Using the relations of $\mathcal{G}(\Lambda_{1})$ in (12), it is straightforward to verify that $\psi$ and $\sigma\psi$ are GL-rack homomorphisms from $\mathcal{G}(\Lambda_{1})$ to $R$.
-Hence, $\operatorname{Col}(\Lambda_{1},R)\geq 2$ (which is actually an equality).
-On the other hand, we claim that $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda_{2}),R)=\emptyset$.
-Using the presentation of $\mathcal{G}(\Lambda_{2})$ in system (12), let $\varphi$ be any GL-rack homomorphism from $\mathcal{G}(\Lambda_{2})$ to $R$.
-Once again, the elements $y_{i}:=\varphi(x_{i})\in Y$ must satisfy the following system of equations in $R$:
-$R\begin{cases}\sigma(y_{7})=\sigma(y_{6}),&\sigma(y_{2})=\sigma(x_{1}),\\ \sigma(y_{2})=y_{3},&\sigma^{2}(y_{5})=y_{6},\\ \sigma(y_{10})=y_{9},&\sigma(y_{4})=y_{3},\\ \sigma(y_{1})=y_{10},&\sigma(y_{4})=y_{5},\\ \sigma(y_{7})=y_{8},&\sigma(y_{9})=y_{8}.\end{cases}$
-Since $\sigma^{2}=\operatorname{id}_{Y}$, we can also rewrite the equalities $\sigma(y_{7})=\sigma(x_{6})$, $\sigma^{2}(y_{5})=y_{6}$, and $\sigma(y_{10})=y_{9}$ as $y_{7}=y_{6}$, $y_{5}=y_{6}$, and $y_{10}=\sigma(y_{9})$, respectively.
-Therefore,
-$y_{7}=y_{6}=y_{5}=\sigma(y_{4})=y_{3}=\sigma(y_{2})=\sigma(y_{1})=y_{10}=\sigma(y_{9})=y_{8}=\sigma(y_{7}),$
-but $\sigma$ has no fixed points in $Y$. Hence, $\varphi$ cannot exist. ∎
-7.3.5. Legendrian $8_{13}$ knots
-Finally, we distinguish the Legendrian $8_{13}$ knots with classical invariants $(\operatorname{tb},\operatorname{rot})=(-6,1)$ in Figure 7.
-This completes the classification of Legendrian $8_{13}$ knots in the extended Legendrian knot atlases of Bhattacharyya et al.
-[atlas-site] and Petkova and Schwartz [new-atlas].
-[^3] Petkova and Schwartz [new-atlas] call the $8_{13}$ knot the $m8a13$ knot.
-At the time of writing, we are unaware of any other proofs that these Legendrian knots are nonequivalent.
-Two oriented Legendrian 8 13 knots with equal classical invariants. All cusped strands and uncusped substrands are labeled.
-Figure 7. Front projections of the two Legendrian representatives of the topological knot $8_{13}$ with $(\operatorname{tb},\operatorname{rot})=(-6,1)$ given in [atlas-site].
-Let $\Lambda_{1}$ and $\Lambda_{2}$ be the Legendrian knots on the left and right of Figure 6, respectively.
-Similarly to before, $\mathcal{G}(\Lambda_{1})$ and $\mathcal{G}(\Lambda_{2})$ are respectively the free GL-racks on the sets $\{x_{1},\dots,x_{10}\}$ and $\{x_{1},\dots,x_{9}\}$ modulo the congruences generated by the following relations:
-(13) $\mathcal{G}(\Lambda_{1})\begin{cases}s_{x_{1}}(x_{5})=\mathtt{u}\mathtt{d}(x_{4}),&s_{x_{4}}(x_{9})=\mathtt{d}(x_{8}),\\ s_{x_{7}}(x_{2})=\mathtt{u}(x_{1}),&s_{x_{4}}(x_{1})=\mathtt{d}^{2}(x_{10}),\\ s_{x_{2}}(x_{7})=x_{6},&s_{x_{5}}(x_{8})=\mathtt{u}(x_{7}),\\ s_{x_{6}}(x_{3})=x_{2},&s_{x_{5}}(x_{10})=x_{9},\\ s_{x_{5}}(x_{4})=x_{3},&s_{x_{10}}(x_{6})=\mathtt{d}(x_{5}).\end{cases}\;\mathcal{G}(\Lambda_{2})\begin{cases}s_{x_{1}}(x_{7})=\mathtt{u}(x_{6}),&s_{x_{3}}(x_{8})=\mathtt{d}(x_{7}),\\ s_{x_{1}}(x_{5})=x_{4},&s_{x_{9}}(x_{4})=\mathtt{d}(x_{3}),\\ s_{x_{4}}(x_{2})=\mathtt{u}(x_{1}),&s_{x_{9}}(x_{6})=x_{5},\\ s_{x_{2}}(x_{9})=x_{8},&s_{x_{6}}(x_{1})=\mathtt{d}^{2}(x_{9}).\\ s_{x_{8}}(x_{3})=x_{2},\end{cases}$
-Theorem 7.6.
-The two Legendrian $8_{13}$ knots with $(\operatorname{tb},\operatorname{rot})=(-6,1)$ in Figure 7 are nonequivalent; they are distinguishable using coloring numbers with respect to a permutation GL-rack of order 3.
-Proof.
-Let $Y=\{1,2,3\}$.
-In cycle notation, let $\sigma\in S_{3}$ be the permutation $(123)$, and let $R$ be the permutation rack $(Y,\sigma,\operatorname{id}_{Y})_{\mathrm{perm}}$, so that $R$ is the tenth GL-rack listed in Table B.2.
-By Proposition 3.11, we have $\mathtt{d}=\sigma^{-1}=(132)$ in $R$.
-Let $X$ be the underlying set of $\mathcal{G}(\Lambda_{1})$ as presented in system (13), and define $\psi:X\to Y$ by
-$\psi(x_{i}):=\begin{cases}1&\text{if }i\in\{1,4,8,10\},\\ 2&\text{if }i\in\{3,5,6,7\},\\ 3&\text{if }i\in\{2,6\}.\end{cases}$
-It is straightforward to verify that $\varphi$, $\sigma\psi$, and $\sigma^{2}\psi$ are GL-rack homomorphisms from $\mathcal{G}(\Lambda_{1})$ to $R$, so $\operatorname{Col}(\Lambda_{1},R)\geq 3$ (which is actually an equality).
-On the other hand, an argument similar to the ones in Theorems 7.3 and 7.5 shows that $\varphi(x_{6})=\sigma\varphi(x_{6})$ for all GL-rack homomorphisms $\varphi\in\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda_{2}),R)$; we leave the details to the reader.
-Since $\sigma$ has no fixed points in $Y$, we conclude that $\operatorname{Col}(\Lambda_{2},R)=0<\operatorname{Col}(\Lambda_{1},R)$, so $\Lambda_{1}$ and $\Lambda_{2}$ are nonequivalent.
-∎
 7. Directions for future work
-In light of our results, we propose questions for further research.
-  1. (1)
-[…]
+We conclude by proposing questions for further research. In the following, let $\mathcal{G}(\Lambda)$ denote the fundamental GL-rack of a Legendrian link $\Lambda$ (see [karmakar]).
   1. (1)
@@
   4. (4)
-    In light of Proposition 7.2, do there exist Legendrian knots $\Lambda_{1}$ and $\Lambda_{2}$ and a medial GL-rack $M$ such that $\operatorname{Col}(\Lambda_{1},M)=\operatorname{Col}(\Lambda_{2},M)$ but $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda_{1}),M)\not\cong\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda_{2}),M)$?
-In the spirit of Corollary 7.4, do there exist such $\Lambda_{1}$ and $\Lambda_{2}$ whose classical invariants are identical?
-A positive answer would show that $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda),M)$ is a proper enhancement of $\operatorname{Col}(\Lambda,M)$ as an invariant of $\Lambda$.
+[…]
   5. (5)
@@
 In 2023, Elhamdadi et al.
-[enhancements]*Theorems 4.2 and 5.1 properly enhanced medial quandle-valued invariants of smooth links using $\mathbb{F}$-algebra homomorphisms between quandle rings and colorings of smooth links by idempotents of quandle rings.
+[enhancements]*Thms.
+4.2 and 5.1 properly enhanced medial quandle-valued invariants of smooth links using $\mathbb{F}$-algebra homomorphisms between quandle rings and colorings of smooth links by idempotents of quandle rings.
 Do similar proper enhancements of $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda),M)$ exist?
   6. (6)
-    In light of Example 6.6, do tensor products of left-distributive racks, medial quandles (cf. [Hom]*Section 8.1), or Latin quandles have any interesting properties?
+    In light of Example 6.6, do tensor products of left-distributive racks, medial quandles (cf. [Hom]*Sec. 8.1), or Latin quandles have any interesting properties?
   7. (7)
-    Theorem 6.9 shows that universal-algebraic tensor products make $\mathsf{Rack}$ and $\mathsf{GLR}$ into symmetric magmoidal categories with units, leading us to ask what their unital nuclei are; see [magmoidal]*Section 2.5.
+    Theorem 6.9 shows that universal-algebraic tensor products make $\mathsf{Rack}$ and $\mathsf{GLR}$ into symmetric magmoidal categories with units, leading us to ask what their unital nuclei are; see [magmoidal]*Sec.
+2.5.
 In particular, is $\otimes$ associative?
@@
   10. (10)
-    Extend the results of Sections 36 and the algorithms in Appendix A to 4-Legendrian racks and 4-Legendrian biracks, which Kimura [leg-cocycle]*Section 4 introduced in 2024.
+    Extend the results of Sections 36 and the algorithms in Appendix A to 4-Legendrian racks and 4-Legendrian biracks, which Kimura [leg-cocycle]*Sec. 4 introduced in 2024.
   11. (11)
@@
   13. (13)
-    Transverse knots are knots that lie everywhere transverse to the standard contact structure on $\mathbb{R}^{3}$; see [etnyre]*Section 2.4.
+    Transverse knots are knots that lie everywhere transverse to the standard contact structure on $\mathbb{R}^{3}$; see [etnyre]*Sec.
+2.4.
 Can one define rack-theoretic invariants of transverse knots?
@@
 For comparison, we also list the corresponding numbers for classical racks and quandles.
-We obtained the numbers $g(n)$ from Algorithm A.1 and $g^{m}(n)$, $g_{q}(n)$, and $g^{m}_{q}(n)$ from Algorithm A.2.
-Meanwhile, the numbers $r(n)$, $r^{m}(n)$, $r_{q}(n)$, and $r^{m}_{q}(n)$ were originally computed by McCarron [rack-count] in 2010, Vojtěchovský and Yang [library] in 2019, Henderson et al.
+We obtained the first four columns from Algorithm A.1, while the last four columns were originally computed by McCarron [rack-count] in 2010, Vojtěchovský and Yang [library] in 2019, Henderson et al.
 [quandle-count] in 2006, and Jedlička et al.
@@
 It appears that each of $g(n)$, $g^{m}(n)$, $g_{q}(n)$, and $g^{m}_{q}(n)$ in Table A.1 grows exponentially and at a much faster rate than its counterpart for classical racks.
+Table A.1.
+The numbers of GL-racks $g(n)$, medial GL-racks $g^{m}(n)$, GL-quandles $g_{q}(n)$, and medial GL-quandles $g^{m}_{q}(n)$ of orders $0\leq n\leq 8$ up to isomorphism, compared against the corresponding numbers of racks $r(n)$, medial racks $r^{m}(n)$, quandles $r_{q}(n)$, and medial quandles $r^{m}_{q}(n)$.
 $n$ 00 $1$ $2$ $3$ $4$ $5$ $6$ $7$ $8$
@@
 $r^{m}_{q}(n)$ $1$ $1$ $1$ $3$ $6$ $18$ $58$ $251$ $1410$
-Table A.1.
-The numbers of GL-racks $g(n)$, medial GL-racks $g^{m}(n)$, GL-quandles $g_{q}(n)$, and medial GL-quandles $g^{m}_{q}(n)$ of orders $0\leq n\leq 8$ up to isomorphism, compared against the corresponding numbers of racks $r(n)$, medial racks $r^{m}(n)$, quandles $r_{q}(n)$, and medial quandles $r^{m}_{q}(n)$.
 For explicit representatives of each GL-rack isomorphism class counted in Table A.1, see Appendix B for those of orders $2\leq n\leq 4$ and the GitHub repository in [my-code] for those of orders $5\leq n\leq 8$.
@@
 A.2. Classification of small GL-racks
-We now distinguish the exhaustive search algorithms in GAP [GAP4] that we used to compute these isomorphism classes.
+We discuss the exhaustive search algorithms in GAP [GAP4] that we used to compute these isomorphism classes.
 We build upon the work of Vojtěchovský and Yang [library] in 2019, who classified racks up to order $11$ [library-site].
@@
 The last line of the above pseudocode uses Proposition 3.11.
-On the author’s personal computer, running our implementation of Algorithm A.1 in GAP with $n=7$ took just under two minutes, while the computation for $n=8$ took seven days.
+On the author’s personal computer, running our implementation of Algorithm A.1 in GAP with $n=7$ took just under two minutes, while the computation for $n=8$ took about $100$ minutes.
 By contrast, a similar program using Definition 3.9 took three hours for the $n=7$ computation.
 This shows that verifying Definition 3.1 is significantly easier than verifying the bi-Legendrian rack axioms.
-Data: List isoClasses of isomorphism classes of GL-racks with underlying set $X=\{1,\dots,n\}$ returned by Algorithm A.1
-Result: Lists $\mathcal{Q}_{n}$, $\mathcal{M}_{n}$, and $\mathcal{B}_{n}$ of isomorphism classes of GL-quandles, medial GL-racks, and medial GL-quandles, respectively, whose underlying set is $X$
-begin
-       $\mathcal{Q}_{n},\mathcal{M}_{n},\mathcal{B}_{n},\textnormal{{seen}}\leftarrow\emptyset$;
-       foreach list $R=[s,\mathtt{u},\mathtt{d}]$ in isoClasses do
-             if seen contains a list $[t,\textnormal{{isQuandle}},\textnormal{{isMedial}}]$ such that $t=s$ then
-                   if isQuandle then Add($\mathcal{Q}_{n}$, $R$);
-                   if isMedial then
-                         Add($\mathcal{M}_{n}$, $R$);
-                         if isQuandle then Add($\mathcal{B}_{n}$, $R$);
-             else
-                   if $\mathtt{d}=\mathtt{u}^{-1}$ then
-                         isQuandle $\leftarrow$ true;
-                         Add($\mathcal{Q}_{n}$, $R$);
-                   else isQuandle $\leftarrow$ false;
-                   isMedial $\leftarrow$ true;
-                   foreach ordered triple $(x,y,z)$ in $X^{3}$ do
-                         if $s_{s_{x}(z)}s_{y}\neq s_{s_{x}(y)}s_{z}$ then
-                               isMedial $\leftarrow$ false;
-                               break;
-                  if isMedial then
-                         Add($\mathcal{M}_{n}$, $R$);
-                         if isQuandle then Add($\mathcal{B}_{n}$, $R$);
-                  Add(seen,$[s,\textnormal{{isQuandle}},\textnormal{{isMedial}}]$);
-Algorithm A.2 Classification of medial GL-racks, all GL-quandles, and medial GL-quandles of order $n$ up to isomorphism, given a classification of GL-racks of order $n$.
-A.2.2. Tabulation of medial GL-racks and GL-quandles
-Algorithm A.2 tests whether or not each GL-rack in the output of Algorithm A.1 is medial or a GL-quandle.
-To test for mediality, the algorithm simply verifies equation (1) for all elements $x,y,z\in X$.
-By Remark 3.10, to test whether a GL-rack $[s,\mathtt{u},\mathtt{d}]$ is a GL-quandle, it suffices to verify whether $\mathtt{d}=\mathtt{u}^{-1}$.
-On the author’s personal computer, running our implementation of Algorithm A.2 in GAP with $n=7$ took just under an hour, while the $n=8$ computation took eight days.
-A.3. Exhaustive searches for GL-rack coloring numbers
-We now describe Algorithm A.3, which computes all colorings of the GL-rack of an oriented Legendrian link $\Lambda$ by each GL-rack in the list $\mathsf{isoClasses}$ computed by Algorithm A.1.
-Before running Algorithm A.3, the user must input a presentation of $\mathcal{G}(\Lambda)$ in terms of crossing relations between generators of $\mathcal{G}(\Lambda)$; see Section 7.3 for examples of such presentations.
-By the discussion in Section 7.3, it suffices for the algorithm to search for all valid solutions in $R$ to the inputted crossing relations.
-Data: List isoClasses of isomorphism classes of GL-racks with underlying set $Y=\{1,\dots,n\}$ from Algorithm A.1 and a presentation of $\mathcal{G}(\Lambda)=[X,s^{\Lambda},\mathtt{u}^{\Lambda},\mathtt{d}^{\Lambda})]$
-Result: List solutions whose elements are lists $[s,\mathtt{u},\mathtt{d},\mathbf{y}]$ such that the mapping $x_{i}\mapsto y_{i}$ defines a GL-rack homomorphism $\mathcal{G}(\Lambda)\to[Y,s,\mathtt{u},\mathtt{d}]$
-begin
-       $m\leftarrow|X_{\Lambda}|$;
-[…]
-       foreach GL-rack $[Y,s,\mathtt{u},\mathtt{d}]$ in isoClasses do
-             foreach ordered $m$-tuple $\mathbf{y}\leftarrow(y_{1},\dots,y_{m})$ in $Y^{m}$ do
-                   if all crossing relations are satisfied after replacing each $x_{i}\in X_{\Lambda}$, $s^{\Lambda}$, $\mathtt{u}^{\Lambda}$, and $\mathtt{d}^{\Lambda}$ with $y_{i}$, $s$, $\mathtt{u}$, and $\mathtt{d}$, respectively then Add(solutions, $[s,\mathtt{u},\mathtt{d},\mathbf{y}]$);
-Algorithm A.3 Computation of colorings of an oriented Legendrian link $\Lambda$ by GL-racks of a given order $1\leq n\leq 11$ as computed by Algorithm A.1.
-In particular, if $R=(Y,s,\mathtt{u},\mathtt{d})$ is a bi-Legendrian rack of order $n\leq 11$, then $\operatorname{Col}(\Lambda,R)$ is simply the number of lists in $\mathsf{solutions}$ produced by Algorithm A.3 whose first three list elements are $s$, $\mathtt{u}$, and $\mathtt{d}$.
-To distinguish two oriented Legendrian links $\Lambda_{1}$ and $\Lambda_{2}$, it suffices to run Algorithm A.3 twice, once inputting $\mathcal{G}(\Lambda_{1})$ and again inputting $\mathcal{G}(\Lambda_{2})$, and find a GL-rack $R$ in $\mathsf{isoClasses}$ such that $\operatorname{Col}(\Lambda_{1},R)\neq\operatorname{Col}(\Lambda_{2},R)$.
-For example, running Algorithm A.3 with $n=2,3$ is how we determined which GL-racks and homomorphisms to employ in Section 7.3.
-Running the algorithm with $n=5$ also gave us the following example.
-Example A.1.
-In this example, we use Algorithm A.3 to once again distinguish the Legendrian $6_{2}$ knots $\Lambda_{1}$ and $\Lambda_{2}$ on the left and right of Figure 5, respectively.
-This time, we use the $222$nd GL-rack $R$ of order $5$ listed in the data linked above, which is a Legendrian rack.
-Let $Y:=\{1,2,3,4,5\}$.
-In cycle notation, let $\sigma,\tau\in S_{5}$ be the $5$-cycles $\sigma:=(12345)$ and $\tau:=(13524)$.
-Let $R:=(Y,\sigma,\tau)_{\mathrm{perm}}$.
-By Corollary 3.13, $R$ is a Legendrian rack.
-We input the relations of $\mathcal{G}(\Lambda_{1})$ in (9) and then those of $\mathcal{G}(\Lambda_{2})$ in (10) into our GAP implentation of Algorithm A.3.
-After running the program with $n=5$, the program outputs the text in Figure A.1 upon reaching $\mathsf{isoClasses}[222]=R$.
-The output states that $\operatorname{Col}(\Lambda_{1},R)=0\neq 5=\operatorname{Col}(\Lambda_{2},R)$, and the images of the generators $(x_{1},\dots,x_{7})$ of $\mathcal{G}(\Lambda_{2})$ under each element of $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda_{2}),R)$ are given by the orbit of $(1,2,3,5,1,4,5)\in Y^{7}$ under the action of the subgroup $\langle\sigma\rangle\leq S_{5}$ on $Y^{7}$.
-In particular, Example A.1 yields an analogue of Corollary 7.4 for Legendrian racks.
-Proposition A.2.
-There exist Legendrian knots sharing the same classical invariants that are distinguished by the Legendrian rack coloring numbers originally defined in [ceniceros]*Proposition 1.
-Finding all colorings of knot 1 by GL-rack 222 of 308…
-Finding all colorings of knot 2 by GL-rack 222 of 308…
-[ [(1,2,3,4,5), … (1,2,3,4,5)], (1,3,5,2,4), (1,3,5,2,4), 1, 2, 3, 5, 1, 4, 5 ]
-[ [(1,2,3,4,5), … (1,2,3,4,5)], (1,3,5,2,4), (1,3,5,2,4), 2, 3, 4, 1, 2, 5, 1 ]
-[ [(1,2,3,4,5), … (1,2,3,4,5)], (1,3,5,2,4), (1,3,5,2,4), 3, 4, 5, 2, 3, 1, 2 ]
-[ [(1,2,3,4,5), … (1,2,3,4,5)], (1,3,5,2,4), (1,3,5,2,4), 4, 5, 1, 3, 4, 2, 3 ]
-[ [(1,2,3,4,5), … (1,2,3,4,5)], (1,3,5,2,4), (1,3,5,2,4), 5, 1, 2, 4, 5, 3, 4 ]
-Number of colorings of knot 1 by GL-rack 222 of 308: 0
-Number of colorings of knot 2 by GL-rack 222 of 308: 5
-Since their GL-rack coloring numbers are distinct, these knots are not Legendrian isotopic.
-Figure A.1.
-Excerpt from the output of our GAP implementation of Algorithm A.3 with $n=5$.
-Here, knots $1$ and $2$ are the Legendrian knots in Figure 5, while GL-rack $222$ of $308$ is the Legendrian rack $R$ described in Example A.1.
+Our GAP implementation of Algorithm A.1 also tests whether the underlying rack of each GL-rack is medial, a quandle, or both.
+The former verifies equation (1), while the latter verifies the condition in Remark 3.10.
+We omit these obvious checks from the above pseudocode.
 Appendix B Tabulation of GL-racks of orders 2, 3, and 4
@@
 These entries denote all valid bi-Legendrian structures $[\mathtt{u},\mathtt{d}]$ up to isomorphism on the rack ($(\{1,\dots,n\}\},s)$, where $s$ is given by the corresponding entry in the first column.
-For example, the $11$th entry in Table B.2 is the permutation GL-rack of order 3 with $\sigma=(123)$ and bi-Legendrian structure $[\mathtt{u},\mathtt{d}]=[(132),\operatorname{id}]$, which we used to prove Theorem 7.3.
+For example, the $11$th entry in Table B.2 is the permutation GL-rack of order 3 with $\sigma=(123)$ and bi-Legendrian structure $[\mathtt{u},\mathtt{d}]=[(132),\operatorname{id}]$.
+Table B.1. The four isomorphism classes of GL-racks of order 2.
 $[s_{1},s_{2}]$ $[\mathtt{u},\mathtt{d}]$ GL-quandle? Medial?
@@
 No Yes
-Table B.1. The four isomorphism classes of GL-racks of order 2.
+Table B.2. The 13 isomorphism classes of GL-racks of order 3.
 $[s_{1},s_{2},s_{3}]$ $[\mathtt{u},\mathtt{d}]$ GL-quandle? Medial?
@@
 $[(23),(13),(12)]$ $[\operatorname{id},\operatorname{id}]$ Yes Yes
-Table B.2. The 13 isomorphism classes of GL-racks of order 3.
 Table B.3. The 62 isomorphism classes of GL-racks of order 4.